### v2

Abstract
Language Models (LMs) are increasingly used for a wide range of prediction tasks, but their training can often neglect rare edge cases, reducing their reliability. Here, we define a stringent standard of trustworthiness whereby the task algorithm and circuit implementation must be verified, accounting for edge cases, with no known failure modes. We show that a transformer model can be trained to meet this standard if built using mathematically and logically specified frameworks. In this paper, we fully verify a model for n-digit integer addition. To exhibit the reusability of verified modules, we insert the trained integer addition model into an untrained model and train the combined model to perform both addition and subtraction. We find extensive reuse of the addition circuits for both tasks, easing verification of the more complex subtractor model. We discuss how inserting verified task modules into LMs can leverage model reuse to improve verifiability and trustworthiness of language models built using them. The reuse of verified circuits reduces the effort to verify more complex composite models which we believe to be a significant step towards safety of language models.
1 Introduction
Figure 1: An overview of our methodology: (1) We trained a perfectly accurate 6-digit integer addition model. (2) We reverse-engineered the model to find the algorithms that were implemented to perform addition. (3) We inserted the addition model into a new model, by copying the weights of the attention heads and MLPs (in brown) into the larger model during initialisation. (4) We then train the new model on 80% subtraction and 20% addition questions. (5) We find that the resulting model reuses the inserted addition circuits for both addition and subtraction questions.
Transformer-based large language models (LLMs) are powerful (Barak et al., 2022) yet largely inscrutable due to their complex, nonlinear interactions in dense layers within high-dimensional spaces. Given this complexity, their deployment in critical settings (Zhang et al., 2022) highlights the need for understanding their behavior. Many argue that making these models interpretable is key to their safe use (Hendrycks & Mazeika, 2022).
Mechanistic interpretability focuses on demystifying and validating the algorithms behind model weights, translating complex computations into more human-understandable components (Raukur et al., 2022). This understanding aids in predicting model behavior in new situations and fixing model errors.
In creating and training a model, we would ideally want the model to have a very high standard of accuracy and trustworthiness. We aim to achieve this by holding the model to a standard that we term known-good. A model performing a task is known-good if:
  1. 1.
    The model’s algorithm for the task and the mechanisms it implements (the “circuits”) are understood.
  2. 2.
    All possible edge cases have been identified and tested. As a heuristic, the model should be able to perform the task one million times without error (“1M Q accuracy”).
  3. 3.
    The model has no known failure modes.
However, exhaustive testing of a model for most tasks is unfeasible. For instance, a task involving the addition of 5-digit integers (e.g., 12345+67890) presents ten billion possible variations. Rather than testing every variation, the task can be conceptualized within a formal framework to pinpoint edge cases the model must handle. For example, in 5-digit addition, the most uncommon edge case is 55555+44445=100000, which requires a carry bit to cascade through all digits, occurring in only 0.002% of cases. A known-good model must incorporate algorithms to manage this and other more frequent edge cases. A known-good model would hence have circuits that are verified to perform the task perfectly.
In this paper, we detail the development and interpretation of a known-good model for addition. Our findings indicate that the model constructs a specific circuit for each edge case, with these circuits sharing intermediate results. We confirm the validity of the entire set of circuits, ensuring they cover all identified edge cases. The model achieves a very low training loss and can do one million predictions without error, making it a known-good model.
Additionally, we develop a “mixed” model capable of both addition and subtraction, incorporating the known-good addition model. This mixed model extensively reuses the addition circuits for both operations and learns new circuits, facilitating the interpretation of the model’s algorithm. This approach marks progress toward a known-good model for both addition and subtraction.
Hence, our main contributions are three-fold:
  • A trustworthy known-good 5-digit addition model, with an algorithm extensible to n-digits.
  • A 6-digit addition and subtraction model that reuses established addition circuits for both operations, indicating steps toward a known-good model for these tasks.
  • A proof of concept for enhancing the trustworthiness of a larger model by integrating a smaller known-good transformer model.
2 Related Work
Mechanistic interpretability aims to reverse engineer neural networks to find interpretable algorithms that are implemented in a model’s weights (Olah et al., 2020). To better represent these algorithms, Elhage et al. (2021a) introduced a mathematical framework to represent how transformer attention heads can work with each other to implement complex algorithms.
To explain how (not where) a model algorithm is implemented, Jenner et al. (2023) recommends documenting a low-level computation graph (causal abstraction), with a mapping from the graph to the model nodes that implement the computation, with experimentation verification.
Investigative techniques such as ablation interventions, activation unembeddings (nostalgebraist, 2020), and sparse autoencoders (Nanda, 2023; Cunningham et al., 2023), underpinned by the more theoretical frameworks (Elhage et al., 2021b; Geva et al., 2022), provide tools to help confirm a mapping.
Investigating pre-trained LMs on Arithmetic. Even though basic arithmetic can be solved following a few simple rules, pre-trained LMs often struggle to solve simple math questions (Hendrycks et al., 2021). Stolfo et al. (2023) used causal mediation analysis to investigate how large pre-trained LMs like Pythia and GPT-J performed addition to solve word problems.
It is also possible to improve a model’s arithmetic abilities through training data enrichment. Liu & Low (2023) used supervised fine tuning (including enriched training data) to convert LLaMA 7B to Goat - a model that is “Good at Arithmetic Tasks” including addition and subtraction. They also showed the model’s performance relied in part on LLaMA’s consistent tokenization of digits.
Studying Toy Models for Arithmetic. Doing mechanistic interpretability on toy transformers can help to better isolate clear, distinct circuits given the highly specific experimental setup for the model studied. Nanda et al. (2023) studied modular addition in a one-layer transformer (where each number was represented by a distinct token) to better understand training dynamics.
Quirke & Barez (2024) detailed how a 1-layer, 3-head transformer model performed 5-digit addition questions. It showed a rare edge case (e.g. “77778+22222=100000” where a “carry 1” cascades through 4 digits) that the model couldn’t handle - showing the importance of understanding and testing all edge cases for trustworthiness.
Michaud et al. (2023) posits that many natural prediction problems decompose into a finite set of knowledge and skills that are “quantized” into discrete chunks (quanta). Models must learn these quanta to reduce loss. They posit that “understanding a network reduces to enumerating its quanta.” Schaeffer et al. (2023) provides useful ways to measure quanta in mathematical prediction problems.
3 Methodology
3.1 Mathematical Framework
Consider the task of adding two $n$-digit numbers together. Let us define the first number as $D=\{D_{n-1},D_{n-2},\ldots,D_{0}\}$ and the second number as $D^{\prime}=\{D^{\prime}_{n-1},D^{\prime}_{n-2},\ldots,D^{\prime}_{0}\}$. The resulting answer can be represented as $A=\{A_{n},A_{n-1},\ldots,A_{0}\}$. For an illustrative example, see Figure 2.
Figure 2: We refer to individual tokens in a 5-digit addition question as D4, .. D0, and D’4, .., D’0 and the answer tokens as A5, .., A0. For n-digits, we use the terms $D_{n}$, $D^{\prime}_{n}$and $A_{n}$.
We categorize the sum of each digit pair into three overlapping groups BA, MC and US defined as:
$BA_{n}=(D_{n}+D^{\prime}_{n})\mod 10$ (1)
$MC_{n}=\begin{cases}1&\text{if }(D_{n}+D^{\prime}_{n})\geq 10,\\ 0&\text{otherwise}.\end{cases}$ (2)
$US_{n}=\begin{cases}1&\text{if }(D_{n}+D^{\prime}_{n})=9,\\ 0&\text{otherwise}.\end{cases}$ (3)
When calculating $(A_{n})$, considering only the carry bit from the two preceding digits, we have:
$A_{n}=\begin{cases}D_{n}+D^{\prime}_{n}+1\underbrace{\text{mod}}_{10},&\text{if }\underbrace{MC}_{n-1}=1\text{ or}\\ &(\underbrace{US}_{n-1}=1\text{ and }\underbrace{MC}_{n-2}=1),\\ D_{n}+D^{\prime}_{n}\underbrace{\text{mod}}_{10},&\text{otherwise.}\end{cases}$ (4)
This approach is aligned with the framework implemented by the addition model in Quirke & Barez (2024). However, it is restricted to cases where the carry bit cascades over two or fewer digits. To address this limitation, we introduce an additional digit-level categorization called TriCase:
$D_{n}.C=\underbrace{\text{TriCase}}_{(D_{n},D^{\prime}_{n})}=\begin{cases}T10&\text{if }(D_{n}+D^{\prime}_{n})\geq 10,\\ T9&\text{if }(D_{n}+D^{\prime}_{n})=9,\\ T8&\text{if }(D_{n}+D^{\prime}_{n})\leq 8.\end{cases}$ (5)
These categorization cases contain carry bit information:
  • If $D_{n}.C=T10$, the digit-pair always causes a MC,
  • if $D_{n}.C=T9$, the digit-pair will cause an MC if the proceeding digit-pair sum generates an MC, and
  • if $D_{n}.C=T8$, the digit-pair never causes an MC .
To compute the first digit of the answer in a problem such as (55555+44445=)(55555+44445=)( 55555 + 44445 = ), the algorithm must synthesize the five TriCase values from the respective digit-pairs. This synthesis is crucial to determine the presence of a carry bit cascading through all five digits.
We introduce a function, TriAdd, to facilitate the transfer of TriCase data between digits. It is defined as follows:
$\underbrace{\text{TriAdd}}_{(D_{n}.C,D_{n-1}.C)}=\begin{cases}T10&\text{if }D_{n}.C=T10\\ &\text{or }D_{n}.C=T9\\ &\text{\quad and }D_{n-1}.C=T10,\\ T9&\text{if }D_{n}.C=T8\\ &\text{\quad and }D_{n-1}.C=T10,\\ T8&\text{otherwise.}\end{cases}$ (6)
TriAdd transfers MC data from $(D_{n-1})$ to $(D_{n})$ by integrating the values of $(D_{n}.C)$ and $(D_{n-1}.C)$. This function can be represented as nine bigram mappings and yields three distinct outputs (details in App. D). Finally, we introduce the recursive function $D_{n}$.$C_{m}$ , which aggregates TriCase data across $(m)$ digits to enhance accuracy.
$D_{n}.C_{m}=\begin{cases}\underbrace{\text{TriAdd}}_{(D_{n}.C,D_{n-1}.C_{m-1})}&\text{if }m>2,\\ \underbrace{\text{TriAdd}}_{(D_{n}.C,D_{n-1}.C)}&\text{if }m=2.\end{cases}$ (7)
Simply put $D_{n}$.$C_{m}$ cascades the carry bit over $m$ digits. Calculating $D_{1}.C_{2}$ means calculating $\text{TriAdd}(D_{1}.C,D_{0}.C)$ combining the $D_{0}$-specific TriCase value and $D_{1}$-specific TriCase value to give a perfectly accurate TriCase value for $D_{1}$. Calculating $D_{4}.C_{5}$ means calculating $\text{TriAdd}(D_{4}.C,D_{3}.C_{4})$ combining the $D_{4}$-specific TriCase value with the perfectly accurate $D_{3}.C_{4}$ TriCase value to give a perfectly accurate TriCase value for $D_{4}$.
In 5-digit addition, the values D4.C5, D3.C4, D2.C3 and D1.C2 are perfectly accurate as they integrate TriCase data all the way back to and including D0.MC . An accurate algorithm must calculate these values. Works on 5-digit model was unable to calculate these $D_{n}$.$C_{m}$ and so was not perfectly accurate (Quirke & Barez, 2024) . As shown later, our model’s algorithm incorporates the (Quirke & Barez, 2024) algorithm, and adds these perfectly accurate $D_{n}$.$C_{m}$ Figure 3 shows the combined algorithm.
Figure 3: In 5-digit addition the model first calculates all $D_{n}.C$ values, then combines them into perfectly accurate $D_{n}$.$C_{m}$ values by P11 (token =) when A5 is predicted. From P12, the remaining answer digits are calculated one by one, using $D_{n}$.$C_{m}$ values if needed. This algorithm can be extended to generalise to n-digits.
3.2 Question Difficulty and Frequency
To analyse question difficulty, we categorised questions by the complexity of the computation required to solve the question, as shown in Tab. 1. The categories are arranged according to the number of digits that a carry bit has to cascade through.
Table 1: We categorise addition questions into non-overlapping “calculation complexity” quanta, ordered by increased computational difficulty (and decreasing occurrence frequency). Five-digit addition questions quanta are S0 to S5. Ten-digit addition question quanta are S0 to S10. S10’s frequency is $\sim 3\times 10^{-4}$ showing the need to enrich training data for rare edge cases.
Name Contains Example Freq
S0 BA 11111+12345=23456 $\sim$5%
S1 BA,MC 11111+9=22230 $\sim$21%
S2 BA,MC,US 11111+89=22300 $\sim$34%
S3 BA,MC,USx2 11111+889=23000 $\sim$28%
S4 BA,MC,USx3 11111+8889=30000 $\sim$11%
S5 BA,MC,USx4 11111+88889=100000 $\sim$2%
3.3 Techniques
In investigating the circuits of the model, we want to understand what each attention head or MLP layer is doing across each token position, and how it relates to our mathematical framework. Hence, we define a node as the computation done by an attention head or MLP layer for a given token position.
Mean ablation. To find out how the model depends on the output of a node, we replace the output of that node with the vector that is the mean of all of its outputs across its batch.
Attention Patterns. To find out what the model attends to at that node, we take the attention pattern at that token position and take the significant tokens attended to ($>0.01$ post softmax).
4 Experiments
4.1 Training & Investigating Five-Digit Addition Model
Training the model. The 5-digit 1-layer addition model in Quirke & Barez (2024) achieved an accuracy of approximately 99%. Our preliminary experiments indicated that a 2-layer, 3-head model was the smallest configuration capable of achieving fully accurate addition (see App. C for alternative configurations tested).
This configuration effectively doubled the computational power compared to the 1-layer model. Moreover, a 2-layer model introduces the capability to “compose” the attention heads in novel ways, facilitating the implementation of more complex algorithms (Elhage et al., 2021b).
By enriching the training dataset with more instances of rare edge cases and using 30 thousand training batches, the 5-digit 2-layer model attained a final training loss of approximately $2.3\times 10^{-8}$. This model demonstrated a 1M Q accuracy (details in App. B).
Investigating the model. To gain insights into the model’s predictive behavior, we conducted an investigation focusing on its activity during prediction calculations. Ablation experiments targeting the nodes revealed that the model primarily depends on nodes located in token positions D’2 to A1, as illustrated in Figs 4 and 5. Across these nine positions, the model engages 36 nodes, comprising 21 attention head nodes and 15 MLP layer nodes. The effects of node ablation on our complexity and answer-impact metrics were analyzed (see Figs 6 and 7), elucidating the specific computations performed at each node.
For each answer digit $A_{n}$using test questions corresponding to the T8, T9 and T10 categories, we performed Principal Component Analysis (PCA) on the nodes yielding interpretable results. Specifically nine “node and answer-digit” combinations (see Figs 13 and 14) showed strong clustering of the questions aligned to the T8, T9 and T10 categories, suggesting these nodes are implementing the TriCase categorisation.
Figure 4: This map shows the % of enriched questions that fail when we ablate each node in the 5-digit 2-layer 3-head addition model. The model only uses nodes in token positions P8 to P16 (tokens D’2 to A1). No attention heads in layer 1 are used. Lower percentages correspond to rarer edge cases. The grey space represents nodes that are not used by the model.
Figure 5: This map shows the input tokens each attention head attends to at each token position. At token position P11 (token =) the model predicts the first answer digit A5. All digit pairs (e.g. D2 D’2) are attended to by P11.
Figure 6: This map shows the lowest complexity quanta S0, S1, etc impacted when we ablate each node in the 5-digit 2-layer 3-head addition model. To answer S0 questions, only the S0 nodes are used. To answer S1 questions, S0 and S1 nodes are used, etc.
Figure 7: This map shows the answer digit(s) A0 .. A5 impacted when we ablate each useful node in the 5-digit 2-layer 3-head addition model. The “column per digit” pattern in the last 6 columns mirrors the 1-layer model’s pattern. Cells containing values like A2..5 (in column D’0) are used in multiple prediction steps to calculate multiple answer digits e.g. A2 to A5.
The algorithm predicts the first answer digit, A5, at position P11 (tokne =). This digit, which can only be 0 or 1, is the most challenging to predict as it may rely on a cascade starting from the digit pair D0 D’0 (e.g., 55555+44445=100000). The algorithm is required to compute this cascade using the nodes located in positions P8 to P11 (input tokens D’2 to =). As illustrated in Figure 5, these nodes attend to all digit pairs from D4 D’4 to D0 D’0 by P11. Additionally, the PCA data, as shown in Figures 13 and 14, suggest that these nodes produce tri-state outputs.
After encountering several initial setbacks (see App. E and F), we discovered that the model utilizes a minimal set of “make carry” information, leading to the development of the TriCase quanta.
The model performs $D_{n}.C$ using bigrams (see App. D) to map two input tokens to one result token e.g. “6” + “7” = T10. In positions P8 to P11, the model does $D_{n}.C$ calculations on all digit pairs from D4 D’4 to D0 D’0.
The mathematical framework behind this is intricate. To validate it, we implemented the algorithm in Python and successfully executed one million additions with zero errors.
For a specific instance of the 5-digit addition model, we mapped the algorithm to individual nodes. Using ablation interventions, we verified each node’s role in executing the expected algorithmic step (see App. G). Table 2 and Figure 1 illustrate how the algorithm was applied to this particular model instance, adhering to all known constraints and accurately calculating all six answer digits.
Table 2: Pseudo-code for accurately predicting all answer digits, with mappings to the model nodes that implement the pseudo-code, for a specific 5-digit model instance.
Value Value calculated as Nodes used
D2.C
TriCase(D2, D’2)
P8.L0.H1 and P8.L0.MLP
D1.C
TriCase(D1, D’1)
P9.L0.H1 and P9.L0.MLP
D1.C2
TriAdd(D1.C, TriCase(D0, D’0))
P10.L0.H1 and P10.L0.MLP
D3.C4
TriAdd(TriCase(D3, D’3), TriAdd(D2.C,D1.C2))
P11.L0.H1
D4.C
TriCase(D4, D’4)
P11.L0.H2
D4.C5
TriAdd(D4.C, D3.C4)
P11.L0.MLP
A5
(D4.C5 == 10)
P11.L1.MLP
D4.BA
(D4 + D’4) % 10
P12.L0.H0 + H2
D3.C4
TriAdd(TriCase(D3, D’3), TriAdd(D2.C,D1.C2))
P12.L0.H1
A4
(D4.BA + (D3.C4 / 10)) % 10
P12.L0.MLP and P12.L1.MLP
D3.BA
(D3 + D’3) % 10
P13.L0.H0 + H2
D2.C3
TriAdd(D2.C,D1.C2)
P13.L0.H1
A3
(D3.BA + (D2.C3 / 10)) % 10
P13.L0.MLP and P13.L1.MLP
D2.BA
(D2 + D’2) % 10
P14.L0.H0 + H2
D1.C2
(D1 + D’1) / 10 + P10.D1.C2 *
P14.L0.H1
A2
(D2.BA + (D1.C2 / 10)) % 10
P14.L0.MLP and P14.L1.MLP
D1.BA
(D1 + D’1) % 10
P15.L0.H0 + H2
D0.MC
(D0 + D’0) / 10
P15.L0.H1
A1
(D1.BA + D0.MC) % 10
P15.L0.MLP and P15.L1.MLP
A0
(D0 + D’0) % 10
P16.L0.H0 + H2 P16.L0.MLP and P16.L1.MLP
Throughout the training of different instances of the 5-digit models, small modifications to the training environment resulted in subtly varied implementations. All variations seen use the Figure 3 algorithm, but each has small implementation differences. For example, in positions P8 to P11, the model might alter which 5 nodes perform the $D_{n}.C$ calculations and in what order. This corresponds to small changes in the Tab. 2 documentation per variation.
In summary, we concluded that the specific instance of the 5-digit addition model documented above is reliably accurate, well-understood, well-functioning and so known-good.
4.2 Case Study: Extending to Six-Digit Addition
Our investigations showed that all instances of the 5-digit addition model we trained appeared to utilize the same algorithm. To determine if this algorithm extends to related models, we trained a 6-digit 2-layer 3-head addition model, employing a different seed and optimizer. Additionally, we modified the answer format to include a ’+’ token (e.g., 111111+222222=+0333333).
Employing similar quanta and training data enrichment techniques, we developed a 6-digit addition model. This model achieved a loss of approximately $9.1\times 10^{-9}$ and demonstrated 1M Q accuracy.
The 5-digit (see Figures 4,  5,  6, and  7) and 6-digit (see Figures 15,  16,  8, and  17) quanta maps show numerous similarities, suggesting the 6-digit model implements the logical extension of the 5-digit algorithm (see Figure 3).
Intervention ablation tests show the 6-digit model does 6 BA and 6 MC calculations in late tokens, and does 6 $D_{n}.C$ calculations in early tokens with PCA showing the expected TriCase output. The 5 $D_{n}$.$C_{m}$ calculations are still to be tested.
Give this model has 1M Q accuracy, if this testing is successfully completed, and the algorithm implementation detail is documented, we can state the 6-digit model is known-good.
Figure 8: This map shows the simpliest addition complexity quanta S0, S1, etc used in each useful node of the 6-digit 2-layer 3-head addition model. Note the similarity to 5-digit Figure 6
4.3 Case Study: Extending the model for subtraction
To explore the reuse of known-good models, we developed a larger (6-digit 3-layer 4-head) model capable of performing both addition and subtraction. We introduced specific subtraction quanta and enriched the training dataset with rare subtraction edge cases.
This new “mixed” model, initially untrained, was initialized with the weights from the (6-digit 2-layer 3-head) addition model. To facilitate its learning of subtraction while maintaining its pre-inserted addition skills, we trained the mixed model with a dataset comprising 80% subtraction and 20% addition problems. This approach resulted in a model achieving 1M Q accuracy for both addition and subtraction tasks. Attempts to “freeze” useful attention heads and/or MLP layers from the addition model by periodically copying their weights back into the mixed model every 100 training epochs, however, led to decreased accuracy.
Subtraction ($a-b$) presents a greater challenge than addition ($a+b$), primarily due to its two distinct cases: $a>b$ (positive answer) and $a<b$ (negative answer). The model must discern this difference to accurately determine the first answer token as either “+” or “-”.
In cases of “positive” subtraction, the process closely mirrors addition. This observation led us to define subtraction quanta (see Tables 3 and 4 (see App. J), paralleling the addition quanta.
Calculating the “-” case in subtraction is more complex than the “+” case. The model might employ distinct circuits for each scenario. Our hypothesis is that after mastering the simpler “+” case, the model will learn to apply the mathematical identity $a-b=-(b-a)$, thus enabling it to handle “-” case questions by re-utilizing the “+” case circuits and then inverting the result. To examine this, we introduced the NG quanta (see Table 4) and again applied mean-ablation techniques to analyze the process.
Table 3: We categorise digit-pair differences (i.e. $D_{n}$- $D^{\prime}_{n}$ abbreviated as $D_{n}$.Diff) into three groups. These subtraction groups parallel the BA, MC and US addition groups. Groups BS and DZ overlap.
Name Abbrev Definition E.g. Like
Base Subtract BS $D_{n}$.Diff $<=0$ 7 - 2 BA
Borrow 1 B1 $D_{n}$.Diff $<0$ 7 - 9 MC
Diff Zero DZ $D_{n}$.Diff = 0 7 - 7 US
Table 4: We categorise subtraction questions into non-overlapping “calculation complexity” quanta, ordered by increased computational difficulty and decreasing occurrence frequency. Five-digit subtraction questions quanta are M0 to M4. Ten-digit subtraction question quanta are M0 to M9, and M9’s frequency is $\sim$0.000001%. Subtraction questions where the answer is negative get their own category NG .
Name Contains Example Freq
M0 BS 33333-22222=+11111 $\sim$25%
M1 BS,B1 33333-22282=+11051 $\sim$15%
M2 BS,B1,DZ 33333-22382=+10951 $\sim$8%
M3 BS,B1,DZx2 33333-23382=+09951 $\sim$2%
M4 BS,B1,DZx3 33333-23338=+09995 $\sim$0.1%
NG N/A 10000-90000=-80000 $\sim$50%
We compared the nodes utilized by the inserted addition model (see Figure 6) to those used by the mixed model when performing addition (see Figure 9). There is a strong correlation between them. Notably, the mixed model predominantly uses the existing attention heads in layers 1 and 2, ignoring many available spare heads.
Further comparison was made between the nodes the mixed model employs for addition and subtraction (see Figures 9 and 10). The nodes used for addition (S*) and those for “subtraction with a positive answer” (M*) show a very strong correlation. This suggests that the model repurposes its addition nodes to facilitate subtraction in some manner. Notably, an additional 15 nodes are activated for subtraction problems that result in a negative answer (represented by the NG quanta).
While the mixed model has 1M Q accuracy for both addition and subtraction tasks, intervation ablation testing and PCA of the mixed model have not been completed, and the algorithm implementation has not been documented. We cannot yet assert that it is known-good.
Figure 9: This map shows the simplest complexity quanta S0, S1, etc used in each useful node of the 6-digit 3-layer 4-head mixed model when doing addition questions.
Figure 10: This map shows the simpliest complexity quanta M0, M1, etc used in each useful node of the 6-digit 3-layer 4-head mixed model for subtraction questions with positive answers. The NG cells are only used in subtractions $a-b$ where $a<b$, likely implementing the $-(b-a)$ identity as described in  4.3.
5 Conclusion
In this paper, we successfully trained a 5-digit 2-layer 3-head addition model that we have verified as known-good. Our investigation revealed that a 6-digit 2-layer 3-head addition model follows the same algorithmic principles, albeit with some implementation variations.
A significant aspect of our work was demonstrating component reuse. We achieved this by integrating an existing addition model into a new “mixed” model designed for both addition and subtraction tasks. This integration allowed us to leverage our in-depth understanding of the addition model’s algorithm to unravel the workings of the mixed model. Our findings indicate that the mixed model employs the nodes of the inserted addition model for conducting both addition and subtraction computations.
This paper lends support to the assertion by Michaud et al. (2023) that many natural prediction problems can be broken down into a finite set of “quanta” computations. These quanta are essential for models to learn in order to minimize loss. Furthermore, it aligns with the concept that comprehending a network’s functionality is fundamentally about identifying and understanding its sub-quanta.
6 Limitations and Future Work
Our research opens up several promising avenues for future exploration. A key objective is to establish the mixed model as known-good by reverse-engineering its algorithm implementation and validating it through experimental methods. Additionally, we aim to develop a comprehensive model that proficiently handles addition, subtraction, multiplication, and division, ensuring its status as known-good.
Currently, our approach requires significant human effort, which may not be feasible for large-scale applications. However, we believe applying this method to critical tasks, such as logical reasoning involving both simple (e.g., AND, OR) and more complex operators (e.g., Converse, Inverse), is justified. This could lead to the creation of more known-good transformer models, potentially extending to frameworks for ethical decision-making.
A promising finding of our research was the successful insertion of a known-good model into a larger, untrained model while retaining its functionality. We think it might be feasible to embed a compact known-good model into an existing Large Language Model (LLM) to enhance its trustworthiness. Previous studies (Stolfo et al., 2023; Kruthoff, 2024) found nodes in LLMs that were responsible for inaccuracies mathematical calculations. Given our findings, it might be possible to replace these nodes with known-good ones. While this poses an engineering challenge, research by (Voita et al., 2023) and (Hu et al., 2021) indicates the presence of “spare” neurons in LLMs and the potential for integrating small-scale modifications to fix erroneous circuits in LLMs.
Moreover, developing methods to incorporate compact known-good models into LLMs, whether untrained or trained, could democratize AI Safety research. It enables small teams to focus on specific functional areas, crafting component models to a defined quality standard for integration into various LLMs.
A significant challenge in our research was the sensitivity of model behavior to minor changes in the training environment. These changes often altered the specific nodes used for calculations, necessitating manual adjustments to our algorithm-validation test suite. A more effective strategy would be a declarative approach to describing each component in an algorithm hypothesis, allowing automatic identification of corresponding model nodes. For instance, a component description could specify nodes that primarily attend to question digits D2 and D’2, influence answer digit A3, and are crucial for S0 but not S1 complexity questions. Such a declarative language and its accompanying tooling would accelerate research progress and ensure that known-good components retain their functionality when integrated into LLMs. An initial version of this tooling has been created, but significant further work is required to fully realize its potential.
Impact Statement
Our work aims to explain the inner workings of transformer-based language models, which may have broad implications for a wide range of applications. A deeper understanding of generative AI has dual usage. While the potential for misuse exists, we discourage it. The knowledge gained can be harnessed to safeguard systems, ensuring they operate as intended. It is our sincere hope that this research will be directed towards the greater good, enriching our society and preventing detrimental effects. We encourage responsible use of AI, aligning with ethical guidelines.
Acknowledgements
The authors would like to thank Apart Labs (https://apartresearch.com/lab) for their support in producing this paper and more generally supporting the entry of people new to AI Safety into the field.
Appendix A Appendix: Model Configuration
Addition experiments were done in Colab notebooks named “Accurate Addition - Train” and “Accurate Addition - Analyse”. They run on a T4 GPU. Training takes $\sim$15 mins. Analysis takes $\sim$8 mins. The key parameters (which are all configurable) are:
  • n_layers = 2: This is a 2-layer Transformer model
  • n_heads = 3: There are 3 attention heads
  • n_digits = 5: Number of digits in the addition question
Training uses a new batch of data each step (aka Infinite Training Data) to minimise memorisation. During a training run the model processes about 1.5 million training datums. For the 5-digit addition problem there are 100,000 squared (that is 10 billion) possible questions. So the training data is much less than 1% of the possible problems. Because US cascades (e.g. 44445+55555=100000, 54321+45679=1000000, 44450+55550=10000, 1234+8769=10003) are exceedingly rare, the data generator was enhanced to increase the likelihood of these cases turning up in the training data
Mixed operation (addition and subtraction) experiments were done in Colab notebooks named “Accurate Math - Train” and “Accurate Math - Analyse”. They run on a T4 GPU. Training takes $\sim$25 mins. Analysis takes $\sim$8 mins. The key parameters (which are all configurable) are:
  • n_layers = 3: This is a 3-layer transformer model
  • n_heads = 4: There are 4 attention heads
  • n_digits = 6: Number of digits in the addition and subtraction questions
Training uses a new batch of data each step (aka Infinite Training Data) to minimise memorisation.
In all Colabs, each digit is represented as a separate token. (Liu & Low, 2023) state that LLaMa’s “remarkable arithmetic ability … is mainly atributed to LLaMA’s consistent tokenization of numbers”.
Appendix B Appendix: Addition Model Loss
The 1-layer model is not very accurate (see Tab. 5). With 2 layers the model gains accuracy (see Tab. 6). The model uses batch_size = 64, n_heads = 3, lr = 0.00008, weight_decay = 0.1. The loss function is simple:
  • Per Digit Loss: For “per digit” graphs and analysis, for a given answer digit, the loss used is negative log likelihood.
  • All Digits Loss: For “all answer digits” graphs and analysis, the loss used is the mean of the “per digit” loss across all the answer digits.
Table 5: Results for 1-layer 3-head model with 5, 10 & 15 digits
No. digits 5 10 15
Training batches 20K 20K 20K
Loss 0.008314 0.040984 0.031746
Heads used 15 29 46
MLPs used 6 11 16
Table 6: Results for 2-layer 3-head model with 5 & 10 digits
No. digits 5 5 10 10
Training batches 20K 30K 20K 30K
Loss 2.7e-6 2.3e-8 9.4e-5 1.2e-4
Heads used 21 19 57 51
MLPs used 15 15 27 28
The 2-layer, 5-digit model trained over 30 thousand batches has a final training loss of $\sim 2.3\times 10^{-8}$.
Appendix C Appendix: Addition Model Shape Choice
While we wanted a very low loss addition model, we also wanted to keep the model compact - intuiting that a smaller model would be easier to understand than a large model. Here are the things we tried to reduce loss that didn’t work:
  • Increasing the frequency of hard (cascading US ) examples in the training data so the model has more hard examples to learn from. This improved training speed but did not reduce loss.
  • Increasing the number of attention heads from 3 to 4 or 5 (while still using 1 layer) to provide more computing power.
  • Changing the question format from “12345+22222=” to “12345+22222equals” giving the model more prediction steps after the question is revealed before it needs to state the first answer digit.
  • With n_layers = 1 increasing the number of attention heads from 3 to 4.
  • Changing the n_layers to 2 and n_heads to 2.
The smallest model shape that did reduce loss significantly was 2 layers with 3 attention heads
Appendix D Appendix: TriAdd Implementation
TriAdd transfers data from $(D_{n-1})$ to $(D_{n})$ by integrating the values of $(D_{n}.C)$ and $(D_{n-1}.C)$. This function can be represented as nine bigram mappings with three possible outputs.
Table 7: $D_{n}$.TriAdd can be calculated from $D_{n}.C$ and $D_{n-1}.C$ through nine bigram mappings and yielding the three distinct outputs T8, T9 and T10
$D_{n}.C$ =T8 $D_{n}.C$ =T9 $D_{n}.C$ =T10
$D_{n-1}.C$ = T8 T8 T9 T10
$D_{n-1}.C$ = T9 T8 T9 T10
 $D_{n-1}.C$ = T10 T9 * T10 T10
Note that in the case $D_{n}.C$ = T9 and $D_{n-1}.C$ = T10, the answer is indeterminate. The result could be T8 or T9 but importantly it can not be T10. We choose to use T9 in our definition, but T8 would work just as well.
Appendix E Appendix: Addition Hypothesis 1
Given the 2-layer attention pattern’s similarity to 1-layer attention pattern, and the above evidence, our first (incorrect) hypothesis was that the 2-layer algorithm:
  • Is based on the same BA, MC and US operations as the 1-layer.
  • Uses the new early positions to (somehow) do the US calculations with higher accuracy than the 1-layer model.
  • The long double staircase still finalises each answer digit’s calculation.
  • The two attention nodes in the long double staircase positions do the BA and MC calculations and pull in US information calculated in the early positions.
If this is correct then the 2-layer algorithm successfully completes these calculations:
  • A0 = D0.BA
  • A1 = D1.BA + D0.MC
  • A2 = D2.BA + (D1.MC or (D1.MS & D0.MC))
  • A3 = D3.BA + (D2.MC or (D2.MS & D1.MC) or (D2.MS & D1.MS & D0.MC))
  • A4 = D4.BA + (D3.MC or (D3.MS & D2.MC) or (D3.MS & D2.MS & D1.MC) or (D3.MS & D2.MS & D1.MS & D0.MC))
  • A5 = D4.MC or (D4.MS & D3.MC) or (D4.MS & D3.MS & D2.MC) or (D4.MS & D3.MS & D2.MS & D1.MC) or (D4.MS & D3.MS & D2.MS & D1.MS & D0.MC)
Our intuition is that there are not enough useful nodes in positions 8 to 11 to complete the A5 calculation this way. So we abandoned this hypothesis.
Appendix F Appendix: Addition Hypothesis 2
Our second (incorrect) hypothesis was that the 2-layer algorithm has a more compact data representation, so it can pack more calculations into each node, allowing it to accurately predict A5 in step 11.
We claimed the model stores the sum of each digit-pair as a single token in the range “0” to “18” (covering 0+0 to 9+9). We name this operator $D_{n}$.T, where T stands for “token addition”:
  • $D_{n}$.T = $D_{n}$+ $D^{\prime}_{n}$
The $D_{n}$.T operation does not understand mathematical addition. Tab. 8 shows how the model implements the T operator as a bigram mapping.
Table 8: Implementing the T operator as a bigram mapping from 2 input tokens to 1 result token.
$D_{n}$vs $D^{\prime}_{n}$ 0 1 4 5 8 9
0 0 1 4 5 8 9
1 1 2 5 6 9 10
4 4 5 8 9 12 13
5 5 6 9 10 13 14
8 8 9 12 13 16 17
9 9 10 13 14 17 18
$D_{n}$.T is a compact way to store data. Tab. 9 show how, if it needs to, the model can convert a $D_{n}$.T value into a one-digit-accuracy BA, MC or MS value.)
Table 9: Converting a $D_{n}$.T value into a BA , MC or MS value.
$D_{n}$.T $D_{n}$.BA $D_{n}$.MC $D_{n}$.MS
0 0 0 0
1 1 0 0
8 8 0 0
9 9 0 1
10 0 1 0
17 7 1 0
18 8 1 0
Our notation shorthand for one-digit-accuracy these “conversion” bigram mappings is:
  • $D_{n}$.BA = ($D_{n}$.T % 10) where % is the modulus operator
  • $D_{n}$.MC = ($D_{n}$.T // 10) where // is the integer division operator
  • $D_{n}$.MS = ($D_{n}$.T == 9) where == is the equality operator
The D0.T value is perfectly accurate. But the other $D_{n}$.T values are not perfectly accurate because each is constrained to information from just one digit. We define another more accurate operator $D_{n}$.T2 that has “two-digit accuracy”. $D_{n}$.T2 is the pair sum for the nth digit plus the carry bit (if any) from the n-1th digit T:
  • $D_{n}$.T2 = $D_{n}$.T + $D_{n-1}$.MC
$D_{n}$.T2 is more accurate than $D_{n}$.T. The $D_{n}$.T2 value is always in the range “0” to “19” (covering 0+0+0 to 9+9+MakeCarry). Tab. 10 show how the model can implement the T2 operator as a mapping.
Table 10: Calculating $D_{n}$.T2 from $D_{n}$.T and $D_{n-1}$.T
$D_{n}$.T $D_{n}$.T2 if $D_{n-1}$.MC==0 $D_{n}$.T2 if $D_{n-1}$.MC==1
0 0 1
1 1 2
9 9 10
10 10 11
17 17 18
18 18 19
Following this pattern, we define operators $D_{n}$.T3, $D_{n}$.T4 and $D_{n}$.T5 with 3, 4 and 5 digit accuracy respectively:
  • $D_{n}$.T3 = $D_{n}$.T + ( $D_{n-1}$.T2 // 10 ) aka $D_{n}$.T + $D_{n-1}$.MC2
  • $D_{n}$.T4 = $D_{n}$.T + ( $D_{n-1}$.T3 // 10 ) aka $D_{n}$.T + $D_{n-1}$.MC3
  • $D_{n}$.T5 = $D_{n}$.T + ( $D_{n-1}$.T4 // 10 ) aka $D_{n}$.T + $D_{n-1}$.MC4
The value D4.T5 is perfectly accurate as it integrates MC and cascading MS data all the way back to and including D0.T. The values D1.T2, D2.T3, D3.T4 are also all perfectly accurate. If the model knows these values it can calculate answer digits with perfect accuracy:
  • A1 = D1.T2 % 10 with zero loss
  • A2 = D2.T3 % 10 with zero loss
  • A3 = D3.T4 % 10 with zero loss
  • A4 = D4.T5 % 10 with zero loss
  • A5 = D4.T5 // 10 with zero loss
Tab. 11 shows how the hypothesis 2 mathematical framework is applied and Figure 11 shows how this maps to the model. There are nodes that the model needs that this framework does not use, so our hypothesis is not 100% right.
Table 11: Accurately predicting A5 in step 11 and A4 in step 12 using addition hypothesis 2
Calculation Node
D1.T = D1 + D’1
P9.L0.H1 and P9.L0.MLP
D2.T = D2 + D’2
P8.L0.H1 and P8.L0.MLP
D1.T2 = D2.T + (D1.T + (D0 + D0’) // 10) // 10
P10.L0.H1 and P10.L0.MLP
D3.T4 = D3 + D’3 + D2.T3 // 10
P11.L0.H1
D4.T = D4 + D’4
P11.L0.H2
D3.MC3 = D3.T4 // 10
P11.L0.MLP
A5 = (D4.T + D3.MC3) // 10
P11.L1.MLP
D4.T5 = D4.T + D3.T4 // 10
P12.L0.H0
A4 = D4.T5 % 10 // 10
P12.L0.MLP
Figure 11: The 5-digit 2-layer integer addition model algorithm under hypothesis 2
In this hypothesis, all the answer digits are perfectly calculated using the nodes in positions 8 to 11. This hypothesis 2 is feasible, elegant and compact - reflecting the authors (human) values for good code.
Experimenation shows the model does not implement this hypothesis. It retains the long staircase BA calculations in positions 11 to 16. Why? Two reasons suggest themselves:
  • Hypothesis 2 is too compact. The model is not optimising for compactness. The long staircase is discovered early in training, and it works for simple questions. Once the overall algorithm gives low loss consistently it stops optimising.
  • Hypothesis 2 accurately predicts all answer digits in step 11 - reflecting the authors (human) values for good code. The model is not motivated to do this. It just needs to accurately predict A5 as 1 or 0 in step 11 and A4 in step 12 - nothing more.
We abandoned this hypothesis.
Appendix G Appendix: Five-Digit Addition Experiments
Figure 12: Our 5-digit integer addition question has 12 question tokens ($D_{n}$,$D^{\prime}_{n}$) followed by 6 answer tokens ($A_{n}$). This known-good 2-layer 3-head model can perform one million predictions without error. The model’s algorithm co-ordinates computation across 9 token positions ($P_{n}$) using 19 attention heads ($H_{n}$) and 15 MLP layers. Three attention heads produce intermediate results that are reused in multiple prediction steps e.g. P10.L0.H1 is reused in 4 prediction steps to calculate answer digits A5 to A2.
The hypothesis 3 pseudo-code was derived iteratively by obtaining experimental results and mapping them to mathematical operations. Figure 12 shows the pseudo-code diagramtically. Some of the experiments and mappings were:
  • Ablation experiments show that the A5 value is accurately calculated in prediction step 11 using 5 attention heads and 5 MLP layers. The pseudo-code accurately calculates A5 while constraining itself to this many steps.
  • Ablating the nodes one by one shows which answer digit(s) are reliant on each node (Ref Figure 7). Most interestingly, ablating P10.L0.H1 impacts the answer digits A5, A4, A3, A2 (but not A1 and A0). This node is used in the calculation of A5, A4, A3, A2 in prediction steps 11, 12, 13 and 14. These relationships are constraints that are all obeyed by the pseudo-code.
  • The pseudo-code has 4 instances where $D_{n}.C$ is calculated using TriCase. PCA of the corresponding nodes (P8.L0.H1, P9.L0.H1, P11.L0.H2 and P14.L0.H1) shows tri-state output for the specified $D_{n}$. (see Figure 13).
  • The pseudo-code has 4 instances where compound functions using TriCase and TriAdd to generate tri-state outputs. PCA of the corresponding nodes (P11.L0.H1, P12.L0.H1 and P13.L0.H1) shows tri-state output for the specified $D_{n}$. (see Figure 13).
  • Activation patching (aka interchange intervention) experiments at attention head level confirmed some aspects of the calculations (see § H for details.
  • The pseudo code includes calculations like D1.C which it says is calculated in P9.L0.H1 and P9.L0.MLP. Ablation tells us both nodes are necessary. For the attention head we use the PCA results for insights. We didn’t implement a similar investigative tool for the MLP layer, so in the pseudo-code we attribute the calculation of D1.C to both nodes.
  • For P10.L0.H1, the attention head PCA could represent either a bi-state or tri-state output (see Figure 14). The MLP layer at P10.L0.MLP could map the attention head output to either a bi-state or tri-state. We cannot see which. The pseudo-code shows a tri-state calculation at P10.L0.MLP, but with small alterations the pseudo-code would work with a bi-state output.
  • For P15.L0.H1 the attention head PCA could represent either a bi-state or tri-state output (see Figure 14). The pseudo-code shows a bi-state calculation D0.MC at P15.L0.H1, but with small alterations the pseudo-code would work with a tri-state output.
  • The calculation of D1.C2 in P14.L0.H1 is a interesting case. The model needs D1.C2 for A2 accuracy. The model could simply reuse the accurate D1.C2 value calculated in P10. Activation patching shows that it does not. Instead the P14 attention heads calculate D1.C1 from D1 and D’1 directly, and only relies on the P10.D1.C2 value in the case where D1.C2 != D1.C. That is, the calculation is “use P14.D1.C1 value else use P10.D1.C2 values”. This aligns with the model learning the P10.D1.C calculation early in training (for 90% accuracy) and later learning that P10.D1.C2 contains additional information it can use to get to 100% accuracy.
Figure 13: For 5-digit addition, for these attention heads, for exactly 1 answer digit $A_{n}$each, PCA shows these interpretable results. The dot colours show the TriCase value of each question. The PCA data and TriCase quanta are both tri-state and strongly correlated.
Figure 14: For 5-digit addition, for these attention heads, for exactly 1 digit $A_{n}$each, PCA shows these interpretable results. The dot colours show the TriCase value of each question. The PCA and TriCase data are strongly correlated, but the PCA data has 2 states.
Appendix H Appendix: Addition Interchange Interventions
To test the hypothesis 3 mapping of the mathematical framework (casual abstraction) to the model attention heads, various “interchange interventions” (aka activation patching) experiments were performed on the model, where
  • A particular claim about an attention head has selected for testing.
  • The model predicted answers for sample test questions, and the attention head activations were recorded (stored).
  • The model then predicted answers for more questions, but this time we intervened during the prediction to override the selected attention head activations with the activations from the previous run.
Using this approach we obtained the findiings in Tab. 12.
Table 12: Interchange Interventions experiments used activation patching to test the claims addition hypothesis 3 made for each attention head. Experimental results are consistent with hypothesis 3 for all nodes.
Nodes Claim: Attention head(s) perform … Finding: Attention head(s) perform a function that …
P8.L0.H1 and MLP
D2.C = TriCase( D2, D’2 ) impacting A4 and A5 accuracy
Based on D2 and D’2. Triggers on a D2 carry value. Provides carry bit used in A5 and A4 calculation.
P9.L0.H1 and MLP
D1.C = TriCase( D1, D’1 ) impacting A5, A4 and A3 accuracy
Based on D1 and D’1. Triggers on a D1 carry value. Provides carry bit used in A5, A4 and A3 calculation.
P10.L0.H1 and MLP
D1.C2 = TriAdd( D1.C, TriCase(D0, D’0) ) impacting A5, A4, A3 and A2 accuracy
Based on D0 and D’0. Triggers on a D0 carry value. Provides carry bit used in A5, A4, A3 and A2 calculation.
P11.L0.H1 and MLP
D3.C4 = TriAdd( TriCase( D3, D’3 ), TriAdd( D2.C, D1.C2)) impacting A5 accuracy
Based on D3 and D’3. Triggers on a D3 carry value. Provides carry bit used in A5 calculations.
P11.L0.H2 and MLP
D4.C = TriCase(D4, D’4) impacting A5 accuracy A4
Based on D4 and D’4. Triggers on a D4 carry value. Provides carry bit used in A5 calculation.
P12.L0.H0+H2 and MLP
D4.BA = (D4 + D’4) % 10 impacting A4 accuracy A4
Sums D4 and D’4. Impacts A4.
P13.L0.H0+H2 and MLP
D3.BA = (D3 + D’3) % 10 impacting A3 accuracy
Sums D3 and D’3. Impacts A3.
P14.L0.H0+H2 and MLP
D2.BA = (D2 + D’2) % 10 impacting A2 accuracy
Sums D2 and D’2. Impacts A2.
P14.L0.H1 and MLP
(D1 + D’1) / 10 + P10.D1.C2 info impacting A2 accuracy
Calculates P10.D1.C1 but add P10.D1.C2 info when D1.C != D1.C2. Impacts A2
P15.L0.H0+H2 and MLP
D1.BA = (D1 + D’1) % 10 impacting A1 accuracy
Sums D1 and D’1. Impacts A1.
P15.L0.H1 and MLP
D0.MC = (D0 + D’0) / 10 impacting A1 accuracy
Triggers when D0 + D’0 >10. Impacts A1 digit by 1
P16.L0.H0+H2 and MLP
D0.BA = (D0 + D’0) % 10 impacting A0 accuracy
Sums D0 and D’0. Impacts A0.
Appendix I Appendix: Six-Digit Addition
We trained a very low loss 6-digit Addition model. We compared the 5-digit and 6-digit model algorithms. Figures 6 and  8 are very instructive. So are Figures 7 and  17. We concluded that the 6-digit model used the same logical algorithm as the 5-digit model. There were some implementation differences:
  • The models selected different attention heads in the early positions to use to do the same logical calculations.
  • The 5-digit model used 2 attention heads per digit to do the BA calculation, whereas 6-digit model only uses one (and so is more compact).
  • The PCA trigrams have difference appearances in the two models. Figures 13 and  18.
  • For 6-digits, for A2 (only), the BA calculation is done in P18, but the MC calculation is not in P18. This is unique. The 6-digit model has optimised out the MC and instead relies solely on the accurate D1.C2 value (and so is more compact).
Figure 15: This map shows the % of questions that fail when we ablate each node in the 6-digit 2-layer 3-head addition model. The model only uses nodes in token positions P11 to P20. Lower percentages correspond to rarer edge cases. The grey space represents nodes that are not useful.
Figure 16: This map shows the input tokens each attention head attends to at each token position. The model predicts the first answer digit A6 in token position P14, and all digit-pair sums (e.g. D2 D’2) are attended to in P11 to P14. Note the similarity to Figure 5.
Figure 17: This map shows the answer digit(s) A0 .. A5 impacted when we ablate each useful node in the 6-digit 2-layer 3-head addition model. Note the similarity to Figure 7.
Figure 18: For 9 attention heads, for exactly 1 answer digit $A_{n}$each, PCA shows these interpretable results for the 6-digit Addition model. The 3 dot colours show the TriCase value of each question. The PCA data and TriCase quanta are both tri-state and strongly correlated.
Appendix J Appendix: Six-Digit Subtraction
Integer subtraction (e.g. 654321-789012=-0134691) and integer addition have similarities. Subtraction is conceptually more complex than integer addition. While $a+b$ and $b+a$ give the same answer, generally $a-b$ and $b-a$ give difference answers. In human subtraction the algorithm used to calculate $a-b$ changes depending on whether $a>b$ or $a<b$.
In our transformer-based models, Subtraction and Addition have some similar quanta. For example:
  • In addition, the first answer digit is always and “1” or “0” and this token is the hardest to calculate accurately because of the “cascading carry 1” edge case. In subtraction, the first answer token is always “+” or “-” and this token is the hardest to calculate accurately because of the “cascading borrow 1” edge case.
  • If $a>b$ then the algorithm to calculate $a-b$ and $a+b$ have strong structural similarities, allowing per-digit calculations with one sub-task moving data between digits (“carry 1” or “borrow 1”)
  • If $a<b$ then the algorithm to calculate $a-b$ can use the mathematical identity that $1a-b=-(b-a)$. Our intuition is that the model will calculate if $a<b$, and if so will reuse the $a>b$ circuit, then negate the answer.
We created a very-low-loss 6-digit addition model. Figure 8 shows the lowest quanta impacted when ablating each node. We created a “mixed operations” model (6-digit, 3-layer, 4-head) to do both Addition and Subtraction. We initialised the untrained model with the addition model and trained it (with 80% subtraction and 20% addition questions) to be very low loss. This mixed model can predict one million addition and one million subtraction questions without error.
Figure 19: The per-digit log-loss curves for a 6-digit 3-layer 4-head mixed model trained to perform subtraction and addition after being initialised with a very low loss addition model.
To investigate what classes of questions the model is always getting wrong, we expanded our existing addition-specific metrics, to include subtraction-specific question categorisations (see Sec. 4.3. Figures 9 and  10 maps these quanta against nodes showing which are used for addition and subtraction.
The model contains a new sub-task that stands out: The algorithm relies on calculations done at token position P0, when the model has only seen one question token! What information can the model gather from just the first token? Intuitively, if the first token is a “8” or “9” then the first answer token is more likely to be a “+” (and not a “-”). The model uses this heuristic even though this probabilistic information is sometimes incorrect and so will work against the model achieving very low loss.
Figure 20: This map shows the answer digit(s) A0 .. A7 impacted when we ablate each useful node in the 6-digit 3-layer 4-head mixed model. A7 is the sign of the answer and calculated first. Note the similarity to the 5 and 6-addition Figures 7 and  17.

### v7

@@
-
+Increasing Trust in Language Models through the Reuse of Verified Circuits
 Abstract
@@
 Here, we define a stringent standard of trustworthiness whereby the task algorithm and circuit implementation must be verified, accounting for edge cases, with no known failure modes.
-We show that a transformer model can be trained to meet this standard if built using mathematically and logically specified frameworks.
-In this paper, we fully verify a model for n-digit integer addition.
-To exhibit the reusability of verified modules, we insert the trained integer addition model into an untrained model and train the combined model to perform both addition and subtraction.
+We show that a model can be trained to meet this standard if built using mathematically and logically specified frameworks.
+In this paper, we fully verify an auto-regressive transformer model that performs n-digit integer addition.
+To exhibit the reusability of verified modules, we insert the trained integer addition model into a larger untrained model and train the combined model to perform both addition and subtraction.
 We find extensive reuse of the addition circuits for both tasks, easing verification of the more complex subtractor model.
 We discuss how inserting verified task modules into LMs can leverage model reuse to improve verifiability and trustworthiness of language models built using them.
-The reuse of verified circuits reduces the effort to verify more complex composite models which we believe to be a significant step towards safety of language models.
+The reuse of verified circuits reduces the effort to verify more complex composite models which we believe to be a significant step towards safety and interpretability of language models.
+Increasing Trust in Language Models through the Reuse of Verified Circuits
+Philip Quirke Apart Research.
+philipquirkenz@gmail.com                        Clement Neo Apart Research.
+Fazl Barez Apart Research.
+Department of Engineering Sciences, University of Oxford fazl@robots.ox.ac.uk
 1 Introduction
-Figure 1: An overview of our methodology: (1) We trained a perfectly accurate 6-digit integer addition model.
+Figure 1: An overview of our methodology: (1) We trained an accurate 6-digit integer addition model.
 (2) We reverse-engineered the model to find the algorithms that were implemented to perform addition.
-(3) We inserted the addition model into a new model, by copying the weights of the attention heads and MLPs (in brown) into the larger model during initialisation.
+(3) We inserted the addition model into a new model, by copying the weights of the attention heads and MLPs (in brown) into the larger model during initialization.
 (4) We then train the new model on 80% subtraction and 20% addition questions.
-(5) We find that the resulting model reuses the inserted addition circuits for both addition and subtraction questions.
+(5) We find that the resulting model predicts accurately and reuses the inserted addition circuits for both addition and subtraction questions.
 Transformer-based large language models (LLMs) are powerful (Barak et al., 2022) yet largely inscrutable due to their complex, nonlinear interactions in dense layers within high-dimensional spaces.
 Given this complexity, their deployment in critical settings (Zhang et al., 2022) highlights the need for understanding their behavior.
-Many argue that making these models interpretable is key to their safe use (Hendrycks & Mazeika, 2022).
+Hendrycks and Mazeika (2022) argue that making these models interpretable is key to their safe use.
 Mechanistic interpretability focuses on demystifying and validating the algorithms behind model weights, translating complex computations into more human-understandable components (Raukur et al., 2022).
@@
 We aim to achieve this by holding the model to a standard that we term known-good.
-A model performing a task is known-good if:
+We define a model performing a task to be known-good if:
   1. 1.
-    The model’s algorithm for the task and the mechanisms it implements (the “circuits”) are understood.
+    The model’s algorithm for the task and the mechanisms it implements (the “circuits") are understood.
   2. 2.
-    All possible edge cases have been identified and tested. As a heuristic, the model should be able to perform the task one million times without error (“1M Q accuracy”).
+    All possible task edge cases have been identified and tested.
   3. 3.
-    The model has no known failure modes.
-However, exhaustive testing of a model for most tasks is unfeasible.
-For instance, a task involving the addition of 5-digit integers (e.g., 12345+67890) presents ten billion possible variations.
-Rather than testing every variation, the task can be conceptualized within a formal framework to pinpoint edge cases the model must handle.
+    Empirically, the model prediction accuracy is 99.9999% (a standard reliability measure used in industry and abbreviated as “six nines").
+That is, it can perform the task one million times with at most 1 wrong prediction.
+Exhaustive testing of a model task may be infeasible.
+For instance, when adding two 5-digit integers (e.g.
+12345+67890) there are ten billion variations.
+Some tasks can be conceptualized within an existing formal framework that allows identification of all edge cases the model must handle.
 For example, in 5-digit addition, the most uncommon edge case is 55555+44445=100000, which requires a carry bit to cascade through all digits, occurring in only 0.002% of cases.
-A known-good model must incorporate algorithms to manage this and other more frequent edge cases.
-A known-good model would hence have circuits that are verified to perform the task perfectly.
+A known-good model must incorporate algorithms to manage all known edge cases.
+A known-good model would hence have verified [^1] In this paper, ’verified’ has the mechanistic interpretability meaning that a specific group of interconnected neurons within a neural network reliably and causally contributes to the model’s output in a meaningful, understandable way, with supporting empirical evidence.
+circuits that perform the task accurately.
 In this paper, we detail the development and interpretation of a known-good model for addition.
@@
 We confirm the validity of the entire set of circuits, ensuring they cover all identified edge cases.
-The model achieves a very low training loss and can do one million predictions without error, making it a known-good model.
-Additionally, we develop a “mixed” model capable of both addition and subtraction, incorporating the known-good addition model.
-This mixed model extensively reuses the addition circuits for both operations and learns new circuits, facilitating the interpretation of the model’s algorithm.
+The model achieves a very low training loss and has six nines (99.9999%) accuracy.
+The model hence achieves our standard of ’known-good’.
+Additionally, we develop a “mixed" model capable of both addition and subtraction, incorporating the known-good addition model.
+This mixed model has six nines accuracy, and extensively reuses the addition circuits for both operations and learns new circuits, facilitating the interpretation of the model’s algorithm.
 This approach marks progress toward a known-good model for both addition and subtraction.
 Hence, our main contributions are three-fold:
-  • A trustworthy known-good 5-digit addition model, with an algorithm extensible to n-digits.
-  • A 6-digit addition and subtraction model that reuses established addition circuits for both operations, indicating steps toward a known-good model for these tasks.
-  • A proof of concept for enhancing the trustworthiness of a larger model by integrating a smaller known-good transformer model.
+  • Defining several known-good n-digit addition models with six nines accuracy which all use the same algorithm.
+  • Demonstrating a proof of concept for re-using a known good model in the training of another larger, more-capable model, simplifying the interpretation of the more-capable model’s algorithm.
+  • Defining several n-digit addition and subtraction models with six nines accuracy, that reuse established addition model circuits for both operations, and detailing progress towards these models being known-good.
 2 Related Work
 Mechanistic interpretability aims to reverse engineer neural networks to find interpretable algorithms that are implemented in a model’s weights (Olah et al., 2020).
-To better represent these algorithms, Elhage et al.
-(2021a) introduced a mathematical framework to represent how transformer attention heads can work with each other to implement complex algorithms.
-To explain how (not where) a model algorithm is implemented, Jenner et al.
-(2023) recommends documenting a low-level computation graph (causal abstraction), with a mapping from the graph to the model nodes that implement the computation, with experimentation verification.
+Mathematical frameworks (Elhage et al., 2021a) explain how transformer attention heads can work with each other to implement complex algorithms.
+Causal Scrubbing (Jenner et al., 2023) recommends explaining a model algorithm by documenting a low-level computation graph, mapping from the graph to the model nodes that implement the computation, and performing experimentation verification.
 Investigative techniques such as ablation interventions, activation unembeddings (nostalgebraist, 2020), and sparse autoencoders (Nanda, 2023; Cunningham et al., 2023), underpinned by the more theoretical frameworks (Elhage et al., 2021b; Geva et al., 2022), provide tools to help confirm a mapping.
@@
 Even though basic arithmetic can be solved following a few simple rules, pre-trained LMs often struggle to solve simple math questions (Hendrycks et al., 2021).
-Stolfo et al.
-(2023) used causal mediation analysis to investigate how large pre-trained LMs like Pythia and GPT-J performed addition to solve word problems.
-It is also possible to improve a model’s arithmetic abilities through training data enrichment.
-Liu & Low (2023) used supervised fine tuning (including enriched training data) to convert LLaMA 7B to Goat - a model that is “Good at Arithmetic Tasks” including addition and subtraction.
-They also showed the model’s performance relied in part on LLaMA’s consistent tokenization of digits.
+Causal mediation analysis (Stolfo et al., 2023) has been used to investigate how large pre-trained LMs like Pythia and GPT-J performed addition to solve word problems.
+It is also possible to improve a model’s arithmetic abilities used supervised fine tuning - including enriched training data (Liu and Low, 2023).
 Studying Toy Models for Arithmetic.
-Doing mechanistic interpretability on toy transformers can help to better isolate clear, distinct circuits given the highly specific experimental setup for the model studied.
-Nanda et al.
-(2023) studied modular addition in a one-layer transformer (where each number was represented by a distinct token) to better understand training dynamics.
-Quirke & Barez (2024) detailed how a 1-layer, 3-head transformer model performed 5-digit addition questions.
-It showed a rare edge case (e.g.
-“77778+22222=100000” where a “carry 1” cascades through 4 digits) that the model couldn’t handle - showing the importance of understanding and testing all edge cases for trustworthiness.
-Michaud et al.
-(2023) posits that many natural prediction problems decompose into a finite set of knowledge and skills that are “quantized” into discrete chunks (quanta).
+Doing mechanistic interpretability on toy transformers can help to better isolate clear, distinct circuits given the highly specific experimental setup for the model studied (Nanda et al., 2023).
+Quirke and Barez (2024) detailed a 1-layer, 3-head transformer model that performs 5-digit addition, showing it failed on rare edge cases (e.g.
+“77778+22222=100000" where a “carry 1" cascades through 4 digits), highlighting the importance of understanding and testing all edge cases for trustworthiness.
+Many natural prediction problems decompose into a finite set of knowledge and skills that are “quantized" into discrete chunks (quanta) (Michaud et al., 2023).
 Models must learn these quanta to reduce loss.
-They posit that “understanding a network reduces to enumerating its quanta.” Schaeffer et al.
-(2023) provides useful ways to measure quanta in mathematical prediction problems.
+Understanding a network reduces to enumerating its quanta.
+Other studies (Schaeffer et al., 2023) prove useful ways to measure quanta in mathematical prediction problems.
 3 Methodology
+A transformer model may learn an addition algorithm that differs significantly from the traditional human process.
+Here we define an alternative mathematical framework for addition, showing it is equivalent to the traditional process and so is mathematically correct.
+Then we show that our model implements this alternative framework.
 3.1 Mathematical Framework
 Consider the task of adding two $n$-digit numbers together.
-Let us define the first number as $D=\{D_{n-1},D_{n-2},\ldots,D_{0}\}$ and the second number as $D^{\prime}=\{D^{\prime}_{n-1},D^{\prime}_{n-2},\ldots,D^{\prime}_{0}\}$.
-The resulting answer can be represented as $A=\{A_{n},A_{n-1},\ldots,A_{0}\}$.
-For an illustrative example, see Figure 2.
-Figure 2: We refer to individual tokens in a 5-digit addition question as D4, ..
-D0, and D’4, .., D’0 and the answer tokens as A5, .., A0.
+We define the first number as $D=\{D_{n-1},D_{n-2},\ldots,D_{0}\}$ and the second number as $D^{\prime}=\{D^{\prime}_{n-1},D^{\prime}_{n-2},\ldots,D^{\prime}_{0}\}$ and the answer as $A=\{A_{n},A_{n-1},\ldots,A_{0}\}$.
+Figure 2 shows an illustrative example.
+Figure 2: For 5-digit addition, our model has 12 input (question) and 6 output (answer) token positions.
+We name the question tokens D4, …, D0, and D’4, …, D’0 and the answer tokens A5, …, A0.
 For n-digits, we use the terms $D_{n}$, $D^{\prime}_{n}$and $A_{n}$.
-We categorize the sum of each digit pair into three overlapping groups BA, MC and US defined as:
-$BA_{n}=(D_{n}+D^{\prime}_{n})\mod 10$ (1)
-$MC_{n}=\begin{cases}1&\text{if }(D_{n}+D^{\prime}_{n})\geq 10,\\ 0&\text{otherwise}.\end{cases}$ (2)
-$US_{n}=\begin{cases}1&\text{if }(D_{n}+D^{\prime}_{n})=9,\\ 0&\text{otherwise}.\end{cases}$ (3)
-When calculating $(A_{n})$, considering only the carry bit from the two preceding digits, we have:
-$A_{n}=\begin{cases}D_{n}+D^{\prime}_{n}+1\underbrace{\text{mod}}_{10},&\text{if }\underbrace{MC}_{n-1}=1\text{ or}\\ &(\underbrace{US}_{n-1}=1\text{ and }\underbrace{MC}_{n-2}=1),\\ D_{n}+D^{\prime}_{n}\underbrace{\text{mod}}_{10},&\text{otherwise.}\end{cases}$ (4)
-This approach is aligned with the framework implemented by the addition model in Quirke & Barez (2024).
-However, it is restricted to cases where the carry bit cascades over two or fewer digits.
-To address this limitation, we introduce an additional digit-level categorization called TriCase:
-$D_{n}.C=\underbrace{\text{TriCase}}_{(D_{n},D^{\prime}_{n})}=\begin{cases}T10&\text{if }(D_{n}+D^{\prime}_{n})\geq 10,\\ T9&\text{if }(D_{n}+D^{\prime}_{n})=9,\\ T8&\text{if }(D_{n}+D^{\prime}_{n})\leq 8.\end{cases}$ (5)
-These categorization cases contain carry bit information:
-  • If $D_{n}.C=T10$, the digit-pair always causes a MC,
-  • if $D_{n}.C=T9$, the digit-pair will cause an MC if the proceeding digit-pair sum generates an MC, and
-  • if $D_{n}.C=T8$, the digit-pair never causes an MC .
-To compute the first digit of the answer in a problem such as (55555+44445=)(55555+44445=)( 55555 + 44445 = ), the algorithm must synthesize the five TriCase values from the respective digit-pairs.
-This synthesis is crucial to determine the presence of a carry bit cascading through all five digits.
-We introduce a function, TriAdd, to facilitate the transfer of TriCase data between digits. It is defined as follows:
-$\underbrace{\text{TriAdd}}_{(D_{n}.C,D_{n-1}.C)}=\begin{cases}T10&\text{if }D_{n}.C=T10\\ &\text{or }D_{n}.C=T9\\ &\text{\quad and }D_{n-1}.C=T10,\\ T9&\text{if }D_{n}.C=T8\\ &\text{\quad and }D_{n-1}.C=T10,\\ T8&\text{otherwise.}\end{cases}$ (6)
-TriAdd transfers MC data from $(D_{n-1})$ to $(D_{n})$ by integrating the values of $(D_{n}.C)$ and $(D_{n-1}.C)$.
-This function can be represented as nine bigram mappings and yields three distinct outputs (details in App.
-D).
-Finally, we introduce the recursive function $D_{n}$.$C_{m}$ , which aggregates TriCase data across $(m)$ digits to enhance accuracy.
-$D_{n}.C_{m}=\begin{cases}\underbrace{\text{TriAdd}}_{(D_{n}.C,D_{n-1}.C_{m-1})}&\text{if }m>2,\\ \underbrace{\text{TriAdd}}_{(D_{n}.C,D_{n-1}.C)}&\text{if }m=2.\end{cases}$ (7)
-Simply put $D_{n}$.$C_{m}$ cascades the carry bit over $m$ digits.
-Calculating $D_{1}.C_{2}$ means calculating $\text{TriAdd}(D_{1}.C,D_{0}.C)$ combining the $D_{0}$-specific TriCase value and $D_{1}$-specific TriCase value to give a perfectly accurate TriCase value for $D_{1}$.
-Calculating $D_{4}.C_{5}$ means calculating $\text{TriAdd}(D_{4}.C,D_{3}.C_{4})$ combining the $D_{4}$-specific TriCase value with the perfectly accurate $D_{3}.C_{4}$ TriCase value to give a perfectly accurate TriCase value for $D_{4}$.
-In 5-digit addition, the values D4.C5, D3.C4, D2.C3 and D1.C2 are perfectly accurate as they integrate TriCase data all the way back to and including D0.MC .
-An accurate algorithm must calculate these values.
-Works on 5-digit model was unable to calculate these $D_{n}$.$C_{m}$ and so was not perfectly accurate (Quirke & Barez, 2024) .
-As shown later, our model’s algorithm incorporates the (Quirke & Barez, 2024) algorithm, and adds these perfectly accurate $D_{n}$.$C_{m}$ Figure 3 shows the combined algorithm.
-Figure 3: In 5-digit addition the model first calculates all $D_{n}.C$ values, then combines them into perfectly accurate $D_{n}$.$C_{m}$ values by P11 (token =) when A5 is predicted.
-From P12, the remaining answer digits are calculated one by one, using $D_{n}$.$C_{m}$ values if needed.
-This algorithm can be extended to generalise to n-digits.
-3.2 Question Difficulty and Frequency
-To analyse question difficulty, we categorised questions by the complexity of the computation required to solve the question, as shown in Tab.
-1.
-The categories are arranged according to the number of digits that a carry bit has to cascade through.
-Table 1: We categorise addition questions into non-overlapping “calculation complexity” quanta, ordered by increased computational difficulty (and decreasing occurrence frequency).
-Five-digit addition questions quanta are S0 to S5.
-Ten-digit addition question quanta are S0 to S10.
-S10’s frequency is $\sim 3\times 10^{-4}$ showing the need to enrich training data for rare edge cases.
-Name Contains Example Freq
-S0 BA 11111+12345=23456 $\sim$5%
-S1 BA,MC 11111+9=22230 $\sim$21%
-S2 BA,MC,US 11111+89=22300 $\sim$34%
-S3 BA,MC,USx2 11111+889=23000 $\sim$28%
-S4 BA,MC,USx3 11111+8889=30000 $\sim$11%
-S5 BA,MC,USx4 11111+88889=100000 $\sim$2%
+First, the model will need to sum pairs of digits. As in Quirke and Barez (2024), we define the “Simple Addition" sub-task $A_{n}.SA$, which naively calculates the sum of pairs of question digits, as:
+$A_{n}.SA=(D_{n}+D^{\prime}_{n})\mod 10$ (1)
+When there is no carry bit from the previous digit $A_{n}=A_{n}.SA$.
+The model will also need learn how to calculate whether the addition creates a carry bit.
+As in Quirke and Barez (2024), we define the "Simple Carry" sub-task $A_{n}.SC$:
+$A_{n}.SC=\begin{cases}1&\text{if }(D_{n}+D^{\prime}_{n})\geq 10,\\ 0&\text{otherwise}.\end{cases}$ (2)
+The model in Quirke and Barez (2024) is limited to the framework described so far.
+While it is capable of handling simple addition and carry bits generated directly from the addition of digit pairs, it encounters difficulties when faced with problems involving ’cascading carry’ bits, wherein a carry bit from one digit position propagates to the next.
+To illustrate this limitation, consider the problem "00144+00056=000210".
+When adding the digits in the tens position, 4+5, no carry bit is generated directly.
+However, due to the carry bit propagating from the ones position, a carry bit is ultimately required in the tens position.
+A model that solely calculates the sums of digit pairs and considers carry bits resulting from the direct addition of those pairs will fail to accurately resolve this type of problem.
+In such cases, it would yield an incorrect result, such as "144+56=110".
+The 1-layer model could cascade the carry bit up to two digits across, but not more.
+3.2 Extending the Mathematical Framework
+To answer "44444+55556=" with "100000", an accurate model must predict the first answer digit A5 as "1".
+To do so, an accurate model must implement a “carry one cascade" circuit, which combines the “carry one" information from all five digits.
+This is especially hard as the model predicts answer tokens from left to right.
+We introduce a digit-level sub-task called TriCase that calculates the essential “carry one" information for a single pair of digits $D_{n}$ and $D^{\prime}_{n}$.
+TriCase has 3 possible outputs representing a definite carry one (ST10), a possible carry one depending on the results of other calculations (ST9), or definitely not a carry one (ST8):
+$\begin{split}A_{n}.ST&=\underbrace{\text{TriCase}}_{(D_{n},D^{\prime}_{n})}=\\ &\begin{cases}ST10&\text{if }(D_{n}+D^{\prime}_{n})\geq 10,\\ ST9&\text{if }(D_{n}+D^{\prime}_{n})=9,\\ ST8&\text{if }(D_{n}+D^{\prime}_{n})\leq 8,\\ \end{cases}\end{split}$ (3)
+To perform the “cascading carry one" calculation, we introduce a TriAdd sub-task.
+It handles the case where a possible carry one becomes a definite carry one because the next lower digit pair generated a carry one.
+TriAdd is defined as:
+$\begin{split}A_{n}.SV&=\underbrace{\text{TriAdd}}_{(A_{n}.ST,A_{n-1}.ST)}=\\ &\begin{cases}ST10&\text{if }A_{n}.ST=ST10\text{ or }\\ &\quad(A_{n}.ST=ST9\text{ and }\\ &\quad A_{n-1}.ST=ST10),\\ ST8&\text{otherwise.}\end{cases}\end{split}$ (4)
+The model can use $ST$ and $SV$ to accurately calculate A5 as 1 or 0 by using A5 = A4.SV = TriAdd(A4.ST, TriAdd(A3.ST, TriAdd(A2.ST, TriAdd(A1.ST, A0.ST)))).
+Note that in calculating A5, the model has also calculated an accurate carry bit for each answer digit.
+For example, the carry bit for A4 is A3.SV = TriAdd(A3.ST, TriAdd(A2.ST, TriAdd(A1.ST, A0.ST))).
+With this framework, the model only needs the sub-tasks $A_{n}$.SA, $A_{n}.ST$ and $A_{n}.SV$ to accurately perform addition (but some models also use the redundant $A_{n}$.SC sub-task).
+This framework, if implemented by a model, is sufficient for the model to perform n-digit addition accurately.
+Figure 3 diagrams how our model’s algorithm uses SA, SC, $ST$ and $SV$ to perform addition.
+This algorithm is like the 99% accurate Quirke and Barez (2024) algorithm but contains an additional circuit (the shaded boxes) to calculate “cascading carry one" data to predict with 99.9999% accuracy.
+Figure 3: To predict with 99.9999% accuracy, the addition algorithm first calculates “carry one" values $(A_{n}.ST)$, combining them into “cascading carry one" values $(A_{n}.SV).$ At the “+" token, $A_{n-1}.SV$ gives the first answer digit as 1 or 0.
+The other answer digits are calculated by combining “base add" $(A_{n}.BA)$ and “carry one" $(A_{n}.SC)$ calculations with the pre-calculated $A_{n}.SV$ values.
 3.3 Techniques
-In investigating the circuits of the model, we want to understand what each attention head or MLP layer is doing across each token position, and how it relates to our mathematical framework.
+In investigating the model circuits, we want to understand what each attention head or MLP layer is doing across each token position, and how it relates to our mathematical framework.
 Hence, we define a node as the computation done by an attention head or MLP layer for a given token position.
-Mean ablation. To find out how the model depends on the output of a node, we replace the output of that node with the vector that is the mean of all of its outputs across its batch.
-Attention Patterns. To find out what the model attends to at that node, we take the attention pattern at that token position and take the significant tokens attended to ($>0.01$ post softmax).
+To investigate what each node is doing, we use the following techniques:
+  1. 1.
+    Intervention Ablation.
+To find out how the model depends on the output of a node, we replace the output of that node with the vector that is the mean of all of its outputs across a batch and measure how that impacts downstream performance.
+We also use (automated, n-digit) intervention ablation tests, specific to each sub-task, to test for the expected sub-task behavior.
+  2. 2.
+    Attention Patterns. To find out what the model attends to at a node, we take the attention pattern at that token position and take the significant tokens attended to ($>0.01$ post softmax).
+  3. 3.
+    Principal Component Analysis (PCA). We use PCA to investigate the outputs of attention heads, especially where our framework suggests the head output may be tri-state or bi-state.
+  4. 4.
+    Question Complexity.
+We categorized questions by the complexity of the computation required to solve the question (Refer App.
+G).
+The addition categories (named S0 to S4) reflect the number of sequential digits that a “carry one" cascades through.
+The subtraction categories (named M0 to M4) reflect the number of sequential digits that a “borrow one" cascades through.
+We analysed which nodes were needed for correct prediction of each category.
 4 Experiments
-4.1 Training & Investigating Five-Digit Addition Model
-Training the model.
-The 5-digit 1-layer addition model in Quirke & Barez (2024) achieved an accuracy of approximately 99%.
-Our preliminary experiments indicated that a 2-layer, 3-head model was the smallest configuration capable of achieving fully accurate addition (see App.
-C for alternative configurations tested).
-This configuration effectively doubled the computational power compared to the 1-layer model.
-Moreover, a 2-layer model introduces the capability to “compose” the attention heads in novel ways, facilitating the implementation of more complex algorithms (Elhage et al., 2021b).
-By enriching the training dataset with more instances of rare edge cases and using 30 thousand training batches, the 5-digit 2-layer model attained a final training loss of approximately $2.3\times 10^{-8}$.
-This model demonstrated a 1M Q accuracy (details in App.
-B).
-Investigating the model.
-To gain insights into the model’s predictive behavior, we conducted an investigation focusing on its activity during prediction calculations.
-Ablation experiments targeting the nodes revealed that the model primarily depends on nodes located in token positions D’2 to A1, as illustrated in Figs 4 and 5.
-Across these nine positions, the model engages 36 nodes, comprising 21 attention head nodes and 15 MLP layer nodes.
-The effects of node ablation on our complexity and answer-impact metrics were analyzed (see Figs 6 and 7), elucidating the specific computations performed at each node.
-For each answer digit $A_{n}$using test questions corresponding to the T8, T9 and T10 categories, we performed Principal Component Analysis (PCA) on the nodes yielding interpretable results.
-Specifically nine “node and answer-digit” combinations (see Figs 13 and 14) showed strong clustering of the questions aligned to the T8, T9 and T10 categories, suggesting these nodes are implementing the TriCase categorisation.
-Figure 4: This map shows the % of enriched questions that fail when we ablate each node in the 5-digit 2-layer 3-head addition model.
-The model only uses nodes in token positions P8 to P16 (tokens D’2 to A1).
-No attention heads in layer 1 are used.
-Lower percentages correspond to rarer edge cases.
-The grey space represents nodes that are not used by the model.
-Figure 5: This map shows the input tokens each attention head attends to at each token position.
-At token position P11 (token =) the model predicts the first answer digit A5.
-All digit pairs (e.g.
-D2 D’2) are attended to by P11.
-Figure 6: This map shows the lowest complexity quanta S0, S1, etc impacted when we ablate each node in the 5-digit 2-layer 3-head addition model.
-To answer S0 questions, only the S0 nodes are used.
-To answer S1 questions, S0 and S1 nodes are used, etc.
-Figure 7: This map shows the answer digit(s) A0 ..
-A5 impacted when we ablate each useful node in the 5-digit 2-layer 3-head addition model.
-The “column per digit” pattern in the last 6 columns mirrors the 1-layer model’s pattern.
-Cells containing values like A2..5 (in column D’0) are used in multiple prediction steps to calculate multiple answer digits e.g.
-A2 to A5.
-The algorithm predicts the first answer digit, A5, at position P11 (tokne =).
-This digit, which can only be 0 or 1, is the most challenging to predict as it may rely on a cascade starting from the digit pair D0 D’0 (e.g., 55555+44445=100000).
-The algorithm is required to compute this cascade using the nodes located in positions P8 to P11 (input tokens D’2 to =).
-As illustrated in Figure 5, these nodes attend to all digit pairs from D4 D’4 to D0 D’0 by P11.
-Additionally, the PCA data, as shown in Figures 13 and 14, suggest that these nodes produce tri-state outputs.
-After encountering several initial setbacks (see App. E and F), we discovered that the model utilizes a minimal set of “make carry” information, leading to the development of the TriCase quanta.
-The model performs $D_{n}.C$ using bigrams (see App.
-D) to map two input tokens to one result token e.g.
-“6” + “7” = T10.
-In positions P8 to P11, the model does $D_{n}.C$ calculations on all digit pairs from D4 D’4 to D0 D’0.
-The mathematical framework behind this is intricate. To validate it, we implemented the algorithm in Python and successfully executed one million additions with zero errors.
-For a specific instance of the 5-digit addition model, we mapped the algorithm to individual nodes.
-Using ablation interventions, we verified each node’s role in executing the expected algorithmic step (see App.
-G).
-Table 2 and Figure 1 illustrate how the algorithm was applied to this particular model instance, adhering to all known constraints and accurately calculating all six answer digits.
-Table 2: Pseudo-code for accurately predicting all answer digits, with mappings to the model nodes that implement the pseudo-code, for a specific 5-digit model instance.
-Value Value calculated as Nodes used
-D2.C
-TriCase(D2, D’2)
-P8.L0.H1 and P8.L0.MLP
-D1.C
-TriCase(D1, D’1)
-P9.L0.H1 and P9.L0.MLP
-D1.C2
-TriAdd(D1.C, TriCase(D0, D’0))
-P10.L0.H1 and P10.L0.MLP
-D3.C4
-TriAdd(TriCase(D3, D’3), TriAdd(D2.C,D1.C2))
-P11.L0.H1
-D4.C
-TriCase(D4, D’4)
-P11.L0.H2
-D4.C5
-TriAdd(D4.C, D3.C4)
-P11.L0.MLP
-A5
-(D4.C5 == 10)
-P11.L1.MLP
-D4.BA
-(D4 + D’4) % 10
-P12.L0.H0 + H2
-D3.C4
-TriAdd(TriCase(D3, D’3), TriAdd(D2.C,D1.C2))
-P12.L0.H1
-A4
-(D4.BA + (D3.C4 / 10)) % 10
-P12.L0.MLP and P12.L1.MLP
-D3.BA
-(D3 + D’3) % 10
-P13.L0.H0 + H2
-D2.C3
-TriAdd(D2.C,D1.C2)
-P13.L0.H1
-A3
-(D3.BA + (D2.C3 / 10)) % 10
-P13.L0.MLP and P13.L1.MLP
-D2.BA
-(D2 + D’2) % 10
-P14.L0.H0 + H2
-D1.C2
-(D1 + D’1) / 10 + P10.D1.C2 *
-P14.L0.H1
-A2
-(D2.BA + (D1.C2 / 10)) % 10
-P14.L0.MLP and P14.L1.MLP
-D1.BA
-(D1 + D’1) % 10
-P15.L0.H0 + H2
-D0.MC
-(D0 + D’0) / 10
-P15.L0.H1
-A1
-(D1.BA + D0.MC) % 10
-P15.L0.MLP and P15.L1.MLP
-A0
-(D0 + D’0) % 10
-P16.L0.H0 + H2 P16.L0.MLP and P16.L1.MLP
-Throughout the training of different instances of the 5-digit models, small modifications to the training environment resulted in subtly varied implementations.
-All variations seen use the Figure 3 algorithm, but each has small implementation differences.
-For example, in positions P8 to P11, the model might alter which 5 nodes perform the $D_{n}.C$ calculations and in what order.
-This corresponds to small changes in the Tab.
-2 documentation per variation.
-In summary, we concluded that the specific instance of the 5-digit addition model documented above is reliably accurate, well-understood, well-functioning and so known-good.
-4.2 Case Study: Extending to Six-Digit Addition
-Our investigations showed that all instances of the 5-digit addition model we trained appeared to utilize the same algorithm.
-To determine if this algorithm extends to related models, we trained a 6-digit 2-layer 3-head addition model, employing a different seed and optimizer.
-Additionally, we modified the answer format to include a ’+’ token (e.g., 111111+222222=+0333333).
-Employing similar quanta and training data enrichment techniques, we developed a 6-digit addition model. This model achieved a loss of approximately $9.1\times 10^{-9}$ and demonstrated 1M Q accuracy.
-The 5-digit (see Figures 4,  5,  6, and  7) and 6-digit (see Figures 15,  16,  8, and  17) quanta maps show numerous similarities, suggesting the 6-digit model implements the logical extension of the 5-digit algorithm (see Figure 3).
-Intervention ablation tests show the 6-digit model does 6 BA and 6 MC calculations in late tokens, and does 6 $D_{n}.C$ calculations in early tokens with PCA showing the expected TriCase output.
-The 5 $D_{n}$.$C_{m}$ calculations are still to be tested.
-Give this model has 1M Q accuracy, if this testing is successfully completed, and the algorithm implementation detail is documented, we can state the 6-digit model is known-good.
-Figure 8: This map shows the simpliest addition complexity quanta S0, S1, etc used in each useful node of the 6-digit 2-layer 3-head addition model. Note the similarity to 5-digit Figure 6
-4.3 Case Study: Extending the model for subtraction
-To explore the reuse of known-good models, we developed a larger (6-digit 3-layer 4-head) model capable of performing both addition and subtraction.
-We introduced specific subtraction quanta and enriched the training dataset with rare subtraction edge cases.
-This new “mixed” model, initially untrained, was initialized with the weights from the (6-digit 2-layer 3-head) addition model.
-To facilitate its learning of subtraction while maintaining its pre-inserted addition skills, we trained the mixed model with a dataset comprising 80% subtraction and 20% addition problems.
-This approach resulted in a model achieving 1M Q accuracy for both addition and subtraction tasks.
-Attempts to “freeze” useful attention heads and/or MLP layers from the addition model by periodically copying their weights back into the mixed model every 100 training epochs, however, led to decreased accuracy.
-Subtraction ($a-b$) presents a greater challenge than addition ($a+b$), primarily due to its two distinct cases: $a>b$ (positive answer) and $a<b$ (negative answer).
-The model must discern this difference to accurately determine the first answer token as either “+” or “-”.
-In cases of “positive” subtraction, the process closely mirrors addition. This observation led us to define subtraction quanta (see Tables 3 and 4 (see App. J), paralleling the addition quanta.
-Calculating the “-” case in subtraction is more complex than the “+” case.
-The model might employ distinct circuits for each scenario.
-Our hypothesis is that after mastering the simpler “+” case, the model will learn to apply the mathematical identity $a-b=-(b-a)$, thus enabling it to handle “-” case questions by re-utilizing the “+” case circuits and then inverting the result.
-To examine this, we introduced the NG quanta (see Table 4) and again applied mean-ablation techniques to analyze the process.
-Table 3: We categorise digit-pair differences (i.e.
-$D_{n}$- $D^{\prime}_{n}$ abbreviated as $D_{n}$.Diff) into three groups.
-These subtraction groups parallel the BA, MC and US addition groups.
-Groups BS and DZ overlap.
-Name Abbrev Definition E.g. Like
-Base Subtract BS $D_{n}$.Diff $<=0$ 7 - 2 BA
-Borrow 1 B1 $D_{n}$.Diff $<0$ 7 - 9 MC
-Diff Zero DZ $D_{n}$.Diff = 0 7 - 7 US
-Table 4: We categorise subtraction questions into non-overlapping “calculation complexity” quanta, ordered by increased computational difficulty and decreasing occurrence frequency.
-Five-digit subtraction questions quanta are M0 to M4.
-Ten-digit subtraction question quanta are M0 to M9, and M9’s frequency is $\sim$0.000001%.
-Subtraction questions where the answer is negative get their own category NG .
-Name Contains Example Freq
-M0 BS 33333-22222=+11111 $\sim$25%
-M1 BS,B1 33333-22282=+11051 $\sim$15%
-M2 BS,B1,DZ 33333-22382=+10951 $\sim$8%
-M3 BS,B1,DZx2 33333-23382=+09951 $\sim$2%
-M4 BS,B1,DZx3 33333-23338=+09995 $\sim$0.1%
-NG N/A 10000-90000=-80000 $\sim$50%
-We compared the nodes utilized by the inserted addition model (see Figure 6) to those used by the mixed model when performing addition (see Figure 9).
-There is a strong correlation between them.
-Notably, the mixed model predominantly uses the existing attention heads in layers 1 and 2, ignoring many available spare heads.
-Further comparison was made between the nodes the mixed model employs for addition and subtraction (see Figures 9 and 10).
-The nodes used for addition (S*) and those for “subtraction with a positive answer” (M*) show a very strong correlation.
-This suggests that the model repurposes its addition nodes to facilitate subtraction in some manner.
-Notably, an additional 15 nodes are activated for subtraction problems that result in a negative answer (represented by the NG quanta).
-While the mixed model has 1M Q accuracy for both addition and subtraction tasks, intervation ablation testing and PCA of the mixed model have not been completed, and the algorithm implementation has not been documented.
-We cannot yet assert that it is known-good.
-Figure 9: This map shows the simplest complexity quanta S0, S1, etc used in each useful node of the 6-digit 3-layer 4-head mixed model when doing addition questions.
-Figure 10: This map shows the simpliest complexity quanta M0, M1, etc used in each useful node of the 6-digit 3-layer 4-head mixed model for subtraction questions with positive answers.
-The NG cells are only used in subtractions $a-b$ where $a<b$, likely implementing the $-(b-a)$ identity as described in  4.3.
+4.1 Training a Five-Digit Addition Model
+The Quirke and Barez (2024) 5-digit 1-layer addition model achieved an accuracy of $\sim$ 99%.
+Our experiments suggested that a 2-layer, 3-head model was the smallest configuration capable of achieving 99.9999% accuracy (see App.
+D for alternatives tested).
+This configuration effectively doubled the computational power compared to the 1-layer model (see App.
+B for mode model configuration details).
+Moreover, a 2-layer model introduces the capability to “compose" the attention heads in novel ways, facilitating the implementation of more complex algorithms (Elhage et al., 2021b).
+We trained a 5-digit, 2-layer, 3-head model, with a 14 token vocabulary ( 0, .., 9, +, -, =, *, / ), batch size of 64, learning rate of 0.00008 and weight decay of 0.1.
+Training used an infinite dataset enriched with rare edge cases.
+Loss was defined as the mean across all answer tokens of their negative log likelihood loss.
+After 30 thousand training batches, the model’s final training loss was $\sim 2.3\times 10^{-8}$.
+Testing showed this model has six nines accuracy.
+(More details in App.
+C and Tab.
+6).
+4.2 Investigating Five-Digit Addition
+Ablation experiments targeting the nodes revealed that the model depends only on nodes located in nine token positions (Figs 6 and 7).
+Further ablation experiments show that for these nine token positions, model uses 36 nodes in predictions.
+The effects of node ablation on our complexity and answer-impact metrics were analyzed (see Figs 5 and Table  1), providing insight into the specific computations performed at each node.
+For each answer digit $A_{n}$, using test questions corresponding to the ST8, ST9 and ST10 categories, we performed PCA on the nodes yielding interpretable results.
+Specifically nine “node and answer-digit” combinations (see Figs 8 and  9) showed strong clustering of the questions aligned to the ST8, ST9 and ST10 categories.
+(P6) (P9) (P10) (P11) (P12) (P13) (P14) (P15) (P16) (P17)
+D’4 D’1 D’0 = + A5 A4 A3 A2 A1
+L0H0 A4 A2 A1
+L0H1 A5 A5..3 A5 A3 A0
+L0H2 A5 A5..1
+L0MLP A5..2 A4 A3 A2 A1 A0
+L1H0
+L1H2 A4 A3 A2
+L1MLP A1 A0
+Table 1: The nodes used in predictions across token positions.
+For all addition and mixed models studied, before the "=" token, each node often calculates data used to predict multiple answer digits.
+After the "=" token, all nodes in a given token position are used to predict a single answer digit.
+For this sample model, the attention heads in token position P10 labelled A5..3 help predict the answer digits A3, A4 and A5.
+The algorithm predicts the first answer digit, A5, at position P11.
+A5, which is always 0 or 1, is the most challenging to predict as it may rely on a long carry one cascade (e.g.
+55555+44445=100000).
+An accurate algorithm must compute this cascade using the nodes located in positions P8 to P11.
+As illustrated in Figure 7, these nodes attend to all digit pairs from D4 D’4 to D0 D’0.
+Additionally, the PCA data, as shown in Figures 8 and 9, suggest that these nodes produce tri-state outputs.
+After our first two algorithm hypothesises failed testing (see App.
+H and I), we discovered that the model utilizes a minimal set of “carry one” information, leading to the development of the TriCase quanta.
+The model performs $A_{n}.ST$ using bigrams (see App.
+F) to map two input tokens to one result token e.g.
+“6” + “7” = ST10.
+In positions P8 to P11, the model does $A_{n}.ST$ calculations on all digit pairs from D4 D’4 to D0 D’0.
+An MLP layer can be thought of as a “key-value pair" memory (Meng et al., 2022; Geva et al., 2021) that can hold many bigrams and trigrams.
+We posit our MLP implements the TriAdd function using bigrams and trigrams to calculate $A_{n}.SV$ values from $A_{n}.ST$ values.
+For a specific 5-digit addition model instance, we mapped the algorithm to individual nodes.
+Using ablation intervention to override each node, we verified each node’s role in executing the expected algorithmic step.
+Figure 3 diagrams the addition algorithm, while App.
+J details each nodes purpose for this model instance.
+The model adheres to all known constraints.
+Empirically, it achieves six nines accuracy.
+In summary, we concluded that this specific instance of the 5-digit addition model documented above has six nines accuracy, is well-understood and well-functioning and hence known-good.
+The 1-layer model uses 21 nodes to achieve two nines (99%) accuracy. This model uses 36 nodes (an increase of 71%) to achieve six nines accuracy.
+4.3 Training n-digit Addition models
+To investigate whether this algorithm is used widely, we first trained seven 2-layer addition models with 5-, 6- and 10-digits, using different seeds, a different optimizer, and changing the answer format to include a sign token (e.g.
+111111+222222=+0333333).
+These seven models all have very low loss (e.g. 1.5e-8) and six nines accuracy. (Details in Tab.6.)
+4.4 Investigating n-digit Addition models
+We implemented a declarative way to describe each sub-task.
+For example, a sub-task can be described as primarily attending to question digits D2 and D’2, influencing answer digit A3, impacting S0 but not S1 complexity questions, with a specific PCA result, and a specific ablation test.
+Using the declaration, we searched for and identified nodes performing all sub-tasks in all the models.
+An algorithm hypothesis (e.g the one described in section 4.2) is then described in terms of the sub-tasks that must exist in the model and their inter-relationships.
+For example, our addition algorithm description states (in part) that the model must implement $A_{n}.ST$ sub-tasks for each question digit before the “=” token.
+We implemented an n-digit algorithm hypothesis declaration framework and a way to test the declaration against the sub-tasks found in each model and map the results.
+(Table 2 shows an example map.)
+(P11) (P12) (P13) (P14) (P15) (P16) (P17) (P18) (P19) (P20)
+D’1 D’0 = + A6 A5 A4 A3 A2 A1
+L0H0 A2.ST A3.ST A1.ST A4.ST A4.SC A3.SC A2.SC A1.SC A0.SC
+L0H1 A1.ST A0.ST A5.SA A4.SA A3.SA A2.SA A1.SA A0.SA
+L0H2 A5.ST
+Table 2: All addition models studied implement our addition algorithm.
+The algorithm SA, SC and $ST$ sub-tasks all exist for each digit and in appropriate token positions.
+For a sample model, this map shows the subtask locations.
+Interestingly, here each SA sub-task is shared across two attention heads.
+We found that our seven 2-layer addition models all implement our addition algorithm.
+Given these seven models have six nines accuracy, and all implement the same valid algorithm, we can state these models are known-good.
+4.5 Training n-digit Mixed models
+To explore reuse, we initialized untrained models with a known-good addition-only model, then trained them to perform both subtraction and addition. We call these “mixed" models.
+Specifically, we trained seven larger (6- or 10-digit, 2- or 3-layer, 3- or 4-head) models after initializing them with the weights from a known-good 2-layer 3-head addition model.
+The first 2 layers, first 3 heads of the mixed model were initialized with the addition model weights.
+We trained the mixed model with 80% subtraction and 20% addition batches.
+We enriched the (infinite) training dataset with rare addition and subtraction edge cases.
+Some models achieved six nines accuracy and the others five nines.
+Attempts to “freeze" the inserted attention heads and/or MLP layers by periodically copying the addition weights back into the mixed model every 100 training steps resulted in lower accuracy.
+(Refer App.M and Tab.6)
+Figure 4: Each mixed model was initialized with the weights from a known-good addition model, then was trained on 80% subtraction and 20% addition batches.
+A sample log loss graph (final loss 8.0e-9) is shown.
+4.6 Investigating n-digit Mixed models
+Unlike addition, subtraction question answers can be either positive or negative.
+Similar looking positive-answer (e.g.
+10009-10000=+000009) and negative-answer (e.g.
+10009-20000=-009991) subtraction questions can give answers that differ at several digit positions.
+We posited that the model treats three distinct question classes “addition", “positive-answer subtraction" and “negative-answer subtraction" differently.
+Ablation of nodes showed that some nodes only help predict one question class, some help predict two classes and some all three classes.
+It also showed that the inserted addition nodes are heavily used and the majority become polysemantic, performing both addition and subtraction calculations.
+(Details in Tab.3.)
+Used Inserted
+Question class No % No %
+All questions 96 48
+Addition 61 64% 42 88%
+Positive-answer sub 70 73% 40 83%
+Negative-answer sub 53 55% 29 60%
+Table 3: Mixed models re-use most inserted addition-model nodes.
+Many inserted nodes become polysemantic during training.
+That is they perform addition (S), positive-answer subtraction (M) and negative-answer subtraction (N) sub-tasks simultaneously.
+For a sample mixed model that uses 96 nodes and had 48 nodes inserted, this table shows inserted node reuse.
+To investigate the subtraction algorithm, and paralleling addition algorithm subtasks, we defined subtraction-specific sub-tasks Base Diff and Borrow One (see Tab.
+4) and complexity measures (see App.
+G and  N).
+Name +ve Sub -ve Sub Definition
+Base Diff MD ND $D_{n}$- $D^{\prime}_{n}$10
+Borrow One MB NB $D_{n}$- $D^{\prime}_{n}$$<0$
+Table 4: We define 2 “positive-answer subtraction" and 2 “negative-answer subtraction" sub-tasks that parallel the addition sub-tasks Base Add SA and Carry One SC
+We found that the inserted nodes that performed SA in the addition model, perform SA, MD and ND in the mixed model.
+(Refer Table 5.) These three sub-tasks are similar in that each performs a mapping from 100 input cases (10 $D_{n}$x 10 $D^{\prime}_{n}$) to 10 output cases (0..9).
+The mixed model “upgraded" these nodes to be polysemantic during training.
+Similarly, some SC addition nodes became polysemantic and now process SC, MB and/or NB .
+(P9) (P10) (P11) (P12) (P13) (P14) (P15) (P16) (P17) (P18) (P19) (P20)
+D’3 D’2 D’1 D’0 = A7 A6 A5 A4 A3 A2 A1
+L0H0 A4.MT A3.MT A4.ST A2.SC A1.SC A0.SC OPR
+A4.GT A3.MT A4.MT A4.SC A3.SC A2.NB A1.MB A0.MB SGN
+OPR A1.NB A0.NB
+L0H1 A2.ST A1.ST A3.ST A0.ST A5.SA
+A4.ST A2.MT A1.MT A3.GT A0.MT A5.MD
+A2.GT A1.GT A0.GT A4.SA A3.SA A2.SA A1.SA A0.SA
+L0H2 A5.ST A5.SA A4.MD A3.MD A2.MD A1.MD A0.MD
+OPR A5.MD A4.ND A3.ND A2.ND A1.ND A0.ND
+SGN A5.ND
+L0H3 OPR OPR OPR OPR OPR OPR
+SGN SGN SGN SGN SGN SGN
+Table 5: For mixed models, in later tokens, polysemantic attention heads simultaneously generate outputs for the three question classes addition S, M and N.
+Other heads calculate the question class by attending to the question operation (OPR) token and the answer sign (SGN) token.
+The MLP layers then select the output appropriate for the class.
+In this sample map, from P16, the first 3 rows contain many polysemantic nodes, while the 4th row calculates the question class.
+An accurate subtraction model must answer “cascading borrow one" questions like 100000-000001=+0099999.
+We define the “essence of borrow one" subtask $MT$ used by both positive-answer (M) and negative answer (N) subtraction questions.
+$MT$ is like addition’s $ST$ subtask:
+$\begin{split}A_{n}.MT&=\underbrace{\text{TriCase}}_{(D_{n},D^{\prime}_{n})}=\\ &\begin{cases}MTN&\text{if }D_{n}<D^{\prime}_{n}\\ MT0&\text{if }D_{n}=D^{\prime}_{n}\\ MT1&\text{if }D_{n}>D^{\prime}_{n}\\ \end{cases}\end{split}$ (5)
+For M questions, MTN is a definite borrow one, MT0 is a possible borrow one (depending on the results of other digit calculations) and MT1 is definitely not a borrow one.
+For N questions, the interpretation is the opposite.
+Paralleling addition’s “cascading carry one" $A_{n}.SV$ calculation, we define “cascading borrow one" calculation sub-tasks $A_{n}.MV$ for positive-answer subtraction and $A_{n}.NV$ for negative-answer subtraction.
+We posit the MLP implements TriAdd-like functions using bigrams and trigrams to calculate $A_{n}.MV$ and $A_{n}.NV$ values from $A_{n}.MT$ values.
+4.7 Mixed model question class detection
+We posit that there is a specific circuit to detect whether a question is in the S, M or N class.
+If the question operator is “+" then the class is S, but if the question operator is “-" then the model must calculate if $D>=D^{\prime}$ to distinguish between the M and N classes.
+For accuracy, it needs this answer by the “=" token to predict the first answer token (the answer sign) as “+" or “-".
+$D>=D^{\prime}$ can be derived from the $A_{n}.MT$ data. Alternatively, this calculation could be a distinct circuit using a new sub-task we define as:
+$A_{n}.GT=\begin{cases}1&\text{if }D_{n}\geq D^{\prime}_{n},\\ 0&\text{otherwise}.\end{cases}$ (6)
+For a 4 digit question, $D>=D^{\prime}$ can be calculated as A3.GT=1 or (A3.GT=0 and (A2.GT=1 or (A2.GT=0 and (A1.GT=1 or (A1.GT=0 and A0.GT=1))).
+Our test for GT is that ablation causes the answer to change sign.
+The $MT$ calculation is very similar.
+We found $A_{n}.MT$ and $A_{n}.GT$ subtasks in all mixed models.
+Usually (but not always) both subtasks are calculated by the same node.
+(Table 2 shows an example.) As both the $MT$ and GT approaches are valid, models can learn valid algorithm sub-task implementations that differ per answer digit.
+We further posit that a node that implements say A2.SA, A2.MD and A2.ND does not know whether it is dealing with a S, M or N question, so it outputs all three possible answers to the residual stream.
+Another node must calculate the S, M or N distinction likely by attending to the question operator (OPR) and the answer sign (SGN).
+We define attention sub-tasks OPR and SGN for these calculations.
+Table 2 shows that for each answer digit one attention head attends to both OPR and SGN.
+We believe these heads transfer sufficient data to the residual stream to allow the MLP layer(s) to calculate the question class, and so select the appropriate output from the polysemantic A2.SA/MD/ND node.
+4.8 Mixed model summary
+The 7 mixed experimental models all have five or six nines accuracy.
+They all contain the same sub-tasks - suggesting they all implement the same algorithm.
+As we don’t fully understand the algorithm yet, we can’t assert that these mixed models are known-good.
 5 Conclusion
-In this paper, we successfully trained a 5-digit 2-layer 3-head addition model that we have verified as known-good.
-Our investigation revealed that a 6-digit 2-layer 3-head addition model follows the same algorithmic principles, albeit with some implementation variations.
+In this paper, we successfully trained 5-, 6- and 10-digit, 2-layer, 3-head addition models that we have verified as known-good.
+They all implement the same algorithm, albeit with some implementation variations.
 A significant aspect of our work was demonstrating component reuse.
-We achieved this by integrating an existing addition model into a new “mixed” model designed for both addition and subtraction tasks.
-This integration allowed us to leverage our in-depth understanding of the addition model’s algorithm to unravel the workings of the mixed model.
-Our findings indicate that the mixed model employs the nodes of the inserted addition model for conducting both addition and subtraction computations.
-This paper lends support to the assertion by Michaud et al.
-(2023) that many natural prediction problems can be broken down into a finite set of “quanta” computations.
-These quanta are essential for models to learn in order to minimize loss.
+We achieved this by integrating an existing addition model into a new larger “mixed” model designed for both addition and subtraction tasks.
+This mixed model has six nines accuracy.
+This integration allowed us to leverage our in-depth understanding of the addition model’s algorithm to unravel the workings of the mixed model’s algorithm.
+Interestingly, the mixed model re-uses most of the inserted addition model nodes, and “upgrades” most of them to become polysemantic - performing the inserted addition subtask and the corresponding subtraction subtasks simultaneously.
+This paper lends support to the assertion by (Michaud et al., 2023) that many natural prediction problems can be broken down into a finite set of “quanta” computations.
+These quanta are essential for models to learn to minimize loss.
 Furthermore, it aligns with the concept that comprehending a network’s functionality is fundamentally about identifying and understanding its sub-quanta.
@@
 Our research opens up several promising avenues for future exploration.
-A key objective is to establish the mixed model as known-good by reverse-engineering its algorithm implementation and validating it through experimental methods.
-Additionally, we aim to develop a comprehensive model that proficiently handles addition, subtraction, multiplication, and division, ensuring its status as known-good.
-Currently, our approach requires significant human effort, which may not be feasible for large-scale applications.
-However, we believe applying this method to critical tasks, such as logical reasoning involving both simple (e.g., AND, OR) and more complex operators (e.g., Converse, Inverse), is justified.
-This could lead to the creation of more known-good transformer models, potentially extending to frameworks for ethical decision-making.
-A promising finding of our research was the successful insertion of a known-good model into a larger, untrained model while retaining its functionality.
-We think it might be feasible to embed a compact known-good model into an existing Large Language Model (LLM) to enhance its trustworthiness.
-Previous studies (Stolfo et al., 2023; Kruthoff, 2024) found nodes in LLMs that were responsible for inaccuracies mathematical calculations.
-Given our findings, it might be possible to replace these nodes with known-good ones.
-While this poses an engineering challenge, research by (Voita et al., 2023) and (Hu et al., 2021) indicates the presence of “spare” neurons in LLMs and the potential for integrating small-scale modifications to fix erroneous circuits in LLMs.
+A key objective is to complete the reverse-engineering of the mixed model’s algorithm.
+We need to understand how the polysemantic nodes and the OPR+SGN node outputs are represented in the residual stream and show the MLP layers combine this information to select the appropriate question class data to accurately predict the answer digit.
+Additionally, we aim to develop a comprehensive model that proficiently handles integer n-digit addition, subtraction, and multiplication, ensuring its status as known-good.
+We believe this approach can be re-used to create known-good models in fields like logical reasoning (for AND, OR, Converse, Inverse etc tasks) and planning.
+Using a smaller model to improve an existing Large Language Model (LLM) trustworthiness and capabilities is an active area of research.
+Approaches include composing multiple models (Bansal et al., 2024), LM up-scaling which essentially emulates fine-tuning a large pre-trained model using a small model (Mitchell et al., 2023), and this paper’s insertion of an accurate model into an untrained model.
+Other research (Voita et al., 2023; Hu et al., 2021) indicates the presence of “spare” neurons in LLMs and the potential for integrating small-scale modifications to fix erroneous circuits in LLMs.
+Further research into the benefits of using known-good models to increase LLM trustworthiness and capabilities is justified.
 Moreover, developing methods to incorporate compact known-good models into LLMs, whether untrained or trained, could democratize AI Safety research.
 It enables small teams to focus on specific functional areas, crafting component models to a defined quality standard for integration into various LLMs.
-A significant challenge in our research was the sensitivity of model behavior to minor changes in the training environment.
-These changes often altered the specific nodes used for calculations, necessitating manual adjustments to our algorithm-validation test suite.
-A more effective strategy would be a declarative approach to describing each component in an algorithm hypothesis, allowing automatic identification of corresponding model nodes.
-For instance, a component description could specify nodes that primarily attend to question digits D2 and D’2, influence answer digit A3, and are crucial for S0 but not S1 complexity questions.
-Such a declarative language and its accompanying tooling would accelerate research progress and ensure that known-good components retain their functionality when integrated into LLMs.
-An initial version of this tooling has been created, but significant further work is required to fully realize its potential.
-Impact Statement
+Our automated framework, for discovering the location of algorithm sub-tasks in models, sped up our research.
+Our framework can be re-used to ensure that known-good components retain their functionality when integrated into LLMs.
+Some of our automated framework is specific to our maths models - much is generic and reuseable on other types of models.
+Our automation focuses more on attention heads and less on MLP layers.
+Extending our framework to integrate the output data from complimentary frameworks (e.g.
+ACDC’s node connectivity strength data (Conmy et al., 2023)) would provide more automated insights into a model’s behavior, further speeding up research.
+We made progress on a declarative language to describe an algorithm in terms of the number and types sub-tasks that must exist in a model for the algorithm to feasibly explain the model’s implementation.
+We made progress on a framework to automatically test an algorithm described in this way against a specific model.
+Further work is required to fully realize its potential.
+7 Impact Statement
 Our work aims to explain the inner workings of transformer-based language models, which may have broad implications for a wide range of applications.
@@
 We encourage responsible use of AI, aligning with ethical guidelines.
-Acknowledgements
-The authors would like to thank Apart Labs (https://apartresearch.com/lab) for their support in producing this paper and more generally supporting the entry of people new to AI Safety into the field.
-Appendix A Appendix: Model Configuration
-Addition experiments were done in Colab notebooks named “Accurate Addition - Train” and “Accurate Addition - Analyse”.
-They run on a T4 GPU.
-Training takes $\sim$15 mins.
-Analysis takes $\sim$8 mins.
-The key parameters (which are all configurable) are:
-  • n_layers = 2: This is a 2-layer Transformer model
-  • n_heads = 3: There are 3 attention heads
-  • n_digits = 5: Number of digits in the addition question
+Appendix A Appendix: Terminology
+These terms and abbreviations are used in this paper and the associated Colabs and python code:
+  • Pn : Model (input or output) token position. Zero-based. e.g. P18, P18L1H0
+  • Ln : Model layer n. Zero-based. e.g. P18L1H2
+  • Hn : Attention head n. Zero-based. e.g. P18L1H2
+  • Mn : MLP neuron n. Zero-based
+  • PnLnHn : Location / name of a single attention head, at a specified layer, at a specific token position
+  • PnLnMn : Location / name of a single MLP neuron, at a specified layer, at a specific token position
+  • D : First number of the pair question numbers
+  • Dn : nth numeric token in the first question number. Zero-based. D0 is the units value
+  • D’ : Second number of the pair question numbers
+  • D’n : nth token in the second question number. Zero-based. D0 is the units value
+  • A : Answer to the question (including answer sign)
+  • An : nth token in the answer. Zero-based. A0 is the units value. The highest token is the “+" or "-" answer sign
+  • S : Prefix for Addition. Think S for Sum. Aka ADD.
+  • SA : Basic Add. An addition sub-task. An.SA is defined as (Dn + D’n) % 10. e.g. 5 + 7 gives 2
+  • SC : Carry One. An addition sub-task. An.SC is defined as Dn + D’n >=>=> = 10. e.g. 5 + 7 gives True
+  • SS : Make Sum 9. An addition sub-task. An.SS is defined as Dn + D’n == 9. e.g. 5 + 7 gives False
+  • ST : TriCase. An addition sub-task. Refer paper 2 for details
+  • ST8, ST9, ST10 : Outputs of the ST TriCase sub-task.
+  • M : Prefix for Subtraction with a positive answer. Think M for Minus. Aka SUB
+  • : Basic Difference. A subtraction sub-task. An.MD is defined as (Dn - D’n) % 10. e.g. 3 - 7 gives 6
+  • : Borrow One. A positive-answer subtraction sub-task. An.MB is defined as Dn - D’n $<$ 0. e.g. 5 - 7 gives True
+  • MZ : Make Zero. A positive-answer subtraction sub-task. An.MZ is defined as Dn - D’n == 0. e.g. 5 - 5 gives True
+  • MT : TriCase. A positive-answer subtraction sub-task.
+  • MT1, MT0, MT-1 : Outputs of the MT TriCase sub-task.
+  • N : Prefix for Subtraction with a negative answer. Think N for Negative. Aka NEG
+  • ND : Basic Difference. A negative-answer subtraction sub-task. An.ND is defined as (Dn - D’n) % 10. e.g. 3 - 7 gives 6
+  • NB : Borrow One. A negative-answer subtraction sub-task. An.NB is defined as Dn - D’n $<$ 0. e.g. 5 - 7 gives True
+  • NZ : Make Zero. A negative-answer subtraction sub-task. An.NZ is defined as Dn - D’n == 0. e.g. 5 - 5 gives True
+  • NT : TriCase. A negative-answer subtraction sub-task.
+  • GT : Greater Than. A (positive-answer or negative-answer) subtraction sub-task. An.GT is defined as Dn $>$ D’n. e.g. 3 $>$ 5 gives False
+  • OPR : Operator. A sub-task that attends to the + or - token in the question (which determines whether the question is addition or subtraction).
+  • SGN : Sign. A sub-task that attends to the first answer token, which is + or -
+  • PCA : Principal Component Analysis
+  • EVR : Explained Variance Ratio. In PCA, EVR represents the percentage of variance explained by each of the selected components.
+Appendix B Appendix: Model Configuration
+Addition, subtraction and mixed (addition and subtraction) training experiments were done in a Colab notebook.
+The Colab runs on a T4 GPU.
+Training takes up to 35 mins.
+The key parameters (and their common configurations) are:
+  • n_layers = 1, 2 or 3: Number of layers.
+  • n_heads = 3 or 4: Number of attention heads.
+  • n_digits = 5, 6 or 10: Number of digits in the question.
+Each digit is represented as a separate token.
+(Liu and Low, 2023) state that LLaMa’s “remarkable arithmetic ability … is mainly atributed to LLaMA’s consistent tokenization of numbers".
+The model’s vocabulary contains 14 tokens ( 0, .., 9, +, -, =, *, / ) to enable this and planned future investigations.
 Training uses a new batch of data each step (aka Infinite Training Data) to minimise memorisation.
-During a training run the model processes about 1.5 million training datums.
+Depending on the configuration, each training run processes 1 to 4 million training datums.
 For the 5-digit addition problem there are 100,000 squared (that is 10 billion) possible questions.
 So the training data is much less than 1% of the possible problems.
-Because US cascades (e.g.
-44445+55555=100000, 54321+45679=1000000, 44450+55550=10000, 1234+8769=10003) are exceedingly rare, the data generator was enhanced to increase the likelihood of these cases turning up in the training data
-Mixed operation (addition and subtraction) experiments were done in Colab notebooks named “Accurate Math - Train” and “Accurate Math - Analyse”.
-They run on a T4 GPU.
-Training takes $\sim$25 mins.
-Analysis takes $\sim$8 mins.
-The key parameters (which are all configurable) are:
-  • n_layers = 3: This is a 3-layer transformer model
-  • n_heads = 4: There are 4 attention heads
-  • n_digits = 6: Number of digits in the addition and subtraction questions
-Training uses a new batch of data each step (aka Infinite Training Data) to minimise memorisation.
-In all Colabs, each digit is represented as a separate token. (Liu & Low, 2023) state that LLaMa’s “remarkable arithmetic ability … is mainly atributed to LLaMA’s consistent tokenization of numbers”.
-Appendix B Appendix: Addition Model Loss
-The 1-layer model is not very accurate (see Tab.
-5).
-With 2 layers the model gains accuracy (see Tab.
-6).
-The model uses batch_size = 64, n_heads = 3, lr = 0.00008, weight_decay = 0.1.
-The loss function is simple:
+Addition and subtraction include rare edge cases.
+For example, the SS cascades (e.g.
+44445+55555=100000, 54321+45679=1000000, 44450+55550=10000, 1234+8769=10003) are exceedingly rare.
+The data generator was enhanced to increase the frequency of all known edges cases.
+This lead to lower model loss.
+Validation test data covering all edge cases was manually constructed. These test cases are not used during training.
+Appendix C Appendix: Model Loss
+The model defaults to batch size = 64, learning rate = 0.00008 and weight decay = 0.1. The loss function is simple:
   • Per Digit Loss: For “per digit” graphs and analysis, for a given answer digit, the loss used is negative log likelihood.
   • All Digits Loss: For “all answer digits” graphs and analysis, the loss used is the mean of the “per digit” loss across all the answer digits.
-Table 5: Results for 1-layer 3-head model with 5, 10 & 15 digits
-No. digits 5 10 15
-Training batches 20K 20K 20K
-Loss 0.008314 0.040984 0.031746
-Heads used 15 29 46
-MLPs used 6 11 16
-Table 6: Results for 2-layer 3-head model with 5 & 10 digits
-No. digits 5 5 10 10
-Training batches 20K 30K 20K 30K
-Loss 2.7e-6 2.3e-8 9.4e-5 1.2e-4
-Heads used 21 19 57 51
-MLPs used 15 15 27 28
-The 2-layer, 5-digit model trained over 30 thousand batches has a final training loss of $\sim 2.3\times 10^{-8}$.
-Appendix C Appendix: Addition Model Shape Choice
+In our experimental models, the number of digits in the question varies from 5 to 10, the number of layers varies from 1 to 4, the number of heads varies from 3 to 4.
+Each experimental model’s loss is detailed in Tab.
+6.
+Appendix D Appendix: Addition Model Shape
 While we wanted a very low loss addition model, we also wanted to keep the model compact - intuiting that a smaller model would be easier to understand than a large model.
 Here are the things we tried to reduce loss that didn’t work:
-  • Increasing the frequency of hard (cascading US ) examples in the training data so the model has more hard examples to learn from. This improved training speed but did not reduce loss.
+  • Increasing the frequency of hard (cascading SS ) examples in the training data so the model has more hard examples to learn from. This improved training speed but did not reduce loss.
   • Increasing the number of attention heads from 3 to 4 or 5 (while still using 1 layer) to provide more computing power.
@@
   • Changing the n_layers to 2 and n_heads to 2.
-The smallest model shape that did reduce loss significantly was 2 layers with 3 attention heads
-Appendix D Appendix: TriAdd Implementation
-TriAdd transfers data from $(D_{n-1})$ to $(D_{n})$ by integrating the values of $(D_{n}.C)$ and $(D_{n-1}.C)$. This function can be represented as nine bigram mappings with three possible outputs.
-Table 7: $D_{n}$.TriAdd can be calculated from $D_{n}.C$ and $D_{n-1}.C$ through nine bigram mappings and yielding the three distinct outputs T8, T9 and T10
-$D_{n}.C$ =T8 $D_{n}.C$ =T9 $D_{n}.C$ =T10
-$D_{n-1}.C$ = T8 T8 T9 T10
-$D_{n-1}.C$ = T9 T8 T9 T10
- $D_{n-1}.C$ = T10 T9 * T10 T10
-Note that in the case $D_{n}.C$ = T9 and $D_{n-1}.C$ = T10, the answer is indeterminate.
-The result could be T8 or T9 but importantly it can not be T10.
-We choose to use T9 in our definition, but T8 would work just as well.
-Appendix E Appendix: Addition Hypothesis 1
+The smallest model shape that did reduce loss significantly was 2 layers with 3 attention heads.
+Appendix E Appendix: Experimental models
+Twenty-one experiment models were trained and analyzed (refer Tab. 6). The models and output of analysis are available on HuggingFace at https://huggingface.co/PhilipQuirke/VerifiedArithmetic
+Num Digits Num Layers Num Heads Training Steps Seed Addition / Sub Accuracy Training loss Heads used MLPs used
+Addition models
+5 1 3 30K 372001 two nines 9.4e-2 15 6
+5 2 3 15K 372001 six nines 1.6e-8 30 16
+5 2 3 40K 372001 six nines 2.0e-9 22 15
+6 2 3 15K 372001 six nines 1.7e-8 31 17
+6 2 3 20K 173289 six nines 1.5e-8 28 17
+6 2 3 20K 572091 six nines 7.0e-9 35 17
+6 2 3 40K 372001 six nines 2.0e-9 29 17
+10 2 3 20K 572091 six nines 7.0e-9 44 28
+Subtraction models
+6 2 3 30K 372001 six nines 5.8e-6 40 21
+10 2 3 75K 173289 two nines 2.0e-3 101 37
+Mixed models
+6 3 4 40K 372001 six / six nines 5.0e-9 54 26
+10 3 4 75K 173289 five / two nines 1.1e-6 143 53
+Mixed models initialized with addition model
+6 2 3 40K 572091 six / five nines 2.4e-8 57 21
+6 3 3 40K 572091 six / five nines 1.8e-8 70 35
+6 3 3 80K 572091 six / five nines 1.6e-8 75 35
+6 3 4 40K 372001 six / six nines 8.0e-9 72 26
+6 3 4 40K 173289 five / five nines 1.4e-8 60 29
+6 3 4 50K 572091 six / five nines 2.9e-8 79 29
+10 3 3 50K 572091 five / five nines 6.3e-7 90 45
+Mixed models initialized with add model. Reset useful heads every 100 steps
+6 4 4 40K 372001 five / five nines 1.7e-8 51 30
+Mixed models initialized with add model. Reset useful heads & MLPs every 100 steps
+6 4 3 40K 372001 four / two nines 3.0e-4 115 53
+Table 6: Main experimental models studied
+For each model the ’VerifiedArithmeticTrain’ Colab notebook generates two files:
+  • A “XXXXXX.pth" file containing the model weights
+  • A “XXXXXX_train.json" file containing configuration information and training loss data
+While, for each model the ’VerifiedArithmeticAnalysis’ Colab notebook generates two more files:
+  • A “XXXXXX_behavior.json" file containing generic “behavior" facts learnt about the model by the Colab e.g. P18L0H0 attends to tokens D3 and D’3
+  • A “XXXXXX_maths.json" file containing “maths-specific" facts learnt about the model by the Colab e.g. P18L0H0 performs the A3.SC sub-task.
+Appendix F Appendix: TriAdd Implementation
+TriAdd transfers data from $A_{n-1}$to $A_{n}$by integrating the values of $A_{n-1}.ST$ and $A_{n}.ST$ This function can be represented as nine bigram mappings with three possible outputs.
+(Refer Tab.7.)
+$A_{n}.ST$ $A_{n}.ST$ $A_{n}.ST$
+$A_{n-1}.ST$ = ST8 = ST9 = ST10
+ST8 ST8 ST9 ST10
+ST9 ST8 ST9 ST10
+ ST10 ST8 * ST10 ST10
+Table 7: $A_{n}$.TriAdd can be calculated from $A_{n}.ST$ and $A_{n-1}.ST$ through nine bigram mappings and yielding the three distinct outputs ST8, ST9 and ST10
+Note that in the case $A_{n}.ST$ = ST9 and $A_{n-1}.ST$ = ST10, the answer is indeterminate.
+The result could be ST8 or ST9 but importantly it can not be ST10.
+We choose to use ST8 in our definition, but ST9 would work just as well.
+Appendix G Appendix: Complexity
+To analyze question difficulty, we categorized addition questions by the complexity of the computation required to solve the question, as shown in Tab.
+8.
+The categories are arranged according to the number of digits that a carry bit has to cascade through.
+Name Contains Example Freq
+S0 SA 11111+12345=23456 $\sim$5%
+S1 SA,SC 11111+9=22230 $\sim$21%
+S2 SA,SCx2 11111+89=22300 $\sim$34%
+S3 SA,SCx3 11111+889=23000 $\sim$28%
+S4 SA,SCx4 11111+8889=30000 $\sim$11%
+S5 SA,SCx5 11111+88889=100000 $\sim$2%
+Table 8: We categorise addition questions into non-overlapping “calculation complexity" quanta, ordered by increased computational difficulty (and decreasing occurrence frequency).
+Five-digit addition questions quanta are S0 to S5.
+Ten-digit addition question quanta are S0 to S10.
+S10’s frequency is $\sim 3e-4$ showing the need to enrich training data for rare edge cases.
+Appendix H Appendix: Addition Hypothesis 1
 Given the 2-layer attention pattern’s similarity to 1-layer attention pattern, and the above evidence, our first (incorrect) hypothesis was that the 2-layer algorithm:
-  • Is based on the same BA, MC and US operations as the 1-layer.
-  • Uses the new early positions to (somehow) do the US calculations with higher accuracy than the 1-layer model.
+  • Is based on the same SA, SC and SS operations as the 1-layer.
+  • Uses the new early positions to (somehow) do the SS calculations with higher accuracy than the 1-layer model.
   • The long double staircase still finalises each answer digit’s calculation.
-  • The two attention nodes in the long double staircase positions do the BA and MC calculations and pull in US information calculated in the early positions.
+  • The two attention nodes in the long double staircase positions do the SA and SC calculations and pull in SS information calculated in the early positions.
 If this is correct then the 2-layer algorithm successfully completes these calculations:
-  • A0 = D0.BA
-  • A1 = D1.BA + D0.MC
-  • A2 = D2.BA + (D1.MC or (D1.MS & D0.MC))
-  • A3 = D3.BA + (D2.MC or (D2.MS & D1.MC) or (D2.MS & D1.MS & D0.MC))
-  • A4 = D4.BA + (D3.MC or (D3.MS & D2.MC) or (D3.MS & D2.MS & D1.MC) or (D3.MS & D2.MS & D1.MS & D0.MC))
-  • A5 = D4.MC or (D4.MS & D3.MC) or (D4.MS & D3.MS & D2.MC) or (D4.MS & D3.MS & D2.MS & D1.MC) or (D4.MS & D3.MS & D2.MS & D1.MS & D0.MC)
+  • A0 = A0.SA
+  • A1 = A1.SA + A0.SC
+  • A2 = A2.SA + (A1.SC or (A1.SS & A0.SC))
+  • A3 = A3.SA + (A2.SC or (A2.SS & A1.SC) or (A2.SS & A1.SS & A0.SC))
+  • A4 = A4.SA + (A3.SC or (A3.SS & A2.SC) or (A3.SS & A2.SS & A1.SC) or (A3.SS & A2.SS & A1.SS & A0.SC))
+  • A5 = A4.SC or (A4.SS & A3.SC) or (A4.SS & A3.SS & A2.SC) or (A4.SS & A3.SS & A2.SS & A1.SC) or (A4.SS & A3.SS & A2.SS & A1.SS & A0.SC)
 Our intuition is that there are not enough useful nodes in positions 8 to 11 to complete the A5 calculation this way. So we abandoned this hypothesis.
-Appendix F Appendix: Addition Hypothesis 2
+Appendix I Appendix: Addition Hypothesis 2
 Our second (incorrect) hypothesis was that the 2-layer algorithm has a more compact data representation, so it can pack more calculations into each node, allowing it to accurately predict A5 in step 11.
-We claimed the model stores the sum of each digit-pair as a single token in the range “0” to “18” (covering 0+0 to 9+9). We name this operator $D_{n}$.T, where T stands for “token addition”:
-  • $D_{n}$.T = $D_{n}$+ $D^{\prime}_{n}$
-The $D_{n}$.T operation does not understand mathematical addition. Tab. 8 shows how the model implements the T operator as a bigram mapping.
-Table 8: Implementing the T operator as a bigram mapping from 2 input tokens to 1 result token.
+We claimed the model stores the sum of each digit-pair as a single token in the range “0” to “18” (covering 0+0 to 9+9). We name this operator $A_{n}$.T, where T stands for “token addition”:
+  • $A_{n}$.T = $D_{n}$+ $D^{\prime}_{n}$
+The $A_{n}$.T operation does not understand mathematical addition. Tab. 9 shows how the model implements the T operator as a bigram mapping.
 $D_{n}$vs $D^{\prime}_{n}$ 0 1 4 5 8 9
@@
 9 9 10 13 14 17 18
-$D_{n}$.T is a compact way to store data. Tab. 9 show how, if it needs to, the model can convert a $D_{n}$.T value into a one-digit-accuracy BA, MC or MS value.)
-Table 9: Converting a $D_{n}$.T value into a BA , MC or MS value.
-$D_{n}$.T $D_{n}$.BA $D_{n}$.MC $D_{n}$.MS
+Table 9: Implementing the T operator as a bigram mapping from 2 input tokens to 1 result token.
+$A_{n}$.T is a compact way to store data. Tab. 10 show how, if it needs to, the model can convert a $A_{n}$.T value into a one-digit-accuracy SA, SC or SS value.)
+$A_{n}$.T $A_{n}$.SA $A_{n}$.SC $A_{n}$.SS
 0 0 0 0
@@
 18 8 1 0
-Our notation shorthand for one-digit-accuracy these “conversion” bigram mappings is:
-  • $D_{n}$.BA = ($D_{n}$.T % 10) where % is the modulus operator
-  • $D_{n}$.MC = ($D_{n}$.T // 10) where // is the integer division operator
-  • $D_{n}$.MS = ($D_{n}$.T == 9) where == is the equality operator
-The D0.T value is perfectly accurate.
-But the other $D_{n}$.T values are not perfectly accurate because each is constrained to information from just one digit.
-We define another more accurate operator $D_{n}$.T2 that has “two-digit accuracy”.
-$D_{n}$.T2 is the pair sum for the nth digit plus the carry bit (if any) from the n-1th digit T:
-  • $D_{n}$.T2 = $D_{n}$.T + $D_{n-1}$.MC
-$D_{n}$.T2 is more accurate than $D_{n}$.T.
-The $D_{n}$.T2 value is always in the range “0” to “19” (covering 0+0+0 to 9+9+MakeCarry).
+Table 10: Converting a $A_{n}$.T value into a SA , SC or SS value.
+Our notation shorthand for one-digit-accuracy these “conversion" bigram mappings is:
+  • $A_{n}$.SA = ($A_{n}$.T % 10) where % is the modulus operator
+  • $A_{n}$.SC = ($A_{n}$.T // 10) where // is the integer division operator
+  • $A_{n}$.SS= ($A_{n}$.T == 9) where == is the equality operator
+The A0.T value is accurate.
+But the other $A_{n}$.T values are not accurate because each is constrained to information from just one digit.
+We define another more accurate operator $A_{n}$.T2 that has “two-digit accuracy”.
+$A_{n}$.T2 is the pair sum for the nth digit plus the carry bit (if any) from the n-1th digit T:
+  • $A_{n}$.T2 = $A_{n}$.T + $A_{n-1}$.SC
+$A_{n}$.T2 is more accurate than $A_{n}$.T.
+The $A_{n}$.T2 value is always in the range “0” to “19” (covering 0+0+0 to 9+9+CarryOne).
 Tab.
 11 show how the model can implement the T2 operator as a mapping.
-Table 10: Calculating $D_{n}$.T2 from $D_{n}$.T and $D_{n-1}$.T
-$D_{n}$.T $D_{n}$.T2 if $D_{n-1}$.MC==0 $D_{n}$.T2 if $D_{n-1}$.MC==1
+$A_{n}$.T $A_{n}$.T2 if $A_{n-1}$.SC==0 $A_{n}$.T2 if $A_{n-1}$.SC==1
 0 0 1
@@
 18 18 19
-Following this pattern, we define operators $D_{n}$.T3, $D_{n}$.T4 and $D_{n}$.T5 with 3, 4 and 5 digit accuracy respectively:
-  • $D_{n}$.T3 = $D_{n}$.T + ( $D_{n-1}$.T2 // 10 ) aka $D_{n}$.T + $D_{n-1}$.MC2
-  • $D_{n}$.T4 = $D_{n}$.T + ( $D_{n-1}$.T3 // 10 ) aka $D_{n}$.T + $D_{n-1}$.MC3
-  • $D_{n}$.T5 = $D_{n}$.T + ( $D_{n-1}$.T4 // 10 ) aka $D_{n}$.T + $D_{n-1}$.MC4
-The value D4.T5 is perfectly accurate as it integrates MC and cascading MS data all the way back to and including D0.T.
-The values D1.T2, D2.T3, D3.T4 are also all perfectly accurate.
-If the model knows these values it can calculate answer digits with perfect accuracy:
-  • A1 = D1.T2 % 10 with zero loss
-  • A2 = D2.T3 % 10 with zero loss
-  • A3 = D3.T4 % 10 with zero loss
-  • A4 = D4.T5 % 10 with zero loss
-  • A5 = D4.T5 // 10 with zero loss
-Tab.
-11 shows how the hypothesis 2 mathematical framework is applied and Figure 11 shows how this maps to the model.
-There are nodes that the model needs that this framework does not use, so our hypothesis is not 100% right.
-Table 11: Accurately predicting A5 in step 11 and A4 in step 12 using addition hypothesis 2
-Calculation Node
-D1.T = D1 + D’1
-P9.L0.H1 and P9.L0.MLP
-D2.T = D2 + D’2
-P8.L0.H1 and P8.L0.MLP
-D1.T2 = D2.T + (D1.T + (D0 + D0’) // 10) // 10
-P10.L0.H1 and P10.L0.MLP
-D3.T4 = D3 + D’3 + D2.T3 // 10
-P11.L0.H1
-D4.T = D4 + D’4
-P11.L0.H2
-D3.MC3 = D3.T4 // 10
-P11.L0.MLP
-A5 = (D4.T + D3.MC3) // 10
-P11.L1.MLP
-D4.T5 = D4.T + D3.T4 // 10
-P12.L0.H0
-A4 = D4.T5 % 10 // 10
-P12.L0.MLP
-Figure 11: The 5-digit 2-layer integer addition model algorithm under hypothesis 2
-In this hypothesis, all the answer digits are perfectly calculated using the nodes in positions 8 to 11.
+Table 11: Calculating $A_{n}$.T2 from $A_{n}$.T and $A_{n-1}$.T
+Following this pattern, we define operators $A_{n}$.T3, $A_{n}$.T4 and $A_{n}$.T5 with 3, 4 and 5 digit accuracy respectively:
+  • $A_{n}$.T3 = $A_{n}$.T + ( $A_{n-1}$.T2 // 10 ) aka $A_{n}$.T + $A_{n-1}$.SC2
+  • $A_{n}$.T4 = $A_{n}$.T + ( $A_{n-1}$.T3 // 10 ) aka $A_{n}$.T + $A_{n-1}$.SC3
+  • $A_{n}$.T5 = $A_{n}$.T + ( $A_{n-1}$.T4 // 10 ) aka $A_{n}$.T + $A_{n-1}$.SC4
+The value A4.T5 is accurate as it integrates SC and cascading SS data all the way back to and including A0.T.
+The values A1.T2, A2.T3, A3.T4 are also all accurate.
+If the model knows these values it can calculate answer digits accurately:
+  • A1 = A1.T2 % 10
+  • A2 = A2.T3 % 10
+  • A3 = A3.T4 % 10
+  • A4 = A4.T5 % 10
+  • A5 = A4.T5 // 10
+In this hypothesis, all the answer digits are accurately calculated using the nodes in positions 8 to 11.
 This hypothesis 2 is feasible, elegant and compact - reflecting the authors (human) values for good code.
-Experimenation shows the model does not implement this hypothesis. It retains the long staircase BA calculations in positions 11 to 16. Why? Two reasons suggest themselves:
+Experimenation shows the model does not implement this hypothesis. It retains the long staircase SA calculations in positions 11 to 16. Why? Two reasons suggest themselves:
   • Hypothesis 2 is too compact.
@@
 We abandoned this hypothesis.
-Appendix G Appendix: Five-Digit Addition Experiments
-Figure 12: Our 5-digit integer addition question has 12 question tokens ($D_{n}$,$D^{\prime}_{n}$) followed by 6 answer tokens ($A_{n}$).
-This known-good 2-layer 3-head model can perform one million predictions without error.
-The model’s algorithm co-ordinates computation across 9 token positions ($P_{n}$) using 19 attention heads ($H_{n}$) and 15 MLP layers.
-Three attention heads produce intermediate results that are reused in multiple prediction steps e.g.
-P10.L0.H1 is reused in 4 prediction steps to calculate answer digits A5 to A2.
-The hypothesis 3 pseudo-code was derived iteratively by obtaining experimental results and mapping them to mathematical operations.
-Figure 12 shows the pseudo-code diagramtically.
-Some of the experiments and mappings were:
+Appendix J Appendix: Addition Hypothesis 3
+The hypothesis 3 pseudo-code was derived iteratively by obtaining experimental results and mapping them to mathematical operations. Some of the experiments and mappings were:
   • Ablation experiments show that the A5 value is accurately calculated in prediction step 11 using 5 attention heads and 5 MLP layers.
 The pseudo-code accurately calculates A5 while constraining itself to this many steps.
-  • Ablating the nodes one by one shows which answer digit(s) are reliant on each node (Ref Figure 7).
+  • Ablating the nodes one by one shows which answer digit(s) are reliant on each node (Ref Table 1).
 Most interestingly, ablating P10.L0.H1 impacts the answer digits A5, A4, A3, A2 (but not A1 and A0).
@@
 These relationships are constraints that are all obeyed by the pseudo-code.
-  • The pseudo-code has 4 instances where $D_{n}.C$ is calculated using TriCase.
+  • The pseudo-code has 4 instances where $A_{n}.ST$ is calculated using TriCase.
 PCA of the corresponding nodes (P8.L0.H1, P9.L0.H1, P11.L0.H2 and P14.L0.H1) shows tri-state output for the specified $D_{n}$.
@@
 (see Figure 8).
-  • Activation patching (aka interchange intervention) experiments at attention head level confirmed some aspects of the calculations (see § H for details.
-  • The pseudo code includes calculations like D1.C which it says is calculated in P9.L0.H1 and P9.L0.MLP.
+  • Activation patching (aka interchange intervention) experiments at attention head level confirmed some aspects of the calculations (see § K for details.
+  • The pseudo code includes calculations like A1.ST which it says is calculated in P9.L0.H1 and P9.L0.MLP.
 Ablation tells us both nodes are necessary.
 For the attention head we use the PCA results for insights.
-We didn’t implement a similar investigative tool for the MLP layer, so in the pseudo-code we attribute the calculation of D1.C to both nodes.
+We didn’t implement a similar investigative tool for the MLP layer, so in the pseudo-code we attribute the calculation of A1.ST to both nodes.
   • For P10.L0.H1, the attention head PCA could represent either a bi-state or tri-state output (see Figure 9).
@@
   • For P15.L0.H1 the attention head PCA could represent either a bi-state or tri-state output (see Figure 9).
-The pseudo-code shows a bi-state calculation D0.MC at P15.L0.H1, but with small alterations the pseudo-code would work with a tri-state output.
-  • The calculation of D1.C2 in P14.L0.H1 is a interesting case.
-The model needs D1.C2 for A2 accuracy.
-The model could simply reuse the accurate D1.C2 value calculated in P10.
+The pseudo-code shows a bi-state calculation A0.SC at P15.L0.H1, but with small alterations the pseudo-code would work with a tri-state output.
+  • The calculation of A1.ST2 in P14.L0.H1 is a interesting case.
+The model needs A1.ST2 for A2 accuracy.
+The model could simply reuse the accurate A1.ST2 value calculated in P10.
 Activation patching shows that it does not.
-Instead the P14 attention heads calculate D1.C1 from D1 and D’1 directly, and only relies on the P10.D1.C2 value in the case where D1.C2 != D1.C.
-That is, the calculation is “use P14.D1.C1 value else use P10.D1.C2 values”.
-This aligns with the model learning the P10.D1.C calculation early in training (for 90% accuracy) and later learning that P10.D1.C2 contains additional information it can use to get to 100% accuracy.
+Instead the P14 attention heads calculate A1.ST1 from D1 and D’1 directly, and only relies on the P10.D1.ST2 value in the case where A1.ST2 != A1.ST.
+That is, the calculation is “use P14.A1.ST1 value else use A1.ST2 values".
+This aligns with the model learning the P10.A1.ST calculation early in training (for 90% accuracy) and later learning that P10.A1.ST2 contains additional information it can use to get to six nines accuracy.
+Figure 5: For a sample 5-digit 2-layer 3-head addition model, this map shows a compacted view of all useful token positions (horizontally) and all useful attention heads and MLP layers (vertically) used in predictions as green cells.
+Each cell shows the simpliest (lowest complexity) quanta S0, S1, etc impacted when we ablate each node.
+To answer S0 questions, only the S0 nodes are used.
+To answer S1 questions, S0 and S1 nodes are used, etc.
+The model only uses nodes in nine token positions.
+Figure 6: This map shows the % of enriched questions that fail when we ablate each node in a 5-digit 2-layer 3-head addition model.
+The model only uses nodes in token positions P8 to P16 (i.e.
+tokens D’2 to A1).
+Lower percentages correspond to rarer edge cases.
+The grey space represents nodes that are not used by the model.
+Figure 7: This map shows the input tokens each attention head attends to at each token position in a 5-digit 2-layer 3-head addition model.
+At token position P11 the model predicts the first answer digit A5.
+All digit pairs (e.g.
+D2 D’2) are attended to by P11.
 Figure 8: For 5-digit addition, for these attention heads, for exactly 1 answer digit $A_{n}$each, PCA shows these interpretable results.
@@
 The PCA and TriCase data are strongly correlated, but the PCA data has 2 states.
-Appendix H Appendix: Addition Interchange Interventions
-To test the hypothesis 3 mapping of the mathematical framework (casual abstraction) to the model attention heads, various “interchange interventions” (aka activation patching) experiments were performed on the model, where
+Appendix K Appendix: Addition Interchange Interventions
+To test the hypothesis 3 mapping of the mathematical framework (casual abstraction) to the model attention heads, various “interchange interventions" (aka activation patching) experiments were performed on the model, where
   • A particular claim about an attention head has selected for testing.
@@
   • The model then predicted answers for more questions, but this time we intervened during the prediction to override the selected attention head activations with the activations from the previous run.
-Using this approach we obtained the findiings in Tab. 12.
-Table 12: Interchange Interventions experiments used activation patching to test the claims addition hypothesis 3 made for each attention head.
+Using this approach we obtained the findings in Tab. 12.
+Nodes Claim: Attention head(s) perform … Finding: Attention head(s) perform a function that …
+P8.L0.H1 and MLP A2.ST = TriCase( D2, D’2 ) impacting A4 and A5 accuracy Based on D2 and D’2. Triggers on a A2 carry value. Provides carry bit used in A5 and A4 calculation.
+P9.L0.H1 and MLP A1.ST = TriCase( D1, D’1 ) impacting A5, A4 and A3 accuracy Based on D1 and D’1. Triggers on a A1 carry value. Provides carry bit used in A5, A4 and A3 calculation.
+P10.L0.H1 and MLP A1.ST2 = TriAdd( A1.ST, TriCase(D0, D’0) ) impacting A5, A4, A3 and A2 accuracy Based on D0 and D’0.
+Triggers on a A0 carry value.
+Provides carry bit used in A5, A4, A3 and A2 calculation.
+P11.L0.H1 and MLP A3.ST4 = TriAdd( TriCase( D3, D’3 ), TriAdd( A2.ST, A1.ST2)) impacting A5 accuracy Based on D3 and D’3. Triggers on a A3 carry value. Provides carry bit used in A5 calculations.
+P11.L0.H2 and MLP A4.ST = TriCase(D4, D’4) impacting A5 accuracy A4 Based on D4 and D’4. Triggers on a A4 carry value. Provides carry bit used in A5 calculation.
+P12.L0.H0+H2 and MLP A4.SA = (D4 + D’4) % 10 impacting A4 accuracy A4 Sums D4 and D’4. Impacts A4.
+P13.L0.H0+H2 and MLP A3.SA = (D3 + D’3) % 10 impacting A3 accuracy Sums D3 and D’3. Impacts A3.
+P14.L0.H0+H2 and MLP A2.SA = (D2 + D’2) % 10 impacting A2 accuracy Sums D2 and D’2. Impacts A2.
+P14.L0.H1 and MLP (D1 + D’1) / 10 + P10.A1.ST2 info impacting A2 accuracy Calculates P10.A1.ST1 but add P10.A1.ST2 info when A1.ST != A1.ST2. Impacts A2
+P15.L0.H0+H2 and MLP A1.SA = (D1 + D’1) % 10 impacting A1 accuracy Sums D1 and D’1. Impacts A1.
+P15.L0.H1 and MLP A0.SC = (D0 + D’0) / 10 impacting A1 accuracy Triggers when D0 + D’0 >10. Impacts A1 digit by 1
+P16.L0.H0+H2 and MLP A0.SA = (D0 + D’0) % 10 impacting A0 accuracy Sums D0 and D’0. Impacts A0.
+Table 12: Interchange Interventions experiments used activation patching to test the claims addition hypothesis 3 made for each attention head in a sample mixed model.
 Experimental results are consistent with hypothesis 3 for all nodes.
-Nodes Claim: Attention head(s) perform … Finding: Attention head(s) perform a function that …
-P8.L0.H1 and MLP
-D2.C = TriCase( D2, D’2 ) impacting A4 and A5 accuracy
-Based on D2 and D’2. Triggers on a D2 carry value. Provides carry bit used in A5 and A4 calculation.
-P9.L0.H1 and MLP
-D1.C = TriCase( D1, D’1 ) impacting A5, A4 and A3 accuracy
-Based on D1 and D’1. Triggers on a D1 carry value. Provides carry bit used in A5, A4 and A3 calculation.
-P10.L0.H1 and MLP
-D1.C2 = TriAdd( D1.C, TriCase(D0, D’0) ) impacting A5, A4, A3 and A2 accuracy
-Based on D0 and D’0. Triggers on a D0 carry value. Provides carry bit used in A5, A4, A3 and A2 calculation.
-P11.L0.H1 and MLP
-D3.C4 = TriAdd( TriCase( D3, D’3 ), TriAdd( D2.C, D1.C2)) impacting A5 accuracy
-Based on D3 and D’3. Triggers on a D3 carry value. Provides carry bit used in A5 calculations.
-P11.L0.H2 and MLP
-D4.C = TriCase(D4, D’4) impacting A5 accuracy A4
-Based on D4 and D’4. Triggers on a D4 carry value. Provides carry bit used in A5 calculation.
-P12.L0.H0+H2 and MLP
-D4.BA = (D4 + D’4) % 10 impacting A4 accuracy A4
-Sums D4 and D’4. Impacts A4.
-P13.L0.H0+H2 and MLP
-D3.BA = (D3 + D’3) % 10 impacting A3 accuracy
-Sums D3 and D’3. Impacts A3.
-P14.L0.H0+H2 and MLP
-D2.BA = (D2 + D’2) % 10 impacting A2 accuracy
-Sums D2 and D’2. Impacts A2.
-P14.L0.H1 and MLP
-(D1 + D’1) / 10 + P10.D1.C2 info impacting A2 accuracy
-Calculates P10.D1.C1 but add P10.D1.C2 info when D1.C != D1.C2. Impacts A2
-P15.L0.H0+H2 and MLP
-D1.BA = (D1 + D’1) % 10 impacting A1 accuracy
-Sums D1 and D’1. Impacts A1.
-P15.L0.H1 and MLP
-D0.MC = (D0 + D’0) / 10 impacting A1 accuracy
-Triggers when D0 + D’0 >10. Impacts A1 digit by 1
-P16.L0.H0+H2 and MLP
-D0.BA = (D0 + D’0) % 10 impacting A0 accuracy
-Sums D0 and D’0. Impacts A0.
-Appendix I Appendix: Six-Digit Addition
-We trained a very low loss 6-digit Addition model.
-We compared the 5-digit and 6-digit model algorithms.
-Figures 6 and  8 are very instructive.
-So are Figures 7 and  17.
-We concluded that the 6-digit model used the same logical algorithm as the 5-digit model.
-There were some implementation differences:
-  • The models selected different attention heads in the early positions to use to do the same logical calculations.
-  • The 5-digit model used 2 attention heads per digit to do the BA calculation, whereas 6-digit model only uses one (and so is more compact).
-  • The PCA trigrams have difference appearances in the two models. Figures 13 and  18.
-  • For 6-digits, for A2 (only), the BA calculation is done in P18, but the MC calculation is not in P18.
-This is unique.
-The 6-digit model has optimised out the MC and instead relies solely on the accurate D1.C2 value (and so is more compact).
++ve Sub Contains -ve Sub Contains Like
+M0 MD N/A N/A S0
+M1 MD,MB N1 ND,NB S1
+M2 MD,MBx2 N2 ND,NBx2 S2
+M3 MD,MBx3 N3 ND,NBx3 S3
+M4 MD,MBx4 N4 ND,NBx4 S4
+Table 13: We define “positive-answer subtraction" and “negative-answer subtraction" calculation complexity quanta that parallel the addition quanta.
+Appendix L Appendix: N-Digit Addition
+The addition models perform addition accurately. Visualizations that provided insights into the behavior of the model, aiding our interpretation of the algorithm, are below:
+Figure 10: This map shows the simpliest (lowest complexity) quanta S0, S1, etc impacted when we ablate each node in the 5-digit 2-layer 3-head addition model.
+To answer S0 questions, only the S0 nodes are used.
+To answer S1 questions, S0 and S1 nodes are used, etc.
 Figure 11: This map shows the % of questions that fail when we ablate each node in the 6-digit 2-layer 3-head addition model.
@@
 The grey space represents nodes that are not useful.
-Figure 16: This map shows the input tokens each attention head attends to at each token position.
-The model predicts the first answer digit A6 in token position P14, and all digit-pair sums (e.g.
-D2 D’2) are attended to in P11 to P14.
-Note the similarity to Figure 5.
-Figure 17: This map shows the answer digit(s) A0 .. A5 impacted when we ablate each useful node in the 6-digit 2-layer 3-head addition model. Note the similarity to Figure 7.
-Figure 18: For 9 attention heads, for exactly 1 answer digit $A_{n}$each, PCA shows these interpretable results for the 6-digit Addition model.
-The 3 dot colours show the TriCase value of each question.
-The PCA data and TriCase quanta are both tri-state and strongly correlated.
-Appendix J Appendix: Six-Digit Subtraction
-Integer subtraction (e.g.
-654321-789012=-0134691) and integer addition have similarities.
-Subtraction is conceptually more complex than integer addition.
-While $a+b$ and $b+a$ give the same answer, generally $a-b$ and $b-a$ give difference answers.
-In human subtraction the algorithm used to calculate $a-b$ changes depending on whether $a>b$ or $a<b$.
-In our transformer-based models, Subtraction and Addition have some similar quanta. For example:
-  • In addition, the first answer digit is always and “1” or “0” and this token is the hardest to calculate accurately because of the “cascading carry 1” edge case.
-In subtraction, the first answer token is always “+” or “-” and this token is the hardest to calculate accurately because of the “cascading borrow 1” edge case.
-  • If $a>b$ then the algorithm to calculate $a-b$ and $a+b$ have strong structural similarities, allowing per-digit calculations with one sub-task moving data between digits (“carry 1” or “borrow 1”)
-  • If $a<b$ then the algorithm to calculate $a-b$ can use the mathematical identity that $1a-b=-(b-a)$.
-Our intuition is that the model will calculate if $a<b$, and if so will reuse the $a>b$ circuit, then negate the answer.
-We created a very-low-loss 6-digit addition model.
-Figure 8 shows the lowest quanta impacted when ablating each node.
-We created a “mixed operations” model (6-digit, 3-layer, 4-head) to do both Addition and Subtraction.
-We initialised the untrained model with the addition model and trained it (with 80% subtraction and 20% addition questions) to be very low loss.
-This mixed model can predict one million addition and one million subtraction questions without error.
-Figure 19: The per-digit log-loss curves for a 6-digit 3-layer 4-head mixed model trained to perform subtraction and addition after being initialised with a very low loss addition model.
-To investigate what classes of questions the model is always getting wrong, we expanded our existing addition-specific metrics, to include subtraction-specific question categorisations (see Sec.
-4.3.
-Figures 9 and  10 maps these quanta against nodes showing which are used for addition and subtraction.
+Some notes about the models:
+  • The models selected different attention heads in the early positions to use to do the same logical calculations.
+  • Some models use 2 attention heads per digit to do the SA calculation, whereas some models only uses one (and so are more compact).
+  • The PCA trigrams have difference appearances in different models (but the same interpretable clusters). Refer Figures 8
+  • Per answer digit, some models use the SC calculation, whereas some models optimize it out and rely solely on the $ST$ value (and so are more compact).
+Appendix M Appendix: Mixed Model Initialization
+We experimented with three approaches to re-using the trained addition model in the “mixed" (addition and subtraction) model:
+  • Initialize Only: Initialize the untrained mixed model with the addition model weights before training begins.
+  • Freeze Attention: As per “Initialize Only", but also every 100 training steps recopy the attention head weights from the addition model into the partially-trained mixed model.
+  • Freeze All: As per “Initialize Only", but also every 100 training steps recopy the entire addition model (attention heads and MLP layers) into the partially-trained mixed model.
+Our intuition was that “Initialize Only" would give the mixed model the most freedom to learn new algorithms, but that the “Freeze Attention" and “Freeze All" approaches would make the resulting trained mixed model easier to interpret (as we could reuse our addition model insights).
+After experimentation we found that the “Initialize Only" approach was the only one that quickly trained to be able to do both addition and subtraction accurately.
+We concluded that the other two methods constrainws the model’s ability to learn new algorithms too much.
+We also experimented with “where" in the model we inserted the addition (6-digit, 2-layer, 3-head) model into the slightly larger (6-digit, 3-layer, 4-head) mixed model.
+That is, do we initialize the first 2 layers or the last 2 layers of the mixed model?
+Also do we initialize the first 3 attention heads or the last 3 attention heads of the mixed model?
+Our intuition was that initializing the first layers and heads would be more likely to cause the model to re-use the addition circuits adding interpretability, so we used this approach.
+Appendix N Appendix: N-Digit Subtraction
+The mixed models perform addition and subtraction accurately. Visualizations that provided insights into the behavior of the model, aiding our interpretation of the algorithm, are below:
+Figure 12: This map of a sample 6-digit mixed model shows the 98 nodes used to predict answers to addition (S), positive-answer subtraction (M) and/or negative-answer subtraction (N) questions.
+Before training the mixed model, 48 nodes were initalized pre-training with a smaller addition model’s weights.
+These are have a red border.
+During mixed model training, 39 of 48 of the initalized monosemantic nodes were generalized (become poly-semantic) and now help predict two or three question classes.
+Figure 13: This map shows the simplest complexity quanta S0, S1, etc used in each useful node of the 6-digit 3-layer 4-head mixed model when doing addition questions.
+Figure 14: This map shows the simpliest complexity quanta M0, M1, etc used in each useful node of the 6-digit 3-layer 4-head mixed model for subtraction questions with positive answers.
+Some notes about the mixed models:
+  • All the notes about the addition model (above) also apply to the mixed model.
   • The model contains a new sub-task that stands out: The algorithm relies on calculations done at token position P0, when the model has only seen one question token!
 What information can the model gather from just the first token?
-Intuitively, if the first token is a “8” or “9” then the first answer token is more likely to be a “+” (and not a “-”).
+Intuitively, if the first token is a “8" or “9" then the first answer token is more likely to be a “+" (and not a “-").
 The model uses this heuristic even though this probabilistic information is sometimes incorrect and so will work against the model achieving very low loss.
-Figure 20: This map shows the answer digit(s) A0 ..
-A7 impacted when we ablate each useful node in the 6-digit 3-layer 4-head mixed model.
-A7 is the sign of the answer and calculated first.
-Note the similarity to the 5 and 6-addition Figures 7 and  17.

### v8

@@
 We find extensive reuse of the addition circuits for both tasks, easing verification of the more complex subtractor model.
-We discuss how inserting verified task modules into LMs can leverage model reuse to improve verifiability and trustworthiness of language models built using them.
-The reuse of verified circuits reduces the effort to verify more complex composite models which we believe to be a significant step towards safety and interpretability of language models.
+We discuss how inserting verified task modules into LMs can leverage model reuse to improve verifiability and trustworthiness of LMs built using them.
+The reuse of verified circuits reduces the effort to verify more complex composite models which we believe to be a significant step towards safety and interpretability of LMs.
 Increasing Trust in Language Models through the Reuse of Verified Circuits
-Philip Quirke Apart Research.
-philipquirkenz@gmail.com                        Clement Neo Apart Research.
-Fazl Barez Apart Research.
-Department of Engineering Sciences, University of Oxford fazl@robots.ox.ac.uk
+Philip Quirkethanks:    Core research contributor., Clement Neo†‡, Fazl Barezfootnotemark: †‡ Apart Research, University of Oxford
 1 Introduction
@@
 This understanding aids in predicting model behavior in new situations and fixing model errors.
-In creating and training a model, we would ideally want the model to have a very high standard of accuracy and trustworthiness.
-We aim to achieve this by holding the model to a standard that we term known-good.
+In creating and training a model, we aim for high accuracy and trustworthiness.
+We achieve this by holding the model to a standard we term known-good.
 We define a model performing a task to be known-good if:
@@
 A known-good model must incorporate algorithms to manage all known edge cases.
-A known-good model would hence have verified [^1] In this paper, ’verified’ has the mechanistic interpretability meaning that a specific group of interconnected neurons within a neural network reliably and causally contributes to the model’s output in a meaningful, understandable way, with supporting empirical evidence.
+A known-good model must have verified [^1] In this paper, ’verified’ has the mechanistic interpretability meaning that a specific group of interconnected neurons within a neural network reliably and causally contributes to the model’s output in a meaningful, understandable way, with supporting empirical evidence.
 circuits that perform the task accurately.
@@
 The model achieves a very low training loss and has six nines (99.9999%) accuracy.
-The model hence achieves our standard of ’known-good’.
+The model hence achieves our known-good standard.
 Additionally, we develop a “mixed" model capable of both addition and subtraction, incorporating the known-good addition model.
-This mixed model has six nines accuracy, and extensively reuses the addition circuits for both operations and learns new circuits, facilitating the interpretation of the model’s algorithm.
-This approach marks progress toward a known-good model for both addition and subtraction.
+This mixed model has six nines accuracy, and extensively reuses the addition circuits for both operations, facilitating the interpretation of the model’s algorithm.
+We make progress toward a known-good model for both addition and subtraction.
 Hence, our main contributions are three-fold:
   • Defining several known-good n-digit addition models with six nines accuracy which all use the same algorithm.
-  • Demonstrating a proof of concept for re-using a known good model in the training of another larger, more-capable model, simplifying the interpretation of the more-capable model’s algorithm.
+  • Demonstrating a proof of concept for re-using a known good model in the training of another larger, more-capable model, simplifying the interpretation of the new model’s algorithm.
   • Defining several n-digit addition and subtraction models with six nines accuracy, that reuse established addition model circuits for both operations, and detailing progress towards these models being known-good.
@@
 3 Methodology
-A transformer model may learn an addition algorithm that differs significantly from the traditional human process.
-Here we define an alternative mathematical framework for addition, showing it is equivalent to the traditional process and so is mathematically correct.
-Then we show that our model implements this alternative framework.
+Transformer models may learn addition algorithms different from traditional human methods.
+We define an alternative, mathematically-equivalent framework for addition and demonstrate our model implements this approach.
 3.1 Mathematical Framework
@@
 For n-digits, we use the terms $D_{n}$, $D^{\prime}_{n}$and $A_{n}$.
-First, the model will need to sum pairs of digits. As in Quirke and Barez (2024), we define the “Simple Addition" sub-task $A_{n}.SA$, which naively calculates the sum of pairs of question digits, as:
+First, we adopt the framework from Quirke and Barez (2024) for our model’s addition process. The “Simple Addition" sub-task $A_{n}.SA$, which naively calculates the sum of digit pairs, is defined as:
 $A_{n}.SA=(D_{n}+D^{\prime}_{n})\mod 10$ (1)
-When there is no carry bit from the previous digit $A_{n}=A_{n}.SA$.
-The model will also need learn how to calculate whether the addition creates a carry bit.
-As in Quirke and Barez (2024), we define the "Simple Carry" sub-task $A_{n}.SC$:
+When there is no carry bit from the previous digit $A_{n}=A_{n}.SA$. The “Simple Carry" sub-task $A_{n}.SC$ determines whether the addition creates a carry bit:
 $A_{n}.SC=\begin{cases}1&\text{if }(D_{n}+D^{\prime}_{n})\geq 10,\\ 0&\text{otherwise}.\end{cases}$ (2)
-The model in Quirke and Barez (2024) is limited to the framework described so far.
-While it is capable of handling simple addition and carry bits generated directly from the addition of digit pairs, it encounters difficulties when faced with problems involving ’cascading carry’ bits, wherein a carry bit from one digit position propagates to the next.
-To illustrate this limitation, consider the problem "00144+00056=000210".
-When adding the digits in the tens position, 4+5, no carry bit is generated directly.
-However, due to the carry bit propagating from the ones position, a carry bit is ultimately required in the tens position.
-A model that solely calculates the sums of digit pairs and considers carry bits resulting from the direct addition of those pairs will fail to accurately resolve this type of problem.
-In such cases, it would yield an incorrect result, such as "144+56=110".
-The 1-layer model could cascade the carry bit up to two digits across, but not more.
+While Quirke et al.’s model is capable of handling simple addition and carry bits generated directly from digit pair addition, it encounters difficulties with ’cascading carry’ bits, where a carry bit from one digit position propagates to the next.
+Consider “00144+00056=000210".
+Adding 4+5 in the tens position doesn’t generate a carry bit directly, but a carry bit propagates from the ones position.
+A model only summing digit pairs and their direct carry bits would fail, producing an incorrect result like “00144+50006=000110".
+The Quirke et al.
+1-layer model could cascade carry bits across two digits, but not three or more.
 3.2 Extending the Mathematical Framework
-To answer "44444+55556=" with "100000", an accurate model must predict the first answer digit A5 as "1".
+To answer “44444+55556=" with “100000", an accurate model must predict the first answer digit A5 as “1".
 To do so, an accurate model must implement a “carry one cascade" circuit, which combines the “carry one" information from all five digits.
@@
 Figure 3 diagrams how our model’s algorithm uses SA, SC, $ST$ and $SV$ to perform addition.
-This algorithm is like the 99% accurate Quirke and Barez (2024) algorithm but contains an additional circuit (the shaded boxes) to calculate “cascading carry one" data to predict with 99.9999% accuracy.
+This algorithm is like the 99% accurate Quirke et al.
+algorithm but contains an additional circuit (the shaded boxes) to calculate “cascading carry one" data to predict with 99.9999% accuracy.
 Figure 3: To predict with 99.9999% accuracy, the addition algorithm first calculates “carry one" values $(A_{n}.ST)$, combining them into “cascading carry one" values $(A_{n}.SV).$ At the “+" token, $A_{n-1}.SV$ gives the first answer digit as 1 or 0.
@@
     Question Complexity.
-We categorized questions by the complexity of the computation required to solve the question (Refer App.
+We categorized questions by computational complexity (App.
 G).
-The addition categories (named S0 to S4) reflect the number of sequential digits that a “carry one" cascades through.
-The subtraction categories (named M0 to M4) reflect the number of sequential digits that a “borrow one" cascades through.
-We analysed which nodes were needed for correct prediction of each category.
+Addition categories (S0-S4) and subtraction categories (M0-M4) reflect the number of sequential digits a “carry one" or “borrow one" cascades through, respectively.
+We analyzed which nodes were necessary for correct predictions in each category.
 4 Experiments
 4.1 Training a Five-Digit Addition Model
-The Quirke and Barez (2024) 5-digit 1-layer addition model achieved an accuracy of $\sim$ 99%.
+The Quirke et al.
+5-digit 1-layer addition model achieved an accuracy of $\sim$ 99%.
 Our experiments suggested that a 2-layer, 3-head model was the smallest configuration capable of achieving 99.9999% accuracy (see App.
@@
 6).
-4.2 Investigating Five-Digit Addition
-Ablation experiments targeting the nodes revealed that the model depends only on nodes located in nine token positions (Figs 6 and 7).
-Further ablation experiments show that for these nine token positions, model uses 36 nodes in predictions.
-The effects of node ablation on our complexity and answer-impact metrics were analyzed (see Figs 5 and Table  1), providing insight into the specific computations performed at each node.
-For each answer digit $A_{n}$, using test questions corresponding to the ST8, ST9 and ST10 categories, we performed PCA on the nodes yielding interpretable results.
-Specifically nine “node and answer-digit” combinations (see Figs 8 and  9) showed strong clustering of the questions aligned to the ST8, ST9 and ST10 categories.
 (P6) (P9) (P10) (P11) (P12) (P13) (P14) (P15) (P16) (P17)
@@
 L1MLP A1 A0
-Table 1: The nodes used in predictions across token positions.
+Table 1: For a sample model, all nodes used in predictions are shown by token position (horizontally) and model layer (vertically), detailing the answer digits they impact.
+Here, the attention heads in token position P10 labelled A5..3 help predict the answer digits A3, A4 and A5.
 For all addition and mixed models studied, before the "=" token, each node often calculates data used to predict multiple answer digits.
 After the "=" token, all nodes in a given token position are used to predict a single answer digit.
-For this sample model, the attention heads in token position P10 labelled A5..3 help predict the answer digits A3, A4 and A5.
+4.2 Investigating Five-Digit Addition
+Ablation experiments targeting the nodes revealed that the model depends only on nodes located in nine token positions (Figs 6 and 7).
+Further ablation experiments show that for these nine token positions, the model uses 36 nodes in predictions.
+The effects of node ablation on our complexity and answer-impact metrics were analyzed (see Figs 5 and Table  1), providing insight into the specific computations performed at each node.
+For each answer digit $A_{n}$, using test questions corresponding to the ST8, ST9 and ST10 categories, we performed PCA on the nodes yielding interpretable results.
+Specifically nine “node and answer-digit” combinations (see Figs 8 and  9) showed strong clustering of the questions aligned to the ST8, ST9 and ST10 categories.
 The algorithm predicts the first answer digit, A5, at position P11.
@@
 We posit our MLP implements the TriAdd function using bigrams and trigrams to calculate $A_{n}.SV$ values from $A_{n}.ST$ values.
-For a specific 5-digit addition model instance, we mapped the algorithm to individual nodes.
-Using ablation intervention to override each node, we verified each node’s role in executing the expected algorithmic step.
-Figure 3 diagrams the addition algorithm, while App.
-J details each nodes purpose for this model instance.
-The model adheres to all known constraints.
-Empirically, it achieves six nines accuracy.
-In summary, we concluded that this specific instance of the 5-digit addition model documented above has six nines accuracy, is well-understood and well-functioning and hence known-good.
+For a specific 5-digit addition model instance, we mapped the algorithm to individual nodes and verified each node’s role using ablation intervention.
+Figure 3 diagrams the algorithm, with node details in App.
+J.
+The model adheres to all known constraints and achieves six nines accuracy.
+We concluded this model instance is well-understood, well-functioning, and hence known-good.
 The 1-layer model uses 21 nodes to achieve two nines (99%) accuracy. This model uses 36 nodes (an increase of 71%) to achieve six nines accuracy.
-4.3 Training n-digit Addition models
-To investigate whether this algorithm is used widely, we first trained seven 2-layer addition models with 5-, 6- and 10-digits, using different seeds, a different optimizer, and changing the answer format to include a sign token (e.g.
-111111+222222=+0333333).
-These seven models all have very low loss (e.g. 1.5e-8) and six nines accuracy. (Details in Tab.6.)
-4.4 Investigating n-digit Addition models
-We implemented a declarative way to describe each sub-task.
-For example, a sub-task can be described as primarily attending to question digits D2 and D’2, influencing answer digit A3, impacting S0 but not S1 complexity questions, with a specific PCA result, and a specific ablation test.
-Using the declaration, we searched for and identified nodes performing all sub-tasks in all the models.
-An algorithm hypothesis (e.g the one described in section 4.2) is then described in terms of the sub-tasks that must exist in the model and their inter-relationships.
-For example, our addition algorithm description states (in part) that the model must implement $A_{n}.ST$ sub-tasks for each question digit before the “=” token.
-We implemented an n-digit algorithm hypothesis declaration framework and a way to test the declaration against the sub-tasks found in each model and map the results.
-(Table 2 shows an example map.)
 (P11) (P12) (P13) (P14) (P15) (P16) (P17) (P18) (P19) (P20)
@@
 Interestingly, here each SA sub-task is shared across two attention heads.
-We found that our seven 2-layer addition models all implement our addition algorithm.
-Given these seven models have six nines accuracy, and all implement the same valid algorithm, we can state these models are known-good.
+4.3 Training n-digit Addition models
+To investigate whether this algorithm is used widely, we first trained seven 2-layer addition models with 5-, 6- and 10-digits, using different seeds, a different optimizer, and changing the answer format to include a sign token (e.g.
+111111+222222=+0333333).
+These seven models all have very low loss (e.g. 1.5e-8) and six nines accuracy. (Details in Tab.6.)
+4.4 Investigating n-digit Addition models
+We developed a declarative method to outline each sub-task.
+For instance, a sub-task may focus on question digits D2 and D’2, affect answer digit A3, influence S0 but not S1 complexity questions, and have specific PCA results and ablation tests.
+Using this declaration, we identified nodes performing these sub-tasks across all models.
+An algorithm hypothesis, such as the one in section 4.2, is described by the required sub-tasks and their relationships.
+For example, our addition algorithm specifies that the model must execute $A_{n}.ST$ sub-tasks for each question digit before the “=” token.
+We created a framework for declaring n-digit algorithm hypotheses and testing them against the sub-tasks found in each model, mapping the results (see Table 2 for an example).
+Our seven 2-layer addition models all implement our addition algorithm. Given their six nines accuracy and implementation of the same algorithm, we can confirm these models as known-good.
 4.5 Training n-digit Mixed models
@@
 Used Inserted
-Question class No % No %
+Question class # % # %
 All questions 96 48
@@
 Table 3: Mixed models re-use most inserted addition-model nodes.
-Many inserted nodes become polysemantic during training.
-That is they perform addition (S), positive-answer subtraction (M) and negative-answer subtraction (N) sub-tasks simultaneously.
+Many inserted nodes become polysemantic during training - performing addition, positive-answer subtraction and negative-answer subtraction sub-tasks simultaneously.
 For a sample mixed model that uses 96 nodes and had 48 nodes inserted, this table shows inserted node reuse.
@@
 4.8 Mixed model summary
-The 7 mixed experimental models all have five or six nines accuracy.
-They all contain the same sub-tasks - suggesting they all implement the same algorithm.
-As we don’t fully understand the algorithm yet, we can’t assert that these mixed models are known-good.
+The 7 mixed experimental models achieve five or six nines accuracy and contain the same sub-tasks, implying a common algorithm.
+However, without fully understanding this algorithm, we can’t confirm these models as known-good.
 5 Conclusion
-In this paper, we successfully trained 5-, 6- and 10-digit, 2-layer, 3-head addition models that we have verified as known-good.
-They all implement the same algorithm, albeit with some implementation variations.
-A significant aspect of our work was demonstrating component reuse.
-We achieved this by integrating an existing addition model into a new larger “mixed” model designed for both addition and subtraction tasks.
-This mixed model has six nines accuracy.
-This integration allowed us to leverage our in-depth understanding of the addition model’s algorithm to unravel the workings of the mixed model’s algorithm.
-Interestingly, the mixed model re-uses most of the inserted addition model nodes, and “upgrades” most of them to become polysemantic - performing the inserted addition subtask and the corresponding subtraction subtasks simultaneously.
-This paper lends support to the assertion by (Michaud et al., 2023) that many natural prediction problems can be broken down into a finite set of “quanta” computations.
-These quanta are essential for models to learn to minimize loss.
-Furthermore, it aligns with the concept that comprehending a network’s functionality is fundamentally about identifying and understanding its sub-quanta.
+We successfully trained and verified known-good 5-, 6- and 10-digit, 2-layer, 3-head addition models implementing the same algorithm with minor variations.
+We demonstrated component reuse by integrating an existing addition model into a larger “mixed" model for both addition and subtraction, achieving six nines accuracy.
+This integration helped us understand the mixed model’s algorithm.
+The mixed model reuses most inserted addition nodes, upgrading many to become polysemantic - performing both addition and subtraction subtasks simultaneously.
+Our work supports (Michaud et al., 2023)’s assertion that many natural prediction problems can be broken into finite "quanta" computations essential for loss minimization.
+It also aligns with the idea that understanding a network’s functionality involves identifying and comprehending its sub-quanta.
 6 Limitations and Future Work
-Our research opens up several promising avenues for future exploration.
-A key objective is to complete the reverse-engineering of the mixed model’s algorithm.
-We need to understand how the polysemantic nodes and the OPR+SGN node outputs are represented in the residual stream and show the MLP layers combine this information to select the appropriate question class data to accurately predict the answer digit.
-Additionally, we aim to develop a comprehensive model that proficiently handles integer n-digit addition, subtraction, and multiplication, ensuring its status as known-good.
-We believe this approach can be re-used to create known-good models in fields like logical reasoning (for AND, OR, Converse, Inverse etc tasks) and planning.
-Using a smaller model to improve an existing Large Language Model (LLM) trustworthiness and capabilities is an active area of research.
-Approaches include composing multiple models (Bansal et al., 2024), LM up-scaling which essentially emulates fine-tuning a large pre-trained model using a small model (Mitchell et al., 2023), and this paper’s insertion of an accurate model into an untrained model.
-Other research (Voita et al., 2023; Hu et al., 2021) indicates the presence of “spare” neurons in LLMs and the potential for integrating small-scale modifications to fix erroneous circuits in LLMs.
-Further research into the benefits of using known-good models to increase LLM trustworthiness and capabilities is justified.
-Moreover, developing methods to incorporate compact known-good models into LLMs, whether untrained or trained, could democratize AI Safety research.
-It enables small teams to focus on specific functional areas, crafting component models to a defined quality standard for integration into various LLMs.
-Our automated framework, for discovering the location of algorithm sub-tasks in models, sped up our research.
-Our framework can be re-used to ensure that known-good components retain their functionality when integrated into LLMs.
-Some of our automated framework is specific to our maths models - much is generic and reuseable on other types of models.
-Our automation focuses more on attention heads and less on MLP layers.
-Extending our framework to integrate the output data from complimentary frameworks (e.g.
-ACDC’s node connectivity strength data (Conmy et al., 2023)) would provide more automated insights into a model’s behavior, further speeding up research.
-We made progress on a declarative language to describe an algorithm in terms of the number and types sub-tasks that must exist in a model for the algorithm to feasibly explain the model’s implementation.
-We made progress on a framework to automatically test an algorithm described in this way against a specific model.
-Further work is required to fully realize its potential.
+Our research suggests several future directions.
+A key goal is to complete reverse-engineering the mixed model’s algorithm, understanding polysemantic nodes and OPR+SGN node output representations in the residual stream, and how MLP layers use this to predict answer digits accurately.
+We aim to develop a comprehensive, known-good model for n-digit addition, subtraction, and multiplication.
+Our approach also could be applied to create known-good models in logical reasoning and planning.
+Using smaller models to enhance LLM trustworthiness and capabilities is an active research area, including model composition (Bansal et al., 2024), LM up-scaling which essentially emulates fine-tuning a large pre-trained model using a small model (Mitchell et al., 2023), and inserting accurate models into untrained ones.
+Research on “spare” neurons in LLMs ((Voita et al., 2023; Hu et al., 2021)) suggests potential for small-scale modifications to fix erroneous circuits.
+Further exploration of using known-good models to improve LLMs is warranted.
+Developing methods to incorporate compact known-good models into LLMs could democratize AI Safety research, allowing small teams to focus on specific areas and create quality components to improve LLMs.
+Our automated framework for discovering algorithm sub-tasks in models accelerated our research and can test whether known-good components retain functionality when integrated into LLMs.
+While some aspects are specific to our math models, much is generic and reusable.
+Extending the framework to integrate output from complementary tools (e.g.
+ACDC’s node connectivity strength data (Conmy et al., 2023)) would provide more automated insights.
+We progressed on a declarative language to describe algorithms in terms of necessary sub-tasks and a framework to test these descriptions against specific models.
+Further work is needed to fully realize its potential.
 7 Impact Statement
@@
 Twenty-one experiment models were trained and analyzed (refer Tab. 6). The models and output of analysis are available on HuggingFace at https://huggingface.co/PhilipQuirke/VerifiedArithmetic
-Num Digits Num Layers Num Heads Training Steps Seed Addition / Sub Accuracy Training loss Heads used MLPs used
+Num Digits Num Layers Num Heads Train Steps Train Seed Train loss Addition Fails / M Subtract Fails / M Heads used MLPs used
 Addition models
-5 1 3 30K 372001 two nines 9.4e-2 15 6
-5 2 3 15K 372001 six nines 1.6e-8 30 16
-5 2 3 40K 372001 six nines 2.0e-9 22 15
-6 2 3 15K 372001 six nines 1.7e-8 31 17
-6 2 3 20K 173289 six nines 1.5e-8 28 17
-6 2 3 20K 572091 six nines 7.0e-9 35 17
-6 2 3 40K 372001 six nines 2.0e-9 29 17
-10 2 3 20K 572091 six nines 7.0e-9 44 28
+5 1 3 30K 372001 9.4e-2 12621 N/A 15 6
+5 2 3 15K 372001 1.6e-8 0 N/A 30 16
+5 2 3 40K 372001 2.0e-9 0 N/A 22 15
+6 2 3 15K 372001 1.7e-8 2 N/A 31 17
+6 2 3 20K 173289 1.5e-8 0 N/A 28 17
+6 2 3 20K 572091 7.0e-9 0 N/A 35 17
+6 2 3 40K 372001 2.0e-9 0 N/A 29 17
+10 2 3 40K 572091 7.0e-9 0 N/A 44 28
 Subtraction models
-6 2 3 30K 372001 six nines 5.8e-6 40 21
-10 2 3 75K 173289 two nines 2.0e-3 101 37
+6 2 3 30K 372001 5.8e-6 N/A 0 40 21
+10 2 3 75K 173289 2.0e-3 N/A 6672 101 37
 Mixed models
-6 3 4 40K 372001 six / six nines 5.0e-9 54 26
-10 3 4 75K 173289 five / two nines 1.1e-6 143 53
+6 3 4 40K 372001 5.0e-9 1 0 54 26
+10 3 4 75K 173289 1.1e-6 2 295 143 53
 Mixed models initialized with addition model
-6 2 3 40K 572091 six / five nines 2.4e-8 57 21
-6 3 3 40K 572091 six / five nines 1.8e-8 70 35
-6 3 3 80K 572091 six / five nines 1.6e-8 75 35
-6 3 4 40K 372001 six / six nines 8.0e-9 72 26
-6 3 4 40K 173289 five / five nines 1.4e-8 60 29
-6 3 4 50K 572091 six / five nines 2.9e-8 79 29
-10 3 3 50K 572091 five / five nines 6.3e-7 90 45
+6 2 3 40K 572091 2.4e-8 0 5 57 21
+6 3 3 40K 572091 1.8e-8 0 3 70 35
+6 3 3 80K 572091 1.6e-8 0 3 75 35
+6 3 4 40K 372001 8.0e-9 0 0 72 26
+6 3 4 40K 173289 1.4e-8 3 2 60 29
+6 3 4 50K 572091 2.9e-8 0 4 79 29
+10 3 3 50K 572091 6.3e-7 6 7 90 45
 Mixed models initialized with add model. Reset useful heads every 100 steps
-6 4 4 40K 372001 five / five nines 1.7e-8 51 30
+6 4 4 40K 372001 1.7e-8 3 8 51 30
 Mixed models initialized with add model. Reset useful heads & MLPs every 100 steps
-6 4 3 40K 372001 four / two nines 3.0e-4 115 53
-Table 6: Main experimental models studied
+6 4 3 40K 372001 3.0e-4 17 3120 115 53
+Table 6: Main experimental models studied. The number of addition and subtraction failures per million questions is shown. The best 5-, 6- and 10-digit models are bolded.
 For each model the ’VerifiedArithmeticTrain’ Colab notebook generates two files:

### v9

@@
-Increasing Trust in Language Models through the Reuse of Verified Circuits
+Arithmetic in Transformers Explained
 Abstract
-Language Models (LMs) are increasingly used for a wide range of prediction tasks, but their training can often neglect rare edge cases, reducing their reliability.
-Here, we define a stringent standard of trustworthiness whereby the task algorithm and circuit implementation must be verified, accounting for edge cases, with no known failure modes.
-We show that a model can be trained to meet this standard if built using mathematically and logically specified frameworks.
-In this paper, we fully verify an auto-regressive transformer model that performs n-digit integer addition.
-To exhibit the reusability of verified modules, we insert the trained integer addition model into a larger untrained model and train the combined model to perform both addition and subtraction.
-We find extensive reuse of the addition circuits for both tasks, easing verification of the more complex subtractor model.
-We discuss how inserting verified task modules into LMs can leverage model reuse to improve verifiability and trustworthiness of LMs built using them.
-The reuse of verified circuits reduces the effort to verify more complex composite models which we believe to be a significant step towards safety and interpretability of LMs.
-Increasing Trust in Language Models through the Reuse of Verified Circuits
-Philip Quirkethanks:    Core research contributor., Clement Neo†‡, Fazl Barezfootnotemark: †‡ Apart Research, University of Oxford
+While recent work has shown transformers can learn addition, previous models exhibit poor prediction accuracy and are limited to small numbers.
+Furthermore, the relationship between single-task and multitask arithmetic capabilities remains unexplored.
+In this work, we analyze 44 autoregressive transformer models trained on addition, subtraction, or both.
+These include 16 addition-only models, 2 subtraction-only models, 8 “mixed” models trained to perform addition and subtraction, and 14 mixed models initialized with parameters from an addition-only model.
+The models span 5- to 15-digit questions, 2 to 4 attention heads, and 2 to 3 layers.
+We show that the addition models converge on a common logical algorithm, with most models achieving >99.999% prediction accuracy.
+We provide a detailed mechanistic explanation of how this algorithm is implemented within the network architecture.
+Subtraction-only models have lower accuracy.
+With the initialized mixed models, through parameter transfer experiments, we explore how multitask learning dynamics evolve, revealing that some features originally specialized for addition become polysemantic, serving both operations, and boosting subtraction accuracy.
+We explain the mixed algorithm mechanically.
+Finally, we introduce a reusable library of mechanistic interpretability tools to define, locate, and visualize these algorithmic circuits across multiple models.
+Machine Learning, ICML
+11footnotetext: Corresponding author: philipquirkenz@gmail.com
 1 Introduction
-Figure 1: An overview of our methodology: (1) We trained an accurate 6-digit integer addition model.
+Figure 1: An overview of our methodology: (1) We trained an accurate 5-digit integer addition model.
 (2) We reverse-engineered the model to find the algorithms that were implemented to perform addition.
@@
 Transformer-based large language models (LLMs) are powerful (Barak et al., 2022) yet largely inscrutable due to their complex, nonlinear interactions in dense layers within high-dimensional spaces.
-Given this complexity, their deployment in critical settings (Zhang et al., 2022) highlights the need for understanding their behavior.
-Hendrycks and Mazeika (2022) argue that making these models interpretable is key to their safe use.
-Mechanistic interpretability focuses on demystifying and validating the algorithms behind model weights, translating complex computations into more human-understandable components (Raukur et al., 2022).
-This understanding aids in predicting model behavior in new situations and fixing model errors.
-In creating and training a model, we aim for high accuracy and trustworthiness.
-We achieve this by holding the model to a standard we term known-good.
-We define a model performing a task to be known-good if:
-  1. 1.
-    The model’s algorithm for the task and the mechanisms it implements (the “circuits") are understood.
-  2. 2.
-    All possible task edge cases have been identified and tested.
-  3. 3.
-    Empirically, the model prediction accuracy is 99.9999% (a standard reliability measure used in industry and abbreviated as “six nines").
-That is, it can perform the task one million times with at most 1 wrong prediction.
-Exhaustive testing of a model task may be infeasible.
-For instance, when adding two 5-digit integers (e.g.
-12345+67890) there are ten billion variations.
-Some tasks can be conceptualized within an existing formal framework that allows identification of all edge cases the model must handle.
-For example, in 5-digit addition, the most uncommon edge case is 55555+44445=100000, which requires a carry bit to cascade through all digits, occurring in only 0.002% of cases.
-A known-good model must incorporate algorithms to manage all known edge cases.
-A known-good model must have verified [^1] In this paper, ’verified’ has the mechanistic interpretability meaning that a specific group of interconnected neurons within a neural network reliably and causally contributes to the model’s output in a meaningful, understandable way, with supporting empirical evidence.
-circuits that perform the task accurately.
-In this paper, we detail the development and interpretation of a known-good model for addition.
-Our findings indicate that the model constructs a specific circuit for each edge case, with these circuits sharing intermediate results.
-We confirm the validity of the entire set of circuits, ensuring they cover all identified edge cases.
-The model achieves a very low training loss and has six nines (99.9999%) accuracy.
-The model hence achieves our known-good standard.
-Additionally, we develop a “mixed" model capable of both addition and subtraction, incorporating the known-good addition model.
-This mixed model has six nines accuracy, and extensively reuses the addition circuits for both operations, facilitating the interpretation of the model’s algorithm.
-We make progress toward a known-good model for both addition and subtraction.
-Hence, our main contributions are three-fold:
-  • Defining several known-good n-digit addition models with six nines accuracy which all use the same algorithm.
-  • Demonstrating a proof of concept for re-using a known good model in the training of another larger, more-capable model, simplifying the interpretation of the new model’s algorithm.
-  • Defining several n-digit addition and subtraction models with six nines accuracy, that reuse established addition model circuits for both operations, and detailing progress towards these models being known-good.
+Given this complexity, their deployment in critical settings (Zhang et al., 2022) highlights the need to understand their behavior.
+Hendrycks & Mazeika (2022) argue that making these models interpretable is key to their safe use.
+Previous work by (Quirke & Barez, 2024) demonstrated that transformers can learn addition on small numbers, but their models had poor accuracy and were limited to a single operation.
+Understanding how transformers can overcome these limitations to perform more accurate mathematical computations across multiple operations is crucial for several reasons.
+First, it provides insights into how neural networks can learn and combine multiple precise algorithms, a key capability for more complex reasoning tasks.
+Second, studying how models transition from single-task to multitask competence while maintaining accuracy reveals important principles about knowledge transfer and representation sharing in neural networks.
+Understanding how models arrive at their predictions is critical for diagnosing errors, enhancing safety, and gaining fundamental insights into modern neural networks.
+Mechanistic interpretability provides a pathway toward this understanding by identifying the “circuits” responsible for particular computations within a network’s layers and attention heads.
+In this work, we use multidigit arithmetic as a revealing testbed for mechanistic interpretability.
+We train and analyze 44 autoregressive transformer models on addition, subtraction, or both, covering question lengths of 5 to 15 digits and varying in their number of layers (2–3) and attention heads (2–4).
+Our experiments encompass three categories of models: addition-only, subtraction-only, and “mixed” models trained to handle both operations.
+Through a systematic investigation of these networks, we show that addition-only models converge on a highly accurate, shared algorithm, yielding near-perfect performance.
+We mechanistically explain the algorithm.
+Subtraction-only models, in contrast, exhibit lower accuracy.
+To explore how multitask learning reshapes these internal algorithms, we then conduct parameter transfer experiments using three distinct approaches, each involving the initialization of mixed models with parameters from addition-only networks.
+Interestingly, one of these approaches boosts the subtraction accuracy beyond what is achieved by mixed models trained from scratch, indicating that features specialized for addition can be effectively repurposed for subtraction.
+We provide a detailed mechanistic explanation of the mixed algorithm, illustrating how certain addition circuits become “polysemantic” under multitask learning, driving behavior for both operations.
+A further contribution of our study is the development of a reusable library of mechanistic interpretability tools.
+This library enables researchers to define and search for algorithmic features, visualize network structures and feature interactions, and conduct cross-model comparisons efficiently.
+Overall, our results demonstrate the power of mechanistic interpretability in shedding light on how transformers learn and reuse algorithmic circuits across related tasks, paving the way for deeper understanding and more reliable deployment of large language models in diverse domains.
 2 Related Work
@@
 Causal mediation analysis (Stolfo et al., 2023) has been used to investigate how large pre-trained LMs like Pythia and GPT-J performed addition to solve word problems.
-It is also possible to improve a model’s arithmetic abilities used supervised fine tuning - including enriched training data (Liu and Low, 2023).
+It is also possible to improve a model’s arithmetic abilities used supervised fine tuning - including enriched training data (Liu & Low, 2023).
 Studying Toy Models for Arithmetic.
-Doing mechanistic interpretability on toy transformers can help to better isolate clear, distinct circuits given the highly specific experimental setup for the model studied (Nanda et al., 2023).
-Quirke and Barez (2024) detailed a 1-layer, 3-head transformer model that performs 5-digit addition, showing it failed on rare edge cases (e.g.
-“77778+22222=100000" where a “carry 1" cascades through 4 digits), highlighting the importance of understanding and testing all edge cases for trustworthiness.
-Many natural prediction problems decompose into a finite set of knowledge and skills that are “quantized" into discrete chunks (quanta) (Michaud et al., 2023).
+Applying mechanistic interpretability to toy transformers can help to better isolate clear, distinct circuits given the highly specific experimental setup for the model studied (Nanda et al., 2023).
+Quirke & Barez (2024) detailed a 1-layer, 3-head transformer model that performs 5-digit addition, showing it failed on edge cases (e.g.
+“77778+22222=100000” where a “carry 1” cascades through 4 digits), incorrectly predicting 1% of questions.
+Many natural prediction problems decompose into a finite set of knowledge and skills that are “quantized” into discrete chunks (quanta) (Michaud et al., 2023).
 Models must learn these quanta to reduce loss.
@@
 3 Methodology
-Transformer models may learn addition algorithms different from traditional human methods.
-We define an alternative, mathematically-equivalent framework for addition and demonstrate our model implements this approach.
+Transformer models learn addition algorithms different from traditional human methods.
+We define an alternative, mathematically-equivalent framework for addition and demonstrate our models implements this approach.
 3.1 Mathematical Framework
 Consider the task of adding two $n$-digit numbers together.
-We define the first number as $D=\{D_{n-1},D_{n-2},\ldots,D_{0}\}$ and the second number as $D^{\prime}=\{D^{\prime}_{n-1},D^{\prime}_{n-2},\ldots,D^{\prime}_{0}\}$ and the answer as $A=\{A_{n},A_{n-1},\ldots,A_{0}\}$.
+We define the first number (or addend) as $D=\{D_{n-1},D_{n-2},\ldots,D_{0}\}$ and the second number as $D^{\prime}=\{D^{\prime}_{n-1},D^{\prime}_{n-2},\ldots,D^{\prime}_{0}\}$ and the answer (or sum) as $A=\{A_{n},A_{n-1},\ldots,A_{0}\}$.
 Figure 2 shows an illustrative example.
@@
 For n-digits, we use the terms $D_{n}$, $D^{\prime}_{n}$and $A_{n}$.
-First, we adopt the framework from Quirke and Barez (2024) for our model’s addition process. The “Simple Addition" sub-task $A_{n}.SA$, which naively calculates the sum of digit pairs, is defined as:
+We adopt the framework from Quirke & Barez (2024) for our model’s addition process, where they break down the task of multi-digit addition into an addition subtask and a carry subtask.
+Firstly, the “Simple Addition” subtask $A_{n}.SA$ naively calculates the sum of corresponding digit pairs in the addends:
 $A_{n}.SA=(D_{n}+D^{\prime}_{n})\mod 10$ (1)
-When there is no carry bit from the previous digit $A_{n}=A_{n}.SA$. The “Simple Carry" sub-task $A_{n}.SC$ determines whether the addition creates a carry bit:
+The “Simple Carry” subtask $A_{n}.SC$ then determines whether the addition between the two digits creates a carry bit:
 $A_{n}.SC=\begin{cases}1&\text{if }(D_{n}+D^{\prime}_{n})\geq 10,\\ 0&\text{otherwise}.\end{cases}$ (2)
-While Quirke et al.’s model is capable of handling simple addition and carry bits generated directly from digit pair addition, it encounters difficulties with ’cascading carry’ bits, where a carry bit from one digit position propagates to the next.
-Consider “00144+00056=000210".
+By combining these two subtasks, the model and framework from Quirke & Barez (2024) is capable of handling simple addition and carry bits generated directly from digit pair addition.
+However, it struggles with problems with ’cascading carry’ bits, where a carry bit from one digit position propagates to the next.
+Consider “00144+00056=000210”.
 Adding 4+5 in the tens position doesn’t generate a carry bit directly, but a carry bit propagates from the ones position.
-A model only summing digit pairs and their direct carry bits would fail, producing an incorrect result like “00144+50006=000110".
-The Quirke et al.
-1-layer model could cascade carry bits across two digits, but not three or more.
+A model only summing digit pairs and their direct carry bits would fail, producing an incorrect result like “00144+50006=000110”.
+The Quirke & Barez (2024) 1-layer model could cascade carry bits across two digits, but not three or more.
 3.2 Extending the Mathematical Framework
-To answer “44444+55556=" with “100000", an accurate model must predict the first answer digit A5 as “1".
-To do so, an accurate model must implement a “carry one cascade" circuit, which combines the “carry one" information from all five digits.
-This is especially hard as the model predicts answer tokens from left to right.
-We introduce a digit-level sub-task called TriCase that calculates the essential “carry one" information for a single pair of digits $D_{n}$ and $D^{\prime}_{n}$.
-TriCase has 3 possible outputs representing a definite carry one (ST10), a possible carry one depending on the results of other calculations (ST9), or definitely not a carry one (ST8):
-$\begin{split}A_{n}.ST&=\underbrace{\text{TriCase}}_{(D_{n},D^{\prime}_{n})}=\\ &\begin{cases}ST10&\text{if }(D_{n}+D^{\prime}_{n})\geq 10,\\ ST9&\text{if }(D_{n}+D^{\prime}_{n})=9,\\ ST8&\text{if }(D_{n}+D^{\prime}_{n})\leq 8,\\ \end{cases}\end{split}$ (3)
-To perform the “cascading carry one" calculation, we introduce a TriAdd sub-task.
-It handles the case where a possible carry one becomes a definite carry one because the next lower digit pair generated a carry one.
-TriAdd is defined as:
+To solve $n$-digit addition with high accuracy, we need to extend the basic framework to handle carry bits that cascade through multiple digits.
+Consider the problem ”44444+55556=100000”.
+For an autoregressive model to correctly predict the first answer digit as ”1”, it needs to track how the carry bit cascades from the rightmost digit all the way to the leftmost position in a single forward pass.
+This is particularly challenging for an autoregressive model that predicts answer tokens from left to right.
+To solve this, we extend the mathematical framework with two key subtasks.
+First, the digit-level subtask ”Sum Type” $A_{n}.ST$ analyzes each pair of digits and classifies their sum as either definitely causing a carry, possibly causing a carry, or definitely not causing a carry.
+This is implemented by a TriCase operation that maps each digit pair ($D_{n}$ and $D^{\prime}_{n}$), TriCase into one of three possible states:
+$\begin{split}A_{n}.ST&=\underbrace{\text{TriCase}}_{(D_{n},D^{\prime}_{n})}=\\ &\begin{cases}ST10&\text{if }(D_{n}+D^{\prime}_{n})\geq 10\text{ (definite carry)}\\ ST9&\text{if }(D_{n}+D^{\prime}_{n})=9\text{ (possible carry)}\\ ST8&\text{if }(D_{n}+D^{\prime}_{n})\leq 8\text{ (no carry)}\\ \end{cases}\end{split}$ (3)
+The key insight is that while some digit pairs definitely generate a carry (like 5+6), and others definitely don’t (like 2+3), pairs that sum to 9 (like 4+5) may or may not generate a carry depending on whether there’s an incoming carry bit from the next lower digit.
+Then, we introduce the ”Sum Validation” subtask $A_{n}.SV$ that processes these carry states sequentially to determine whether any ”possible carries” become ”definite carries”.
+This subtask is implemented using a TriAdd operation that examines each digit’s sum type along with the sum type of the next lower digit:
 $\begin{split}A_{n}.SV&=\underbrace{\text{TriAdd}}_{(A_{n}.ST,A_{n-1}.ST)}=\\ &\begin{cases}ST10&\text{if }A_{n}.ST=ST10\text{ or }\\ &\quad(A_{n}.ST=ST9\text{ and }\\ &\quad A_{n-1}.ST=ST10),\\ ST8&\text{otherwise.}\end{cases}\end{split}$ (4)
-The model can use $ST$ and $SV$ to accurately calculate A5 as 1 or 0 by using A5 = A4.SV = TriAdd(A4.ST, TriAdd(A3.ST, TriAdd(A2.ST, TriAdd(A1.ST, A0.ST)))).
-Note that in calculating A5, the model has also calculated an accurate carry bit for each answer digit.
-For example, the carry bit for A4 is A3.SV = TriAdd(A3.ST, TriAdd(A2.ST, TriAdd(A1.ST, A0.ST))).
-With this framework, the model only needs the sub-tasks $A_{n}$.SA, $A_{n}.ST$ and $A_{n}.SV$ to accurately perform addition (but some models also use the redundant $A_{n}$.SC sub-task).
-This framework, if implemented by a model, is sufficient for the model to perform n-digit addition accurately.
-Figure 3 diagrams how our model’s algorithm uses SA, SC, $ST$ and $SV$ to perform addition.
-This algorithm is like the 99% accurate Quirke et al.
-algorithm but contains an additional circuit (the shaded boxes) to calculate “cascading carry one" data to predict with 99.9999% accuracy.
-Figure 3: To predict with 99.9999% accuracy, the addition algorithm first calculates “carry one" values $(A_{n}.ST)$, combining them into “cascading carry one" values $(A_{n}.SV).$ At the “+" token, $A_{n-1}.SV$ gives the first answer digit as 1 or 0.
-The other answer digits are calculated by combining “base add" $(A_{n}.BA)$ and “carry one" $(A_{n}.SC)$ calculations with the pre-calculated $A_{n}.SV$ values.
+Revisiting the problem ”44444+55556=100000”, the model can accurately determine whether the first answer digit A5 should be 1 or 0 by examining all digit pairs in sequence.
+This chain of operations uses TriAdd to process the Sum Types from least to most significant digits:
+$A1.SV$ $=\text{TriAdd}(A1.ST,A0.ST)$
+$A2.SV$ $=\text{TriAdd}(A2.ST,A1.SV)$
+$A3.SV$ $=\text{TriAdd}(A3.ST,A2.SV)$
+$A4.SV$ $=\text{TriAdd}(A4.ST,A3.SV)$
+$A5$ $=(A4.SV=ST10)$
+Most importantly, by using this cascading calculation to calculating A5, the model also simulatenously calculates the accurate carry bits for each answer digit.
+The carry bits are simply the SV values we computed along the way—we already have A3.SV for A4’s carry bit, A2.SV for A3’s carry bit, and so on.
+Within our extended framework, the model only requires the subtasks $A_{n}$.SA, $A_{n}.ST$ and $A_{n}.SV$ to accurately perform $n$-digit addition accurately.
+Some models may also use the redundant $A_{n}$.SC subtask.
+Figure 3 shows how our model’s algorithm uses SA, SC, $ST$ and $SV$ to perform addition.
+This algorithm is like the 99% accurate Quirke & Barez (2024) algorithm, but contains an additional circuit (the shaded boxes) to calculate “cascading carry one” data to solve $n$-digit addition with >99.999% accuracy.
+Figure 3: To predict with >99.999% accuracy, the addition algorithm first calculates “carry one” values $(A_{n}.ST)$, combining them into “cascading carry one” values $(A_{n}.SV).$ At the “+” token, $A_{n-1}.SV$ gives the first answer digit as 1 or 0.
+The other answer digits are calculated by combining “base add” $(A_{n}.BA)$ and “carry one” $(A_{n}.SC)$ calculations with the pre-calculated $A_{n}.SV$ values.
 3.3 Techniques
-In investigating the model circuits, we want to understand what each attention head or MLP layer is doing across each token position, and how it relates to our mathematical framework.
-Hence, we define a node as the computation done by an attention head or MLP layer for a given token position.
-To investigate what each node is doing, we use the following techniques:
+To undetsand how the model implements arithmetic operations, we analyzed both attention heads and MLP layers at each token position.
+We define a node as either an attention head or MLP layer at a specific token position (e.g., the first attention head in the first layer processing the third digit).
+We employed four techniques to understand each node’s function:
   1. 1.
     Intervention Ablation.
-To find out how the model depends on the output of a node, we replace the output of that node with the vector that is the mean of all of its outputs across a batch and measure how that impacts downstream performance.
-We also use (automated, n-digit) intervention ablation tests, specific to each sub-task, to test for the expected sub-task behavior.
+To determine the role of a node’s output, we replace its output with the mean vector of its outputs across a batch, then measured how this impacted downstream predictions.
+For each subtask, we developed automated intervention ablation tests across $n$-digits to verify whether a node implemented that subtask.
   2. 2.
-    Attention Patterns. To find out what the model attends to at a node, we take the attention pattern at that token position and take the significant tokens attended to ($>0.01$ post softmax).
+    Attention Patterns.
+For each attention head node, we examined which input tokens received significant attention (defined as post-softmax attention score $>0.01$).
+This revealed which parts of the input each attention head considered relevant.
   3. 3.
-    Principal Component Analysis (PCA). We use PCA to investigate the outputs of attention heads, especially where our framework suggests the head output may be tri-state or bi-state.
+    Principal Component Analysis (PCA).
+We applied PCA to node outputs, especially when investigating attention heads that our framework suggests might have descrete states (e.g.
+tri-state or bi-state outputs).
   4. 4.
@@
 G).
-Addition categories (S0-S4) and subtraction categories (M0-M4) reflect the number of sequential digits a “carry one" or “borrow one" cascades through, respectively.
+For addition, categories S0-S4 reflect how many sequential digits a “carry one” operation cascades through.
+For subtraction, categories M0-M4 reflect a similar cascading of ”borrow one” operations.
 We analyzed which nodes were necessary for correct predictions in each category.
@@
 4.1 Training a Five-Digit Addition Model
-The Quirke et al.
-5-digit 1-layer addition model achieved an accuracy of $\sim$ 99%.
-Our experiments suggested that a 2-layer, 3-head model was the smallest configuration capable of achieving 99.9999% accuracy (see App.
-D for alternatives tested).
-This configuration effectively doubled the computational power compared to the 1-layer model (see App.
-B for mode model configuration details).
-Moreover, a 2-layer model introduces the capability to “compose" the attention heads in novel ways, facilitating the implementation of more complex algorithms (Elhage et al., 2021b).
-We trained a 5-digit, 2-layer, 3-head model, with a 14 token vocabulary ( 0, .., 9, +, -, =, *, / ), batch size of 64, learning rate of 0.00008 and weight decay of 0.1.
-Training used an infinite dataset enriched with rare edge cases.
-Loss was defined as the mean across all answer tokens of their negative log likelihood loss.
-After 30 thousand training batches, the model’s final training loss was $\sim 2.3\times 10^{-8}$.
-Testing showed this model has six nines accuracy.
-(More details in App.
+Previous work by Quirke & Barez (2024) achieved $\sim$ 99% accuracy on 5-digit addition using a 1-layer addition model.
+To improve upon this, we systematically explored different model architectures and found that a transformer with 2 layers and 3 attention heads per layer was the minimal configuration needed to achieve >99.999% accuracy (see App.
+D for our exploration of alternative architectures, and App.
+B for model configuration details.).
+This 2-layer architecture doubles the model’s computational capacity and enables cross-layer attention composition for more sophiscated algorithms (Elhage et al., 2021b).
+We trained our model to perform 5-digit addition using an architecture with 2 layers and 3 attention heads per layer.
+The model’s vocabulary consisted of 14 tokens: the digits 0-9 plus the mathematical operators +, -, =, *, and /.
+We used a batch size of 64, learning rate of 0.00008, and weight decay of 0.1.
+To ensure robust performance on edge cases, we trained on an infinite dataset that was deliberately enriched with rare numerical combinations.
+The loss function was defined as the mean negative log likelihood across all output tokens.
+After 30,000 training batches, the model achieved a final training loss of approximately $2.3\times 10^{-8}$, and subsequent testing confirmed accuracy exceeding 99.999%.
+For comprehensive training details and results, see App.
 C and Tab.
-6).
+6.
+Table 1: For a sample model, all nodes used in predictions are shown by token position (horizontally) and model layer (vertically), detailing the answer digits they impact.
+Here, the attention heads in token position P10 labelled A5..3 help predict the answer digits A3, A4 and A5.
+For all addition and mixed models studied, before the ”=” token, each node often calculates data used to predict multiple answer digits.
+After the ”=” token, all nodes in a given token position are used to predict a single answer digit.
 (P6) (P9) (P10) (P11) (P12) (P13) (P14) (P15) (P16) (P17)
@@
 L1MLP A1 A0
-Table 1: For a sample model, all nodes used in predictions are shown by token position (horizontally) and model layer (vertically), detailing the answer digits they impact.
-Here, the attention heads in token position P10 labelled A5..3 help predict the answer digits A3, A4 and A5.
-For all addition and mixed models studied, before the "=" token, each node often calculates data used to predict multiple answer digits.
-After the "=" token, all nodes in a given token position are used to predict a single answer digit.
 4.2 Investigating Five-Digit Addition
-Ablation experiments targeting the nodes revealed that the model depends only on nodes located in nine token positions (Figs 6 and 7).
-Further ablation experiments show that for these nine token positions, the model uses 36 nodes in predictions.
-The effects of node ablation on our complexity and answer-impact metrics were analyzed (see Figs 5 and Table  1), providing insight into the specific computations performed at each node.
-For each answer digit $A_{n}$, using test questions corresponding to the ST8, ST9 and ST10 categories, we performed PCA on the nodes yielding interpretable results.
-Specifically nine “node and answer-digit” combinations (see Figs 8 and  9) showed strong clustering of the questions aligned to the ST8, ST9 and ST10 categories.
-The algorithm predicts the first answer digit, A5, at position P11.
-A5, which is always 0 or 1, is the most challenging to predict as it may rely on a long carry one cascade (e.g.
-55555+44445=100000).
-An accurate algorithm must compute this cascade using the nodes located in positions P8 to P11.
-As illustrated in Figure 7, these nodes attend to all digit pairs from D4 D’4 to D0 D’0.
-Additionally, the PCA data, as shown in Figures 8 and 9, suggest that these nodes produce tri-state outputs.
-After our first two algorithm hypothesises failed testing (see App.
-H and I), we discovered that the model utilizes a minimal set of “carry one” information, leading to the development of the TriCase quanta.
-The model performs $A_{n}.ST$ using bigrams (see App.
-F) to map two input tokens to one result token e.g.
-“6” + “7” = ST10.
-In positions P8 to P11, the model does $A_{n}.ST$ calculations on all digit pairs from D4 D’4 to D0 D’0.
-An MLP layer can be thought of as a “key-value pair" memory (Meng et al., 2022; Geva et al., 2021) that can hold many bigrams and trigrams.
-We posit our MLP implements the TriAdd function using bigrams and trigrams to calculate $A_{n}.SV$ values from $A_{n}.ST$ values.
+Through ablation experiments, we found that the model relies on just 36 nodes distributed across nine token positions (Figs 7 and 8).
+We analyzed how these nodes affect computation complexity and answer accuracy through ablation studies (see Figs 6 and Table 1).
+Core Algorithm Discovery.
+The model’s most challenging task is predicting A5, the first answer digit, at position P11.
+This digit, which can only be 0 or 1, depends on potential carry-bit cascades through multiple positions (e.g., in 55555+44445=100000).
+After two unsuccessful algorithm hypotheses (detailed in App.H andI), we discovered the model’s fundamental mechanism: it uses a minimal representation of carry information through what we call the TriCase quanta.
+The model computes $A_{n}.ST$ using bigram operations that map digit pairs to result tokens (e.g., ”6” + ”7” = ST10) in positions P8 to P11.
 For a specific 5-digit addition model instance, we mapped the algorithm to individual nodes and verified each node’s role using ablation intervention.
@@
 J.
-The model adheres to all known constraints and achieves six nines accuracy.
-We concluded this model instance is well-understood, well-functioning, and hence known-good.
-The 1-layer model uses 21 nodes to achieve two nines (99%) accuracy. This model uses 36 nodes (an increase of 71%) to achieve six nines accuracy.
+The model adheres to all known constraints and achieves >99.999% accuracy using 36 nodes, as compared to the 1-layer model uses 21 nodes to achieve 99% accuracy.
+Table 2: All addition models studied implement our addition algorithm.
+The algorithm SA, SC and $ST$ subtasks all exist for each digit and in appropriate token positions.
+For a sample model, this map shows the subtask locations.
+Interestingly, here each SA subtask is shared across two attention heads.
 (P11) (P12) (P13) (P14) (P15) (P16) (P17) (P18) (P19) (P20)
@@
 L0H2 A5.ST
-Table 2: All addition models studied implement our addition algorithm.
-The algorithm SA, SC and $ST$ sub-tasks all exist for each digit and in appropriate token positions.
-For a sample model, this map shows the subtask locations.
-Interestingly, here each SA sub-task is shared across two attention heads.
-4.3 Training n-digit Addition models
-To investigate whether this algorithm is used widely, we first trained seven 2-layer addition models with 5-, 6- and 10-digits, using different seeds, a different optimizer, and changing the answer format to include a sign token (e.g.
+4.3 Investigating n-digit Addition models
+To test whether this algorithm generalises beyond 5-digit addition, we conducted a comprehensive study of larger addition models.
+We trained 15 2-layer transformer models to handle addition problems ranging from 5- to 15-digits.
+These models used varying random seeds and optimizers, and included a sign token in their output format (e.g.
 111111+222222=+0333333).
-These seven models all have very low loss (e.g. 1.5e-8) and six nines accuracy. (Details in Tab.6.)
-4.4 Investigating n-digit Addition models
-We developed a declarative method to outline each sub-task.
-For instance, a sub-task may focus on question digits D2 and D’2, affect answer digit A3, influence S0 but not S1 complexity questions, and have specific PCA results and ablation tests.
-Using this declaration, we identified nodes performing these sub-tasks across all models.
-An algorithm hypothesis, such as the one in section 4.2, is described by the required sub-tasks and their relationships.
-For example, our addition algorithm specifies that the model must execute $A_{n}.ST$ sub-tasks for each question digit before the “=” token.
-We created a framework for declaring n-digit algorithm hypotheses and testing them against the sub-tasks found in each model, mapping the results (see Table 2 for an example).
-Our seven 2-layer addition models all implement our addition algorithm. Given their six nines accuracy and implementation of the same algorithm, we can confirm these models as known-good.
+Figure 4: The 5- to 13-digit 2-layer 3-head addition models have very low loss and >99.999% accuracy. With more digits the model loss and accuracy is worse. Details in Tab.6.
+Analysis Framework.
+To systematically analyse these models, we first created a declarative method for specifying each subtask within the addition process.
+Each subtask definition includes (1) the input digits it processes, (2) the answer digit it affects, (3) the complexity level it influences, and (4) expected PCA patterns and ablation test results.
+Using this declaration, we identified the nodes performing these subtasks across all models.
+Since our framework lays out the addition algorithm in terms of its required subtasks., we can also verify whether a given model implements this algorithm by checking for the presence and proper sequencing of all required subtasks.
+For example, our addition algorithm specifies that the model must execute $A_{n}.ST$ subtasks for each question digit before the first answer token.
+Table 2 shows how an addition model fully implements all required subtasks in its nodes.
+Using this framework, we found that all fifteen 2-layer addition models, regardless of their size, implemented the same core algorithm we identified in our 5-digit model.
+This consistent pattern suggests we have identified a robust and generalizable algorithm that naturally arises when transformer models are trained on addition in our setup where the digits are tokenized individually.
 4.4 Training n-digit Mixed models
-To explore reuse, we initialized untrained models with a known-good addition-only model, then trained them to perform both subtraction and addition. We call these “mixed" models.
-Specifically, we trained seven larger (6- or 10-digit, 2- or 3-layer, 3- or 4-head) models after initializing them with the weights from a known-good 2-layer 3-head addition model.
-The first 2 layers, first 3 heads of the mixed model were initialized with the addition model weights.
-We trained the mixed model with 80% subtraction and 20% addition batches.
-We enriched the (infinite) training dataset with rare addition and subtraction edge cases.
-Some models achieved six nines accuracy and the others five nines.
-Attempts to “freeze" the inserted attention heads and/or MLP layers by periodically copying the addition weights back into the mixed model every 100 training steps resulted in lower accuracy.
-(Refer App.M and Tab.6)
-Figure 4: Each mixed model was initialized with the weights from a known-good addition model, then was trained on 80% subtraction and 20% addition batches.
-A sample log loss graph (final loss 8.0e-9) is shown.
+To explore how transformers can learn multiple arithmetic operations simultaneously, we conducted experiments with models capable of both addition and subtraction.
+We were particularly interested in whether knowledge from addition-specialized models could transfer to this expanded task.
+We trained 14 mixed models with various architectures, including 6- and 10-digit models with 2-3 layers and 3-4 attention heads.
+Each model was initialized using weights from a highly accurate addition-only model, specifically copying the weights into the first two layers and first three heads of the mixed model.
+During training, we used a curriculum of 80% subtraction and 20% addition examples, with deliberate enrichment of rare edge cases for both operations.
+This initialization approach proved effective, with most models achieving high accuracy on both operations.
+We explored alternative training strategies, such as periodically ”freezing” the inserted attention heads or MLPs by restoring their original addition-model weights every 100 training steps.
+However, these more constrained approaches resulted in lower accuracy, suggesting that some flexibility in modifying the initialized weights was necessary for optimal performance.
+The initialized mixed experimental models achieve >99.999% accuracy and contain the same subtasks, implying a common algorithm.
+Figure 5: The 5- to 8-digit initialized mixed models have very low loss and >99.999% accuracy. With more digits the model loss and accuracy is worse. Details in Tab.7.
 4.5 Investigating n-digit Mixed models
-Unlike addition, subtraction question answers can be either positive or negative.
-Similar looking positive-answer (e.g.
-10009-10000=+000009) and negative-answer (e.g.
-10009-20000=-009991) subtraction questions can give answers that differ at several digit positions.
-We posited that the model treats three distinct question classes “addition", “positive-answer subtraction" and “negative-answer subtraction" differently.
-Ablation of nodes showed that some nodes only help predict one question class, some help predict two classes and some all three classes.
-It also showed that the inserted addition nodes are heavily used and the majority become polysemantic, performing both addition and subtraction calculations.
-(Details in Tab.3.)
+The mixed models revealed fascinating patterns in how they handled different types of arithmetic problems.
+Our hypothesis was that the model handled three distinct question classes: addition, positive-answer subtraction (e.g., 10009-10000=+000009), and negative-answer subtraction (e.g., 10009-20000=-009991).
+Through ablation studies, we found that many nodes specialized in handling specific question classes, while others became ”polysemantic,” capable of processing multiple classes simultaneously (More details in Tab.3).
+Most notably, the majority of nodes initialized from the addition model became polysemantic during training.
+Rather than developing entirely new circuits for subtraction, the models adapted their addition circuits to handle both operations.
+Table 3: Mixed models re-use most inserted addition-model nodes.
+Many inserted nodes become polysemantic during training - performing addition, positive-answer subtraction and negative-answer subtraction subtasks simultaneously.
+For a sample mixed model that uses 96 nodes and had 48 nodes inserted, this table shows inserted node reuse.
 Used Inserted
@@
 Negative-answer sub 53 55% 29 60%
-Table 3: Mixed models re-use most inserted addition-model nodes.
-Many inserted nodes become polysemantic during training - performing addition, positive-answer subtraction and negative-answer subtraction sub-tasks simultaneously.
-For a sample mixed model that uses 96 nodes and had 48 nodes inserted, this table shows inserted node reuse.
-To investigate the subtraction algorithm, and paralleling addition algorithm subtasks, we defined subtraction-specific sub-tasks Base Diff and Borrow One (see Tab.
-4) and complexity measures (see App.
+Defining Subtraction Subtasks.
+Subtraction requires two distinct computational paths based on the relative magnitude of the numbers.
+When D $\geq$ D’, as in 325 - 123 = +202, we perform direct subtraction with borrowing.
+When D $<$ D’, as in 325 - 456 = -131, we must instead compute the magnitude of the difference and mark it negative.
+Hence, this computational difference necessitates separate frameworks for positive-answer (M) and negative-answer (N) subtraction, each with parallel operations to addition’s Base Add ($SA$) and Carry One ($SC$), which we define in Tab.
+4.
+We also define complexity measures for subtraction (see App.
 G and  N).
+Table 4: We define 2 “positive-answer subtraction” and 2 “negative-answer subtraction” subtasks that parallel the addition subtasks, Base Add SA and Carry One SC .
 Name +ve Sub -ve Sub Definition
@@
 Borrow One MB NB $D_{n}$- $D^{\prime}_{n}$$<0$
-Table 4: We define 2 “positive-answer subtraction" and 2 “negative-answer subtraction" sub-tasks that parallel the addition sub-tasks Base Add SA and Carry One SC
-We found that the inserted nodes that performed SA in the addition model, perform SA, MD and ND in the mixed model.
-(Refer Table 5.) These three sub-tasks are similar in that each performs a mapping from 100 input cases (10 $D_{n}$x 10 $D^{\prime}_{n}$) to 10 output cases (0..9).
-The mixed model “upgraded" these nodes to be polysemantic during training.
-Similarly, some SC addition nodes became polysemantic and now process SC, MB and/or NB .
+Just as addition must handle ”cascading carry” cases (e.g., 99999+1=100000), subtraction must handle ”cascading borrow” cases.
+Consider 100000-1=099999, where a borrow cascades through multiple zeros.
+To handle such cases efficiently, we define a three-state classification $MT$ that parallels addition’s $ST$ :
+$\begin{split}A_{n}.MT&=\underbrace{\text{TriCase}}_{(D_{n},D^{\prime}_{n})}=\\ &\begin{cases}MTN&\text{if }D_{n}<D^{\prime}_{n}\\ MT0&\text{if }D_{n}=D^{\prime}_{n}\\ MT1&\text{if }D_{n}>D^{\prime}_{n}\\ \end{cases}\end{split}$ (5)
+For M questions, MTN is a definite borrow one, MT0 is a possible borrow one (depending on the results of other digit calculations) and MT1 is definitely not a borrow one.
+For N questions, the interpretation is the opposite.
+Finally, paralleling the ”cascading carry one” $An.SV$ operation in addition, we define cascading borrow calculations $A_{n}.MV$ and $A_{n}.NV$ for positive and negative cases.
+Evolution of Polysemantic Nodes.
+During training, the mixed model adapted the nodes it inherited from the addition model.
+Nodes that originally only performed Simple Addition ($SA$) developed the ability to simultaneously handle three distinct operations: addition ($SA$), positive-answer subtraction ($MD$), and negative-answer subtraction ($ND$).
+(Refer to Table 5.) We believe that this adaptation was likely because these three operations share a common structure: each maps a pair of digits (10 possibilities for $D_{n}$× 10 possibilities for $D^{\prime}_{n}$= 100 input cases) to a single-digit result (10 possible outputs).
+The model similarly adapted nodes that originally handled carry bits ($SC$) to also process borrow operations for both positive ($MB$) and negative ($NB$) subtraction cases.
+Table 5: For mixed models, in later tokens, polysemantic attention heads simultaneously generate outputs for the three question classes addition S, M and N.
+Other heads calculate the question class by attending to the question operation (OPR) token and the answer sign (SGN) token.
+The MLP layers then select the output appropriate for the class.
+In this sample map, from P16, the first 3 rows contain many polysemantic nodes, while the 4th row calculates the question class.
 (P9) (P10) (P11) (P12) (P13) (P14) (P15) (P16) (P17) (P18) (P19) (P20)
@@
 SGN SGN SGN SGN SGN SGN
-Table 5: For mixed models, in later tokens, polysemantic attention heads simultaneously generate outputs for the three question classes addition S, M and N.
-Other heads calculate the question class by attending to the question operation (OPR) token and the answer sign (SGN) token.
-The MLP layers then select the output appropriate for the class.
-In this sample map, from P16, the first 3 rows contain many polysemantic nodes, while the 4th row calculates the question class.
-An accurate subtraction model must answer “cascading borrow one" questions like 100000-000001=+0099999.
-We define the “essence of borrow one" subtask $MT$ used by both positive-answer (M) and negative answer (N) subtraction questions.
-$MT$ is like addition’s $ST$ subtask:
-$\begin{split}A_{n}.MT&=\underbrace{\text{TriCase}}_{(D_{n},D^{\prime}_{n})}=\\ &\begin{cases}MTN&\text{if }D_{n}<D^{\prime}_{n}\\ MT0&\text{if }D_{n}=D^{\prime}_{n}\\ MT1&\text{if }D_{n}>D^{\prime}_{n}\\ \end{cases}\end{split}$ (5)
-For M questions, MTN is a definite borrow one, MT0 is a possible borrow one (depending on the results of other digit calculations) and MT1 is definitely not a borrow one.
-For N questions, the interpretation is the opposite.
-Paralleling addition’s “cascading carry one" $A_{n}.SV$ calculation, we define “cascading borrow one" calculation sub-tasks $A_{n}.MV$ for positive-answer subtraction and $A_{n}.NV$ for negative-answer subtraction.
-We posit the MLP implements TriAdd-like functions using bigrams and trigrams to calculate $A_{n}.MV$ and $A_{n}.NV$ values from $A_{n}.MT$ values.
 4.6 Mixed model question class detection
-We posit that there is a specific circuit to detect whether a question is in the S, M or N class.
-If the question operator is “+" then the class is S, but if the question operator is “-" then the model must calculate if $D>=D^{\prime}$ to distinguish between the M and N classes.
-For accuracy, it needs this answer by the “=" token to predict the first answer token (the answer sign) as “+" or “-".
-$D>=D^{\prime}$ can be derived from the $A_{n}.MT$ data. Alternatively, this calculation could be a distinct circuit using a new sub-task we define as:
+Mixed models must identify whether each input belongs to one of three classes: addition (S), positive-answer subtraction (M), or negative-answer subtraction (N).
+For addition questions (marked by ”+”), classification is immediate.
+For subtraction questions (marked by ”-”), the model must determine whether D $\geq$ D’ to distinguish between M and N classes.
+This classification must occur by the ”=” token since it determines whether the first answer token will be ”+” or ”-”.
+The model can determine D $\geq$ D’ in two ways.
+First, it can use the $A_{n}.MT$ values already computed for subtraction.
+Alternatively, it can use a dedicated comparison circuit based on a new subtask:
 $A_{n}.GT=\begin{cases}1&\text{if }D_{n}\geq D^{\prime}_{n},\\ 0&\text{otherwise}.\end{cases}$ (6)
-For a 4 digit question, $D>=D^{\prime}$ can be calculated as A3.GT=1 or (A3.GT=0 and (A2.GT=1 or (A2.GT=0 and (A1.GT=1 or (A1.GT=0 and A0.GT=1))).
-Our test for GT is that ablation causes the answer to change sign.
-The $MT$ calculation is very similar.
-We found $A_{n}.MT$ and $A_{n}.GT$ subtasks in all mixed models.
-Usually (but not always) both subtasks are calculated by the same node.
-(Table 2 shows an example.) As both the $MT$ and GT approaches are valid, models can learn valid algorithm sub-task implementations that differ per answer digit.
-We further posit that a node that implements say A2.SA, A2.MD and A2.ND does not know whether it is dealing with a S, M or N question, so it outputs all three possible answers to the residual stream.
-Another node must calculate the S, M or N distinction likely by attending to the question operator (OPR) and the answer sign (SGN).
-We define attention sub-tasks OPR and SGN for these calculations.
-Table 2 shows that for each answer digit one attention head attends to both OPR and SGN.
-We believe these heads transfer sufficient data to the residual stream to allow the MLP layer(s) to calculate the question class, and so select the appropriate output from the polysemantic A2.SA/MD/ND node.
-4.8 Mixed model summary
-The 7 mixed experimental models achieve five or six nines accuracy and contain the same sub-tasks, implying a common algorithm.
-However, without fully understanding this algorithm, we can’t confirm these models as known-good.
+For a 4-digit number, this comparison cascades through the digits: D $\geq$ D’ is true if either $A3.GT=1$, or if $A3.GT=0$ and $A2.GT=1$, and so on through $A0.GT$.
+We verify this circuit through ablation tests that show its disruption causes answer signs to flip.
+We found all mixed models incorporate the $A_{n}.MT$ and $A_{n}.GT$ subtasks to determine whether the answer is postiive or negative.
+Typically (but not always), the same node computes both subtasks (see Table 2).
+Since both the $MT$ and GT approaches can validly distinguish between positive and negative outcomes, the model sometimes chooses one approach for some digits and the other approach for others.
+A single node can implement multiple arithmetic operations (e.g., $A2.SA$, $A2.MD$, $A2.ND$) but has no intrinsic way of knowing whether the current question is addition, positive subtraction, or negative subtraction.
+Hence, we hypothesise that it outputs all three possible results into the residual stream.
+A different node must then select the correct result by identifying the question’s operation and required sign.
+In practice, we observe that one or more attention heads attend to the question operator token (OPR) and the initial answer sign token (SGN), encoding the necessary information for the downstream MLP layers to pick which of the three outputs is correct.
+As Table 2 shows, each answer digit typically has one attention head that focuses on both the OPR and SGN tokens, thereby enabling the MLP to filter out the irrelevant outputs and retain the correct one.
 5 Conclusion
-We successfully trained and verified known-good 5-, 6- and 10-digit, 2-layer, 3-head addition models implementing the same algorithm with minor variations.
-We demonstrated component reuse by integrating an existing addition model into a larger “mixed" model for both addition and subtraction, achieving six nines accuracy.
-This integration helped us understand the mixed model’s algorithm.
-The mixed model reuses most inserted addition nodes, upgrading many to become polysemantic - performing both addition and subtraction subtasks simultaneously.
-Our work supports (Michaud et al., 2023)’s assertion that many natural prediction problems can be broken into finite "quanta" computations essential for loss minimization.
-It also aligns with the idea that understanding a network’s functionality involves identifying and comprehending its sub-quanta.
-6 Limitations and Future Work
-Our research suggests several future directions.
-A key goal is to complete reverse-engineering the mixed model’s algorithm, understanding polysemantic nodes and OPR+SGN node output representations in the residual stream, and how MLP layers use this to predict answer digits accurately.
-We aim to develop a comprehensive, known-good model for n-digit addition, subtraction, and multiplication.
-Our approach also could be applied to create known-good models in logical reasoning and planning.
-Using smaller models to enhance LLM trustworthiness and capabilities is an active research area, including model composition (Bansal et al., 2024), LM up-scaling which essentially emulates fine-tuning a large pre-trained model using a small model (Mitchell et al., 2023), and inserting accurate models into untrained ones.
-Research on “spare” neurons in LLMs ((Voita et al., 2023; Hu et al., 2021)) suggests potential for small-scale modifications to fix erroneous circuits.
-Further exploration of using known-good models to improve LLMs is warranted.
-Developing methods to incorporate compact known-good models into LLMs could democratize AI Safety research, allowing small teams to focus on specific areas and create quality components to improve LLMs.
-Our automated framework for discovering algorithm sub-tasks in models accelerated our research and can test whether known-good components retain functionality when integrated into LLMs.
-While some aspects are specific to our math models, much is generic and reusable.
-Extending the framework to integrate output from complementary tools (e.g.
-ACDC’s node connectivity strength data (Conmy et al., 2023)) would provide more automated insights.
-We progressed on a declarative language to describe algorithms in terms of necessary sub-tasks and a framework to test these descriptions against specific models.
-Further work is needed to fully realize its potential.
-7 Impact Statement
+Our analysis of 44 transformer models demonstrates their ability to learn highly accurate mathematical algorithms, achieving over 99.999% accuracy on addition tasks up to 13 digits.
+By investigating how models implement computations using attention heads and MLPs, we uncovered novel computational approaches distinct from traditional human calculation methods.
+Parameter transfer experiments revealed that features originally specialized for addition could be repurposed for subtraction, with many nodes becoming polysemantic.
+Our research provides insights into how transformers learn precise algorithmic computations across related tasks, developing a reusable library of mechanistic interpretability tools that advances our understanding of neural network computational learning.
+Impact Statement
 Our work aims to explain the inner workings of transformer-based language models, which may have broad implications for a wide range of applications.
@@
   • A : Answer to the question (including answer sign)
-  • An : nth token in the answer. Zero-based. A0 is the units value. The highest token is the “+" or "-" answer sign
+  • An : nth token in the answer. Zero-based. A0 is the units value. The highest token is the “+” or ”-” answer sign
   • S : Prefix for Addition. Think S for Sum. Aka ADD.
-  • SA : Basic Add. An addition sub-task. An.SA is defined as (Dn + D’n) % 10. e.g. 5 + 7 gives 2
-  • SC : Carry One. An addition sub-task. An.SC is defined as Dn + D’n >=>=> = 10. e.g. 5 + 7 gives True
-  • SS : Make Sum 9. An addition sub-task. An.SS is defined as Dn + D’n == 9. e.g. 5 + 7 gives False
-  • ST : TriCase. An addition sub-task. Refer paper 2 for details
-  • ST8, ST9, ST10 : Outputs of the ST TriCase sub-task.
+  • SA : Basic Add. An addition subtask. An.SA is defined as (Dn + D’n) % 10. e.g. 5 + 7 gives 2
+  • SC : Carry One. An addition subtask. An.SC is defined as Dn + D’n >=>=> = 10. e.g. 5 + 7 gives True
+  • SS : Make Sum 9. An addition subtask. An.SS is defined as Dn + D’n == 9. e.g. 5 + 7 gives False
+  • ST : TriCase. An addition subtask. Refer paper 2 for details
+  • ST8, ST9, ST10 : Outputs of the ST TriCase subtask.
   • M : Prefix for Subtraction with a positive answer. Think M for Minus. Aka SUB
-  • : Basic Difference. A subtraction sub-task. An.MD is defined as (Dn - D’n) % 10. e.g. 3 - 7 gives 6
-  • : Borrow One. A positive-answer subtraction sub-task. An.MB is defined as Dn - D’n $<$ 0. e.g. 5 - 7 gives True
-  • MZ : Make Zero. A positive-answer subtraction sub-task. An.MZ is defined as Dn - D’n == 0. e.g. 5 - 5 gives True
-  • MT : TriCase. A positive-answer subtraction sub-task.
-  • MT1, MT0, MT-1 : Outputs of the MT TriCase sub-task.
+  • : Basic Difference. A subtraction subtask. An.MD is defined as (Dn - D’n) % 10. e.g. 3 - 7 gives 6
+  • : Borrow One. A positive-answer subtraction subtask. An.MB is defined as Dn - D’n $<$ 0. e.g. 5 - 7 gives True
+  • MZ : Make Zero. A positive-answer subtraction subtask. An.MZ is defined as Dn - D’n == 0. e.g. 5 - 5 gives True
+  • MT : TriCase. A positive-answer subtraction subtask.
+  • MT1, MT0, MT-1 : Outputs of the MT TriCase subtask.
   • N : Prefix for Subtraction with a negative answer. Think N for Negative. Aka NEG
-  • ND : Basic Difference. A negative-answer subtraction sub-task. An.ND is defined as (Dn - D’n) % 10. e.g. 3 - 7 gives 6
-  • NB : Borrow One. A negative-answer subtraction sub-task. An.NB is defined as Dn - D’n $<$ 0. e.g. 5 - 7 gives True
-  • NZ : Make Zero. A negative-answer subtraction sub-task. An.NZ is defined as Dn - D’n == 0. e.g. 5 - 5 gives True
-  • NT : TriCase. A negative-answer subtraction sub-task.
-  • GT : Greater Than. A (positive-answer or negative-answer) subtraction sub-task. An.GT is defined as Dn $>$ D’n. e.g. 3 $>$ 5 gives False
-  • OPR : Operator. A sub-task that attends to the + or - token in the question (which determines whether the question is addition or subtraction).
-  • SGN : Sign. A sub-task that attends to the first answer token, which is + or -
+  • ND : Basic Difference. A negative-answer subtraction subtask. An.ND is defined as (Dn - D’n) % 10. e.g. 3 - 7 gives 6
+  • NB : Borrow One. A negative-answer subtraction subtask. An.NB is defined as Dn - D’n $<$ 0. e.g. 5 - 7 gives True
+  • NZ : Make Zero. A negative-answer subtraction subtask. An.NZ is defined as Dn - D’n == 0. e.g. 5 - 5 gives True
+  • NT : TriCase. A negative-answer subtraction subtask.
+  • GT : Greater Than. A (positive-answer or negative-answer) subtraction subtask. An.GT is defined as Dn $>$ D’n. e.g. 3 $>$ 5 gives False
+  • OPR : Operator. A subtask that attends to the + or - token in the question (which determines whether the question is addition or subtraction).
+  • SGN : Sign. A subtask that attends to the first answer token, which is + or -
   • PCA : Principal Component Analysis
@@
 The Colab runs on a T4 GPU.
-Training takes up to 35 mins.
+Each training run takes up to 60 mins.
 The key parameters (and their common configurations) are:
@@
 Each digit is represented as a separate token.
-(Liu and Low, 2023) state that LLaMa’s “remarkable arithmetic ability … is mainly atributed to LLaMA’s consistent tokenization of numbers".
+(Liu & Low, 2023) state that LLaMa’s “remarkable arithmetic ability … is mainly atributed to LLaMA’s consistent tokenization of numbers”.
 The model’s vocabulary contains 14 tokens ( 0, .., 9, +, -, =, *, / ) to enable this and planned future investigations.
@@
 Validation test data covering all edge cases was manually constructed. These test cases are not used during training.
+The Colabs will be made available on publication.
 Appendix C Appendix: Model Loss
@@
   • All Digits Loss: For “all answer digits” graphs and analysis, the loss used is the mean of the “per digit” loss across all the answer digits.
-In our experimental models, the number of digits in the question varies from 5 to 10, the number of layers varies from 1 to 4, the number of heads varies from 3 to 4.
-Each experimental model’s loss is detailed in Tab.
-6.
+In our experimental models, the number of digits in the question varies from 5 to 13, the number of layers varies from 1 to 4, the number of heads varies from 3 to 4.
+Each experimental model’s loss is detailed in Tabs.
+6 and  7.
 Appendix D Appendix: Addition Model Shape
@@
 Appendix E Appendix: Experimental models
-Twenty-one experiment models were trained and analyzed (refer Tab. 6). The models and output of analysis are available on HuggingFace at https://huggingface.co/PhilipQuirke/VerifiedArithmetic
+Forty-four models were trained and analyzed (Refer Tabs.
+6 and  7).
+The models and output of analysis are available on HuggingFace at https://huggingface.co/PhilipQuirke in the quanta_maths_* collections to support further research in AI Safety.
+Table 6: Addition-only models studied.
+The number of addition failures per million questions is shown.
+The number of useful attention heads at token position and useful MLP layers at token position are shown - summarizing the data shown in figures like Fig 7.
+Num Digits Num Layers Num Heads Train Steps Train Seed Train loss Addition Fails / M Heads used MLPs used
+5 1 3 30K 372001 9.4e-2 12621 15 6
+5 2 3 15K 372001 1.6e-8 0 30 16
+5 2 3 40K 372001 2.0e-9 0 22 15
+6 2 3 15K 372001 1.7e-8 2 31 17
+6 2 3 20K 173289 1.5e-8 0 28 17
+6 2 3 20K 572091 7.0e-9 0 35 17
+6 2 3 40K 372001 2.0e-9 0 29 17
+7 2 3 45K 173289 3.0e-9 0 31 20
+8 2 3 45K 173289 3.0e-9 0 35 22
+9 2 3 45K 173289 3.0e-9 0 54 27
+10 2 3 40K 572091 7.0e-9 0 44 28
+11 2 3 50K 173289 8.0e-9 2 56 29
+12 2 3 50K 173289 5.0e-9 3 50 33
+13 2 3 50K 173289 6.3e-8 1 66 31
+14 2 3 60K 173289 5.5e-6 199 68 35
+15 2 3 80K 572091 8.6e-8 10 93 58
+Table 7: Subtraction-only and mixed models studied.
+The number of addition and subtraction failures per million questions is shown.
+The number of useful attention heads at token position and useful MLP layers at token position are shown - summarizing the data shown in figures like Fig 7.
 Num Digits Num Layers Num Heads Train Steps Train Seed Train loss Addition Fails / M Subtract Fails / M Heads used MLPs used
-Addition models
-5 1 3 30K 372001 9.4e-2 12621 N/A 15 6
-5 2 3 15K 372001 1.6e-8 0 N/A 30 16
-5 2 3 40K 372001 2.0e-9 0 N/A 22 15
-6 2 3 15K 372001 1.7e-8 2 N/A 31 17
-6 2 3 20K 173289 1.5e-8 0 N/A 28 17
-6 2 3 20K 572091 7.0e-9 0 N/A 35 17
-6 2 3 40K 372001 2.0e-9 0 N/A 29 17
-10 2 3 40K 572091 7.0e-9 0 N/A 44 28
 Subtraction models
@@
 Mixed models
+5 3 4 40K 372001 9.0e-9 0 0 45 21
 6 3 4 40K 372001 5.0e-9 1 0 54 26
+7 3 4 50K 372001 2.0e-8 2 6 108 40
+8 3 4 60K 173289 4.7e-8 0 7 123 45
+9 3 4 60K 173289 3.2e-7 1 33 140 46
 10 3 4 75K 173289 1.1e-6 2 295 143 53
+11 3 4 80K 572091 3.9e-8 0 13 138 50
+12 3 4 85K 572091 1.7e-8 2 10 167 55
+13 3 4 685K 572091 9.5e-6 399 4164 197 64
 Mixed models initialized with addition model
+5 2 3 40K 572091 2.5e-7 1 52 49 20
 6 2 3 40K 572091 2.4e-8 0 5 57 21
@@
 6 3 4 50K 572091 2.9e-8 0 4 79 29
+7 3 4 50K 572091 1.8e-8 4 1 104 38
+8 3 4 70K 572091 4.3e-5 50 1196 116 42
+9 3 4 70K 572091 5.4e-8 1 4 160 50
 10 3 3 50K 572091 6.3e-7 6 7 90 45
+11 3 4 75K 572091 5.9e-5 11066 1120 141 47
 Mixed models initialized with add model. Reset useful heads every 100 steps
@@
 6 4 3 40K 372001 3.0e-4 17 3120 115 53
-Table 6: Main experimental models studied. The number of addition and subtraction failures per million questions is shown. The best 5-, 6- and 10-digit models are bolded.
-For each model the ’VerifiedArithmeticTrain’ Colab notebook generates two files:
-  • A “XXXXXX.pth" file containing the model weights
-  • A “XXXXXX_train.json" file containing configuration information and training loss data
-While, for each model the ’VerifiedArithmeticAnalysis’ Colab notebook generates two more files:
-  • A “XXXXXX_behavior.json" file containing generic “behavior" facts learnt about the model by the Colab e.g. P18L0H0 attends to tokens D3 and D’3
-  • A “XXXXXX_maths.json" file containing “maths-specific" facts learnt about the model by the Colab e.g. P18L0H0 performs the A3.SC sub-task.
+For each model the ’QuantaMathsTrain’ Colab notebook generates two files:
+  • A “model.pth” file containing the model weights
+  • A “training_loss.json” file containing configuration information and training loss data
+While, for each model the ’QuantaMathsAnalysis’ Colab notebook generates two more files:
+  • A “behaviors.json” file containing generic “behavior” facts learnt about the model by the Colab e.g. P18L0H0 attends to tokens D3 and D’3
+  • A “features.json” file containing maths-specific ”feature” facts learnt about the model by the Colab e.g. P18L0H0 performs the A3.SC subtask.
 Appendix F Appendix: TriAdd Implementation
@@
 (Refer Tab.8.)
+Table 8: $A_{n}$.TriAdd can be calculated from $A_{n}.ST$ and $A_{n-1}.ST$ through nine bigram mappings and yielding the three distinct outputs ST8, ST9 and ST10
 $A_{n}.ST$ $A_{n}.ST$ $A_{n}.ST$
@@
  ST10 ST8 * ST10 ST10
-Table 7: $A_{n}$.TriAdd can be calculated from $A_{n}.ST$ and $A_{n-1}.ST$ through nine bigram mappings and yielding the three distinct outputs ST8, ST9 and ST10
 Note that in the case $A_{n}.ST$ = ST9 and $A_{n-1}.ST$ = ST10, the answer is indeterminate.
@@
 To analyze question difficulty, we categorized addition questions by the complexity of the computation required to solve the question, as shown in Tab.
-8.
+9.
 The categories are arranged according to the number of digits that a carry bit has to cascade through.
+Table 9: We categorise addition questions into non-overlapping “calculation complexity” quanta, ordered by increased computational difficulty (and decreasing occurrence frequency).
+Five-digit addition questions quanta are S0 to S5.
+Ten-digit addition question quanta are S0 to S10.
+S10’s frequency is $\sim 3e-4$ showing the need to enrich training data for rare edge cases.
 Name Contains Example Freq
@@
 S5 SA,SCx5 11111+88889=100000 $\sim$2%
-Table 8: We categorise addition questions into non-overlapping “calculation complexity" quanta, ordered by increased computational difficulty (and decreasing occurrence frequency).
-Five-digit addition questions quanta are S0 to S5.
-Ten-digit addition question quanta are S0 to S10.
-S10’s frequency is $\sim 3e-4$ showing the need to enrich training data for rare edge cases.
 Appendix H Appendix: Addition Hypothesis 1
@@
   • Uses the new early positions to (somehow) do the SS calculations with higher accuracy than the 1-layer model.
-  • The long double staircase still finalises each answer digit’s calculation.
+  • The long double staircase still finalizes each answer digit’s calculation.
   • The two attention nodes in the long double staircase positions do the SA and SC calculations and pull in SS information calculated in the early positions.
@@
 The $A_{n}$.T operation does not understand mathematical addition. Tab. 10 shows how the model implements the T operator as a bigram mapping.
+Table 10: Implementing the T operator as a bigram mapping from 2 input tokens to 1 result token.
 $D_{n}$vs $D^{\prime}_{n}$ 0 1 4 5 8 9
@@
 9 9 10 13 14 17 18
-Table 9: Implementing the T operator as a bigram mapping from 2 input tokens to 1 result token.
 $A_{n}$.T is a compact way to store data. Tab. 11 show how, if it needs to, the model can convert a $A_{n}$.T value into a one-digit-accuracy SA, SC or SS value.)
+Table 11: Converting a $A_{n}$.T value into a SA , SC or SS value.
 $A_{n}$.T $A_{n}$.SA $A_{n}$.SC $A_{n}$.SS
@@
 18 8 1 0
-Table 10: Converting a $A_{n}$.T value into a SA , SC or SS value.
-Our notation shorthand for one-digit-accuracy these “conversion" bigram mappings is:
+Our notation shorthand for one-digit-accuracy these “conversion” bigram mappings is:
   • $A_{n}$.SA = ($A_{n}$.T % 10) where % is the modulus operator
@@
 Instead the P14 attention heads calculate A1.ST1 from D1 and D’1 directly, and only relies on the P10.D1.ST2 value in the case where A1.ST2 != A1.ST.
-That is, the calculation is “use P14.A1.ST1 value else use A1.ST2 values".
-This aligns with the model learning the P10.A1.ST calculation early in training (for 90% accuracy) and later learning that P10.A1.ST2 contains additional information it can use to get to six nines accuracy.
+That is, the calculation is “use P14.A1.ST1 value else use A1.ST2 values”.
+This aligns with the model learning the P10.A1.ST calculation early in training (for 90% accuracy) and later learning that P10.A1.ST2 contains additional information it can use to get to >99.999% accuracy.
 Figure 6: For a sample 5-digit 2-layer 3-head addition model, this map shows a compacted view of all useful token positions (horizontally) and all useful attention heads and MLP layers (vertically) used in predictions as green cells.
@@
 Appendix K Appendix: Addition Interchange Interventions
-To test the hypothesis 3 mapping of the mathematical framework (casual abstraction) to the model attention heads, various “interchange interventions" (aka activation patching) experiments were performed on the model, where
+To test the hypothesis 3 mapping of the mathematical framework (casual abstraction) to the model attention heads, various “interchange interventions” (aka activation patching) experiments were performed on the model, where
   • A particular claim about an attention head has selected for testing.
@@
 M4 MD,MBx4 N4 ND,NBx4 S4
-Table 13: We define “positive-answer subtraction" and “negative-answer subtraction" calculation complexity quanta that parallel the addition quanta.
+Table 14: We define “positive-answer subtraction” and “negative-answer subtraction” calculation complexity quanta that parallel the addition quanta.
 Appendix L Appendix: N-Digit Addition
@@
 Appendix M Appendix: Mixed Model Initialization
-We experimented with three approaches to re-using the trained addition model in the “mixed" (addition and subtraction) model:
+We experimented with three approaches to re-using the trained addition model in the “mixed” (addition and subtraction) model:
   • Initialize Only: Initialize the untrained mixed model with the addition model weights before training begins.
-  • Freeze Attention: As per “Initialize Only", but also every 100 training steps recopy the attention head weights from the addition model into the partially-trained mixed model.
-  • Freeze All: As per “Initialize Only", but also every 100 training steps recopy the entire addition model (attention heads and MLP layers) into the partially-trained mixed model.
-Our intuition was that “Initialize Only" would give the mixed model the most freedom to learn new algorithms, but that the “Freeze Attention" and “Freeze All" approaches would make the resulting trained mixed model easier to interpret (as we could reuse our addition model insights).
-After experimentation we found that the “Initialize Only" approach was the only one that quickly trained to be able to do both addition and subtraction accurately.
-We concluded that the other two methods constrainws the model’s ability to learn new algorithms too much.
-We also experimented with “where" in the model we inserted the addition (6-digit, 2-layer, 3-head) model into the slightly larger (6-digit, 3-layer, 4-head) mixed model.
+  • Freeze Attention: As per “Initialize Only”, but also every 100 training steps recopy the attention head weights from the addition model into the partially-trained mixed model.
+  • Freeze All: As per “Initialize Only”, but also every 100 training steps recopy the entire addition model (attention heads and MLP layers) into the partially-trained mixed model.
+Our intuition was that “Initialize Only” would give the mixed model the most freedom to learn new algorithms, but that the “Freeze Attention” and “Freeze All” approaches would make the resulting trained mixed model easier to interpret (as we could reuse our addition model insights).
+After experimentation we found that the “Initialize Only” approach was the only one that quickly trained to be able to do both addition and subtraction accurately.
+We concluded that the other two methods constrain the model’s ability to learn new algorithms too much.
+We also experimented with “where” in the model we inserted the addition (6-digit, 2-layer, 3-head) model into the slightly larger (6-digit, 3-layer, 4-head) mixed model.
 That is, do we initialize the first 2 layers or the last 2 layers of the mixed model?
@@
   • All the notes about the addition model (above) also apply to the mixed model.
-  • The model contains a new sub-task that stands out: The algorithm relies on calculations done at token position P0, when the model has only seen one question token!
+  • The model contains a new subtask that stands out: The algorithm relies on calculations done at token position P0, when the model has only seen one question token!
 What information can the model gather from just the first token?
-Intuitively, if the first token is a “8" or “9" then the first answer token is more likely to be a “+" (and not a “-").
+Intuitively, if the first token is a “8” or “9” then the first answer token is more likely to be a “+” (and not a “-”).
 The model uses this heuristic even though this probabilistic information is sometimes incorrect and so will work against the model achieving very low loss.

### v10

@@
-Arithmetic in Transformers Explained
+Understanding Addition and Subtraction in Transformers
 Abstract
-While recent work has shown transformers can learn addition, previous models exhibit poor prediction accuracy and are limited to small numbers.
-Furthermore, the relationship between single-task and multitask arithmetic capabilities remains unexplored.
-In this work, we analyze 44 autoregressive transformer models trained on addition, subtraction, or both.
-These include 16 addition-only models, 2 subtraction-only models, 8 “mixed” models trained to perform addition and subtraction, and 14 mixed models initialized with parameters from an addition-only model.
-The models span 5- to 15-digit questions, 2 to 4 attention heads, and 2 to 3 layers.
-We show that the addition models converge on a common logical algorithm, with most models achieving >99.999% prediction accuracy.
-We provide a detailed mechanistic explanation of how this algorithm is implemented within the network architecture.
-Subtraction-only models have lower accuracy.
-With the initialized mixed models, through parameter transfer experiments, we explore how multitask learning dynamics evolve, revealing that some features originally specialized for addition become polysemantic, serving both operations, and boosting subtraction accuracy.
-We explain the mixed algorithm mechanically.
-Finally, we introduce a reusable library of mechanistic interpretability tools to define, locate, and visualize these algorithmic circuits across multiple models.
-Machine Learning, ICML
-11footnotetext: Corresponding author: philipquirkenz@gmail.com
+Transformers are widely deployed in large language models (LLMs), yet most models still fail on basic arithmetic tasks such as multidigit addition.
+In contrast, we show that small transformers trained from scratch can solve n-digit addition and subtraction with 99.999% accuracy.
+Building directly on prior work that uncovered addition circuits, we extend the analysis to subtraction and present a unified mechanistic account based on cascading carry and borrow circuits.
+Using a suite of 49 trained models, we apply systematic ablations and node-level constraints to validate the learned mechanisms, and release a reproducible interpretability toolkit for studying arithmetic circuits.
+Finally, surveying 180 publicly available LLMs, we find that only 7% can reliably perform addition, underscoring the gap between specialized small models and general-purpose LLMs.
+Our results show that arithmetic can be implemented exactly by tiny transformers, offering a tractable case study for mechanistic interpretability and a cautionary contrast with the persistent arithmetic failures of much larger models.
 1 Introduction
-Figure 1: An overview of our methodology: (1) We trained an accurate 5-digit integer addition model.
-(2) We reverse-engineered the model to find the algorithms that were implemented to perform addition.
-(3) We inserted the addition model into a new model, by copying the weights of the attention heads and MLPs (in brown) into the larger model during initialization.
-(4) We then train the new model on 80% subtraction and 20% addition questions.
-(5) We find that the resulting model predicts accurately and reuses the inserted addition circuits for both addition and subtraction questions.
-Transformer-based large language models (LLMs) are powerful (Barak et al., 2022) yet largely inscrutable due to their complex, nonlinear interactions in dense layers within high-dimensional spaces.
-Given this complexity, their deployment in critical settings (Zhang et al., 2022) highlights the need to understand their behavior.
-Hendrycks & Mazeika (2022) argue that making these models interpretable is key to their safe use.
-Previous work by (Quirke & Barez, 2024) demonstrated that transformers can learn addition on small numbers, but their models had poor accuracy and were limited to a single operation.
-Understanding how transformers can overcome these limitations to perform more accurate mathematical computations across multiple operations is crucial for several reasons.
-First, it provides insights into how neural networks can learn and combine multiple precise algorithms, a key capability for more complex reasoning tasks.
-Second, studying how models transition from single-task to multitask competence while maintaining accuracy reveals important principles about knowledge transfer and representation sharing in neural networks.
-Understanding how models arrive at their predictions is critical for diagnosing errors, enhancing safety, and gaining fundamental insights into modern neural networks.
-Mechanistic interpretability provides a pathway toward this understanding by identifying the “circuits” responsible for particular computations within a network’s layers and attention heads.
-In this work, we use multidigit arithmetic as a revealing testbed for mechanistic interpretability.
-We train and analyze 44 autoregressive transformer models on addition, subtraction, or both, covering question lengths of 5 to 15 digits and varying in their number of layers (2–3) and attention heads (2–4).
-Our experiments encompass three categories of models: addition-only, subtraction-only, and “mixed” models trained to handle both operations.
-Through a systematic investigation of these networks, we show that addition-only models converge on a highly accurate, shared algorithm, yielding near-perfect performance.
-We mechanistically explain the algorithm.
-Subtraction-only models, in contrast, exhibit lower accuracy.
-To explore how multitask learning reshapes these internal algorithms, we then conduct parameter transfer experiments using three distinct approaches, each involving the initialization of mixed models with parameters from addition-only networks.
-Interestingly, one of these approaches boosts the subtraction accuracy beyond what is achieved by mixed models trained from scratch, indicating that features specialized for addition can be effectively repurposed for subtraction.
-We provide a detailed mechanistic explanation of the mixed algorithm, illustrating how certain addition circuits become “polysemantic” under multitask learning, driving behavior for both operations.
-A further contribution of our study is the development of a reusable library of mechanistic interpretability tools.
-This library enables researchers to define and search for algorithmic features, visualize network structures and feature interactions, and conduct cross-model comparisons efficiently.
-Overall, our results demonstrate the power of mechanistic interpretability in shedding light on how transformers learn and reuse algorithmic circuits across related tasks, paving the way for deeper understanding and more reliable deployment of large language models in diverse domains.
+Large language models (LLMs) achieve remarkable performance across diverse tasks, but still struggle with basic arithmetic.
+In a survey of 180 LLMs ranging from 1B to 405B parameters, 7% can reliably perform addition.
+The few that succeed typically rely on external tools, suggesting that the underlying models have not truly learned arithmetic (See App.
+E).
+In contrast, we show that transformers trained from scratch can master both addition and subtraction with near-perfect accuracy - without using external tools.
+Figure 1: Our n-digit addition algorithm is mathematically sound.
+It uses 4 features: It calculates single-digit carry-one values $(ST_{n}$), combining them into multidigit carry-one values $(SV_{n}$).
+Any $SV_{n}$ uncertain (U) values are refined to 0 or 1 over tokens (highlighted).
+By the “+” token, $SV2$ gives the A3 value as 0 or 1.
+The other answer digits $A_{n}$ are calculated from $SA_{n}$ and $SV_{n-1}$ values.
+We construct (effectively infinite) synthetic training data enriched with the hardest edge cases, and train 49 small transformer models (2–3 layers, 3–4 attention heads) on 5 to 12-digit arithmetic.
+These models converge rapidly, reaching >>99.999% accuracy.
+They succeed on the cascading carry one and borrow one cases where LLMs fail e.g.
+555555555 + 444444448 = 1000000003.
+Our largest model has  10M parameters—over four orders of magnitude smaller than GPT-3—and trains to near-perfect accuracy in under an hour on a single GPU, providing an efficient, reproducible, and interpretable testbed for mechanistic analysis.
+Figure 2: We score the addition and subtraction capability of 180 public LLMs.
+A score of 5 means the LLM handled two 5 digits numbers correctly but failed with 6 digit numbers.
+7% of addition and 12% of subtraction models get the maximum test score 15.
+The top models can call external tools.
+We introduce a novel, simple, mathematically exact algorithm for n-digit addition (Figure 1).
+The algorithm respects the transformer’s left-to-right processing constraint, handling high-value digits before low-value ones.
+We also introduce a novel, exact algorithm for subtraction.
+We then ask whether our trained models actually implement these algorithms.
+To do so, a model must realize each algorithmic subtask.
+We hypothesize the existence of specific calculation “nodes” capable of carrying out these subtasks, and develop code to automatically search for candidate nodes.
+We evaluate these candidates through systematic ablations.
+Each ablation is designed around a specific subtask, allowing us to precisely predict how the model’s output should change when we intervene.
+The predictions match perfectly: intervening on the identified node produces exactly the expected effect.
+Every model we study contains nodes corresponding to all subtasks, though in some cases a single logical role is distributed across two attention heads.
+The algorithms also impose strict constraints on node ordering and relationships.
+For example, certain nodes must appear earlier in the computation sequence, while others depend on them downstream (see Figure 1).
+We test these constraints and find they are consistently satisfied across seeds, model sizes, and digit ranges.
+Together, these results show that all our accurate models converge on the same algorithmic solution.
+To make this analysis reproducible, we release a general-purpose interpretability library.
+https://github.com/PhilipQuirke/quanta_mech_interp.
+It provides tools for (1) node characterization, search, and ablation, (2) systematic testing of node relationships and constraints, and (3) visualizations.
+It is a foundational tool used by this project https://github.com/PhilipQuirke/quanta_maths
+Our contributions are fivefold: (i) a large-scale survey of LLM arithmetic capabilities, (ii) enriched datasets for addition and subtraction, (iii) 49 small transformer models trained to near-perfect accuracy, (iv) human-comprehensible exact addition and subtraction algorithms and (v) a reusable interpretability toolkit demonstrating that these models implement the algorithms.
+We conclude that transformers can implement exact n-digit addition and subtraction through interpretable circuits.
+With appropriate training and data, larger LLMs could in principle adopt the same mechanisms.
 2 Related Work
-Mechanistic interpretability aims to reverse engineer neural networks to find interpretable algorithms that are implemented in a model’s weights (Olah et al., 2020).
-Mathematical frameworks (Elhage et al., 2021a) explain how transformer attention heads can work with each other to implement complex algorithms.
-Causal Scrubbing (Jenner et al., 2023) recommends explaining a model algorithm by documenting a low-level computation graph, mapping from the graph to the model nodes that implement the computation, and performing experimentation verification.
-Investigative techniques such as ablation interventions, activation unembeddings (nostalgebraist, 2020), and sparse autoencoders (Nanda, 2023; Cunningham et al., 2023), underpinned by the more theoretical frameworks (Elhage et al., 2021b; Geva et al., 2022), provide tools to help confirm a mapping.
-Investigating pre-trained LMs on Arithmetic.
-Even though basic arithmetic can be solved following a few simple rules, pre-trained LMs often struggle to solve simple math questions (Hendrycks et al., 2021).
-Causal mediation analysis (Stolfo et al., 2023) has been used to investigate how large pre-trained LMs like Pythia and GPT-J performed addition to solve word problems.
-It is also possible to improve a model’s arithmetic abilities used supervised fine tuning - including enriched training data (Liu & Low, 2023).
-Studying Toy Models for Arithmetic.
-Applying mechanistic interpretability to toy transformers can help to better isolate clear, distinct circuits given the highly specific experimental setup for the model studied (Nanda et al., 2023).
-Quirke & Barez (2024) detailed a 1-layer, 3-head transformer model that performs 5-digit addition, showing it failed on edge cases (e.g.
-“77778+22222=100000” where a “carry 1” cascades through 4 digits), incorrectly predicting 1% of questions.
-Many natural prediction problems decompose into a finite set of knowledge and skills that are “quantized” into discrete chunks (quanta) (Michaud et al., 2023).
-Models must learn these quanta to reduce loss.
-Understanding a network reduces to enumerating its quanta.
-Other studies (Schaeffer et al., 2023) prove useful ways to measure quanta in mathematical prediction problems.
+Mechanistic understanding of arithmetic in transformers.
+Early work by Nanda et al.
+(2023) revealed that one-layer models implement modular addition through discrete Fourier transforms, converting addition into rotations in frequency space - showing transformers can find unintuitive algorithms for basic mathematics.
+Quirke and Barez (2024) detailed an algorithm for n-digit addition in a single-layer model that achieved 99% accuracy, identifying a cascading carry-one failure mode.
+Subsequent work on training dynamics (Musat, 2024) and the finding that even random transformers exhibit algorithmic capabilities (Zhong and Andreas, 2024) suggest architectural biases toward systematic computation.
+Architectural modifications and specialized models.
+Several approaches achieve near-perfect arithmetic accuracy through architectural changes.
+Reformatting inputs to present least-significant digits first by adding position embeddings and zero-padding strategies (McLeish et al., 2024).
+Cho et al.
+(2024) align computational flow with carry propagation.
+Specialized models demonstrate superior performance: MathGLM’s 10M parameter model achieves 100% accuracy on integer addition through step-by-step reasoning (Yang et al., 2023), while Qiu et al.
+(2024) reached 99.9% accuracy on 5-digit multiplication with a tiny transformer that outperforms GPT-4.
+Subtraction is much less studied, though Zhang et al.
+(2024) identified symmetries between addition and subtraction circuits, with the same attention heads handling both operations.
+Production systems versus specialized models.
+The Claude 3.5 Haiku circuit analysis Lindsey et al.
+(2025) reveals how a frontier model implements addition.
+Rather than the clean n-digit algorithms found in specialized models, Claude employs a “bespoke” circuit combining lookup tables for memorized facts with magnitude estimation pathways.
+This distributed circuit achieves perfect accuracy on 2-digit addition using mechanisms distinct from both human methods and specialized models.
+The gap between Claude’s complex implementation on short addition questions and the elegant n-digit algorithms in specialized models illustrates a fundamental challenge: production LLMs trained on diverse tasks develop hybrid strategies rather than the interpretable, generalizable algorithms that specialized models achieve.
 3 Methodology
 Transformer models learn addition algorithms different from traditional human methods.
-We define an alternative, mathematically-equivalent framework for addition and demonstrate our models implements this approach.
+We define an alternative, mathematically-equivalent framework for addition and demonstrate our models implement this approach.
 3.1 Mathematical Framework
-Consider the task of adding two $n$-digit numbers together.
-We define the first number (or addend) as $D=\{D_{n-1},D_{n-2},\ldots,D_{0}\}$ and the second number as $D^{\prime}=\{D^{\prime}_{n-1},D^{\prime}_{n-2},\ldots,D^{\prime}_{0}\}$ and the answer (or sum) as $A=\{A_{n},A_{n-1},\ldots,A_{0}\}$.
-Figure 2 shows an illustrative example.
-Figure 2: For 5-digit addition, our model has 12 input (question) and 6 output (answer) token positions.
-We name the question tokens D4, …, D0, and D’4, …, D’0 and the answer tokens A5, …, A0.
-For n-digits, we use the terms $D_{n}$, $D^{\prime}_{n}$and $A_{n}$.
-We adopt the framework from Quirke & Barez (2024) for our model’s addition process, where they break down the task of multi-digit addition into an addition subtask and a carry subtask.
-Firstly, the “Simple Addition” subtask $A_{n}.SA$ naively calculates the sum of corresponding digit pairs in the addends:
-$A_{n}.SA=(D_{n}+D^{\prime}_{n})\mod 10$ (1)
-The “Simple Carry” subtask $A_{n}.SC$ then determines whether the addition between the two digits creates a carry bit:
-$A_{n}.SC=\begin{cases}1&\text{if }(D_{n}+D^{\prime}_{n})\geq 10,\\ 0&\text{otherwise}.\end{cases}$ (2)
-By combining these two subtasks, the model and framework from Quirke & Barez (2024) is capable of handling simple addition and carry bits generated directly from digit pair addition.
-However, it struggles with problems with ’cascading carry’ bits, where a carry bit from one digit position propagates to the next.
-Consider “00144+00056=000210”.
-Adding 4+5 in the tens position doesn’t generate a carry bit directly, but a carry bit propagates from the ones position.
-A model only summing digit pairs and their direct carry bits would fail, producing an incorrect result like “00144+50006=000110”.
-The Quirke & Barez (2024) 1-layer model could cascade carry bits across two digits, but not three or more.
-3.2 Extending the Mathematical Framework
-To solve $n$-digit addition with high accuracy, we need to extend the basic framework to handle carry bits that cascade through multiple digits.
-Consider the problem ”44444+55556=100000”.
-For an autoregressive model to correctly predict the first answer digit as ”1”, it needs to track how the carry bit cascades from the rightmost digit all the way to the leftmost position in a single forward pass.
-This is particularly challenging for an autoregressive model that predicts answer tokens from left to right.
-To solve this, we extend the mathematical framework with two key subtasks.
-First, the digit-level subtask ”Sum Type” $A_{n}.ST$ analyzes each pair of digits and classifies their sum as either definitely causing a carry, possibly causing a carry, or definitely not causing a carry.
-This is implemented by a TriCase operation that maps each digit pair ($D_{n}$ and $D^{\prime}_{n}$), TriCase into one of three possible states:
-$\begin{split}A_{n}.ST&=\underbrace{\text{TriCase}}_{(D_{n},D^{\prime}_{n})}=\\ &\begin{cases}ST10&\text{if }(D_{n}+D^{\prime}_{n})\geq 10\text{ (definite carry)}\\ ST9&\text{if }(D_{n}+D^{\prime}_{n})=9\text{ (possible carry)}\\ ST8&\text{if }(D_{n}+D^{\prime}_{n})\leq 8\text{ (no carry)}\\ \end{cases}\end{split}$ (3)
-The key insight is that while some digit pairs definitely generate a carry (like 5+6), and others definitely don’t (like 2+3), pairs that sum to 9 (like 4+5) may or may not generate a carry depending on whether there’s an incoming carry bit from the next lower digit.
-Then, we introduce the ”Sum Validation” subtask $A_{n}.SV$ that processes these carry states sequentially to determine whether any ”possible carries” become ”definite carries”.
-This subtask is implemented using a TriAdd operation that examines each digit’s sum type along with the sum type of the next lower digit:
-$\begin{split}A_{n}.SV&=\underbrace{\text{TriAdd}}_{(A_{n}.ST,A_{n-1}.ST)}=\\ &\begin{cases}ST10&\text{if }A_{n}.ST=ST10\text{ or }\\ &\quad(A_{n}.ST=ST9\text{ and }\\ &\quad A_{n-1}.ST=ST10),\\ ST8&\text{otherwise.}\end{cases}\end{split}$ (4)
-Revisiting the problem ”44444+55556=100000”, the model can accurately determine whether the first answer digit A5 should be 1 or 0 by examining all digit pairs in sequence.
-This chain of operations uses TriAdd to process the Sum Types from least to most significant digits:
-$A1.SV$ $=\text{TriAdd}(A1.ST,A0.ST)$
-$A2.SV$ $=\text{TriAdd}(A2.ST,A1.SV)$
-$A3.SV$ $=\text{TriAdd}(A3.ST,A2.SV)$
-$A4.SV$ $=\text{TriAdd}(A4.ST,A3.SV)$
-$A5$ $=(A4.SV=ST10)$
-Most importantly, by using this cascading calculation to calculating A5, the model also simulatenously calculates the accurate carry bits for each answer digit.
-The carry bits are simply the SV values we computed along the way—we already have A3.SV for A4’s carry bit, A2.SV for A3’s carry bit, and so on.
-Within our extended framework, the model only requires the subtasks $A_{n}$.SA, $A_{n}.ST$ and $A_{n}.SV$ to accurately perform $n$-digit addition accurately.
-Some models may also use the redundant $A_{n}$.SC subtask.
-Figure 3 shows how our model’s algorithm uses SA, SC, $ST$ and $SV$ to perform addition.
-This algorithm is like the 99% accurate Quirke & Barez (2024) algorithm, but contains an additional circuit (the shaded boxes) to calculate “cascading carry one” data to solve $n$-digit addition with >99.999% accuracy.
-Figure 3: To predict with >99.999% accuracy, the addition algorithm first calculates “carry one” values $(A_{n}.ST)$, combining them into “cascading carry one” values $(A_{n}.SV).$ At the “+” token, $A_{n-1}.SV$ gives the first answer digit as 1 or 0.
-The other answer digits are calculated by combining “base add” $(A_{n}.BA)$ and “carry one” $(A_{n}.SC)$ calculations with the pre-calculated $A_{n}.SV$ values.
-3.3 Techniques
-To undetsand how the model implements arithmetic operations, we analyzed both attention heads and MLP layers at each token position.
-We define a node as either an attention head or MLP layer at a specific token position (e.g., the first attention head in the first layer processing the third digit).
-We employed four techniques to understand each node’s function:
-  1. 1.
-    Intervention Ablation.
-To determine the role of a node’s output, we replace its output with the mean vector of its outputs across a batch, then measured how this impacted downstream predictions.
-For each subtask, we developed automated intervention ablation tests across $n$-digits to verify whether a node implemented that subtask.
-  2. 2.
-    Attention Patterns.
-For each attention head node, we examined which input tokens received significant attention (defined as post-softmax attention score $>0.01$).
-This revealed which parts of the input each attention head considered relevant.
-  3. 3.
-    Principal Component Analysis (PCA).
-We applied PCA to node outputs, especially when investigating attention heads that our framework suggests might have descrete states (e.g.
-tri-state or bi-state outputs).
-  4. 4.
-    Question Complexity.
-We categorized questions by computational complexity (App.
-G).
-For addition, categories S0-S4 reflect how many sequential digits a “carry one” operation cascades through.
-For subtraction, categories M0-M4 reflect a similar cascading of ”borrow one” operations.
-We analyzed which nodes were necessary for correct predictions in each category.
+For addition and subtraction of two $n$-digit numbers, we use the Fig. 3 notation. (Detail in App. A).
+Figure 3: For 5-digit addition and subtraction, our notation for the main input tokens is $D4$, …, $D0$ and $D^{\prime}4$, …, $D^{\prime}0$.
+For output tokens it is $A6$, .., $A0$.
+For n-digits, we use the notation $D_{n}$, .., $D0$, $D^{\prime}_{n}$, .., $D^{\prime}0$ and $A_{n}$, .., $A0$.
+3.2 Addition Algorithm Description
+For addition, aligned with Quirke and Barez (2024), we define the “Base Add” subtask $SA_{n}$ to compute the digit-wise sum modulo 10:
+$SA_{n}=(D_{n}+D^{\prime}_{n})\mod 10$ (1)
+To solve $n$-digit addition with high accuracy, the framework must handle carry bits that cascade through multiple digits.
+For an autoregressive model to correctly predict the first answer digit in “555+448=+1003”, as “1”, it must cascade the carry bit from the rightmost digit to the leftmost digit in a single forward pass.
+This is particularly challenging for an autoregressive model that processes tokens from left to right.
+Aligned to this challenge we define two new subtasks.
+First, we introduce subtask ST.
+It classifies a digit pair sum as definitely causing a carry (e.g.
+6+7), definitely not causing a carry (e.g.
+2+3), or possibly causing a carry (e.g.
+5+4).
+The 5+4 case is uncertain as the digits sum to 9 and a carry from the next-lower-value digit would cause a carry:
+[…]
+The “TriAdd” function combines two $ST_{n}$ style values X and Y.
+Only if both X and Y are U, does TriAdd return U (there is still uncertainty), otherwise it returns 0 or 1 (the carry bit value is known):
+[…]
+Finally, we introduce the SV subtask that uses TriAdd to handle multidigit carry cascades. We describe $SV_{n}$ using the first 3 examples:
+$SV0$ $=ST0$ (4)
+$SV1$ $=\text{TriAdd}(ST1,ST0)$ (5)
+$SV2$ $=\text{TriAdd}(\text{TriAdd}(ST2,ST1),ST0)$ (6)
+The highlighted (solid-color) part of Figure 1 illustrates the SV2 calculation.
+Note that the calculation first combines ST2 and ST1 (the higher value digits) before combining the result with ST0.
+This formulation mirrors the left to right order in which the model processes tokens i.e.
+from higher value digits to lower value digits.
+The last term in all $SV_{n}$ calculations is ST0 which is always 0 or 1, so all $SV_{n}$ evaluate 0 or 1.
+That is, at the end of the $SV_{n}$ calculation any uncertainty has been resolved.
+The $SV_{n}$ are perfectly accurate multidigit cascade carry values.
+Their calculation occurs across multiple tokens and is completed by the “=” token.
+Combining the $SV_{n}$ values with the $SA_{n}$ values as shown in Figure 1 gives perfectly accurate answer tokens.
+3.3 Subtraction Algorithm Description
+Our subtraction algorithm mirrors our addition algorithm, replacing addition’s “cascading carry one” with subtraction’s “cascading borrow one”.
+Both operations use the same iterative approach to reducing “cascading” uncertainty token by token.
+We introduce “Base Diff” $MD_{n}$ defined as $D_{n}$- $D^{\prime}_{n}$ modulo 10 (paralleling $SA_{n}$). We also introduce $MB_{n}$ for the single-digit “borrow one” case (paralleling $ST_{n}$):
+[…]
+Finally, we introduce $MV_{n}$ which parallels $SV_{n}$ to handle the “cascading borrow one” edge case.
+With this formulation, the addition and subtraction algorithms have the same structure (just replacing $SA_{n}$ with $MD_{n}$, $ST_{n}$ with $MB_{N}$, and $SV_{n}$ with $MV_{N}$).
+Subtraction poses an additional algorithmic difficulty - some questions give a positive answer (e.g.
+325-129=+196) and others a negative answer (e.g.
+325-329=-004).
+Note that all the answer digits are different in the above examples.
+While there are multiple ways this could be handled mathematicaly, we introduce “Neg Diff” $ND_{n}$ defined as $D^{\prime}_{n}$- $D_{n}$ modulo 10 (the opposite of $MD_{n}$).
+At the “=” token, the $MV_{n}$ avalues are known, and the algorithm selects $MD_{n}$ or $MD_{n}$ values as appropriate.
+Figure 4: Our subtraction algorithm parallels our addition algorithm but it refines “cascading borrow one” uncertainty over multiple tokens.
+The refined $MV2$ value determines both the answer sign (“+” or “-”) and which digit values ($MD_{n}$ or $ND_{n}$) to use for the final output.
+In Fig.
+4, the value $MV2$ determines the answer sign.
+While the following calculations could use $MV2$ too, in fact they attend to the answer sign itself.
+We introduce the subtask SGN to reflect this.
+3.4 Mixed Algorithm Description
+For mixed models handling both operations, we add the subtask OPR meaning “attends to the prompt operation token”.
+The mixed models uses OPR and SGN to determine whether to output addition, positive-subtraction, or negative-subtraction answer digits.
 4 Experiments
-4.1 Training a Five-Digit Addition Model
-Previous work by Quirke & Barez (2024) achieved $\sim$ 99% accuracy on 5-digit addition using a 1-layer addition model.
-To improve upon this, we systematically explored different model architectures and found that a transformer with 2 layers and 3 attention heads per layer was the minimal configuration needed to achieve >99.999% accuracy (see App.
-D for our exploration of alternative architectures, and App.
-B for model configuration details.).
-This 2-layer architecture doubles the model’s computational capacity and enables cross-layer attention composition for more sophiscated algorithms (Elhage et al., 2021b).
-We trained our model to perform 5-digit addition using an architecture with 2 layers and 3 attention heads per layer.
-The model’s vocabulary consisted of 14 tokens: the digits 0-9 plus the mathematical operators +, -, =, *, and /.
-We used a batch size of 64, learning rate of 0.00008, and weight decay of 0.1.
-To ensure robust performance on edge cases, we trained on an infinite dataset that was deliberately enriched with rare numerical combinations.
+4.1 Training Models
+Our models’ vocabulary consisted of 14 tokens: the digits 0-9 plus the mathematical operators +, -, =, *, and /.
+(Also see App.
+C).
+We created an infinite training dataset enriched with rare numerical combinations (mainly cascading use cases).
+(Also see App.
+D)
+We trained addition-only, subtraction-only, and mixed (addition and subtraction) models.
+We systematically explored different model architectures and settled on 2 layers and 3 or 4 attention heads.
+(Also see App.
+F)
+We trained used a batch size of 64, learning rate of 0.00008, and weight decay of 0.1.
 The loss function was defined as the mean negative log likelihood across all output tokens.
-After 30,000 training batches, the model achieved a final training loss of approximately $2.3\times 10^{-8}$, and subsequent testing confirmed accuracy exceeding 99.999%.
-For comprehensive training details and results, see App.
-C and Tab.
+Training was stopped when the loss was very low e.g.
+$2\times 10^{-8}$.
+(Also see App.
+G, Fig.
+5, Fig.
+9).
+Each model’s accuracy was tested on 1 million questions.
+Robust statistical accuracy ranges were calculated, but for ease of presentation, we say a model has 99.999% accuracy if it got less than 10 of the 1 million questions wrong.
+Most models achieved this.
+The resulting 49 models, including detailed accuracy results, are summarized in Tab.
+5 and Tab.
 6.
-Table 1: For a sample model, all nodes used in predictions are shown by token position (horizontally) and model layer (vertically), detailing the answer digits they impact.
-Here, the attention heads in token position P10 labelled A5..3 help predict the answer digits A3, A4 and A5.
-For all addition and mixed models studied, before the ”=” token, each node often calculates data used to predict multiple answer digits.
-After the ”=” token, all nodes in a given token position are used to predict a single answer digit.
-(P6) (P9) (P10) (P11) (P12) (P13) (P14) (P15) (P16) (P17)
-D’4 D’1 D’0 = + A5 A4 A3 A2 A1
-L0H0 A4 A2 A1
-L0H1 A5 A5..3 A5 A3 A0
-L0H2 A5 A5..1
-L0MLP A5..2 A4 A3 A2 A1 A0
-L1H0
-L1H2 A4 A3 A2
-L1MLP A1 A0
-4.2 Investigating Five-Digit Addition
-Through ablation experiments, we found that the model relies on just 36 nodes distributed across nine token positions (Figs 7 and 8).
-We analyzed how these nodes affect computation complexity and answer accuracy through ablation studies (see Figs 6 and Table 1).
-Core Algorithm Discovery.
-The model’s most challenging task is predicting A5, the first answer digit, at position P11.
-This digit, which can only be 0 or 1, depends on potential carry-bit cascades through multiple positions (e.g., in 55555+44445=100000).
-After two unsuccessful algorithm hypotheses (detailed in App.H andI), we discovered the model’s fundamental mechanism: it uses a minimal representation of carry information through what we call the TriCase quanta.
-The model computes $A_{n}.ST$ using bigram operations that map digit pairs to result tokens (e.g., ”6” + ”7” = ST10) in positions P8 to P11.
-For a specific 5-digit addition model instance, we mapped the algorithm to individual nodes and verified each node’s role using ablation intervention.
-Figure 3 diagrams the algorithm, with node details in App.
-J.
-The model adheres to all known constraints and achieves >99.999% accuracy using 36 nodes, as compared to the 1-layer model uses 21 nodes to achieve 99% accuracy.
-Table 2: All addition models studied implement our addition algorithm.
-The algorithm SA, SC and $ST$ subtasks all exist for each digit and in appropriate token positions.
-For a sample model, this map shows the subtask locations.
-Interestingly, here each SA subtask is shared across two attention heads.
+Figure 5: The 5- to 15-digit 2-layer 3-head addition models have very low loss. With more digits training takes longer. Details in Tab.5.
+4.1.1 Training mixed models
+To explore how transformers can learn multiple arithmetic operations simultaneously, we conducted experiments with models capable of both addition and subtraction.
+We were particularly interested in whether knowledge from addition-specialized models could transfer to this expanded task.
+We trained 14 mixed models with various architectures, including 6- and 10-digit models with 2-3 layers and 3-4 attention heads (see Tab.6).
+Each model was initialized using weights from a highly accurate addition-only model, specifically copying the weights into the first two layers and first three heads of the mixed model.
+During training, we used a curriculum of 80% subtraction and 20% addition examples, using enriched data for both operations (See Fig.
+6).
+This initialization approach proved effective, with most models quickly achieving high accuracy on both operations.
+We explored alternative training strategies, such as periodically “freezing” the inserted attention heads or MLPs by restoring their original addition-model weights every 100 training steps.
+However, these more constrained approaches resulted in lower accuracy, suggesting that some flexibility in modifying the initialized weights was necessary for optimal performance (Also see App.
+K)
+Figure 6: Mixed model initialization and analysis: (1) Train an accurate 5-digit addition model.
+(2) Reverse-engineer it to identify the implemented algorithm.
+(3) Insert its attention head and MLP weights (in brown) into a larger model.
+(4) Train the new model on 80% subtraction and 20% addition.
+(5) The resulting model predicts accurately and reuses the inserted addition circuits for both tasks.
+We did a seed sensitivity analysis. App. L includes a visualization Fig. 15 of the range of training losses for different model categories.
+4.2 Experimental Results
+To verify whether our models implement the algorithms described above, we developed a systematic framework for analyzing model internals.
+For each trained model, we investigated whether it contains the specific computational components our algorithms require and whether these components interact according to the algorithmic constraints.
+4.2.1 Implementation Analysis Framework
+Consider a 6-digit addition model. If it implements our algorithm, it must satisfy these constraints:
+  • Functional accuracy: The model must correctly add any two 6-digit numbers, selecting the unique correct answer from nearly 2 million possible answers.
+  • Complete component coverage: The model must implement the required subtasks for each digit position i.e. 6 $SA$, 6 $ST$, 6 $SV$, and 7 answer digit finalization nodes.
+  • Component-level behavior: Each node implementing a subtask must obey certain conditions. For instance, a node implementing $ST2$ must:
+    • Attend primarily to tokens $D_{2}$ and $D^{\prime}_{2}$ (post-softmax attention >0.01>0.01)
+    • Be positioned after $D^{\prime}_{2}$ but before the “=” token
+    • A Primary Component Analysis (PCA) should show 3 distinct clusters. (See Fig. 7.)
+    • Given three test question sets, created specifically to trigger the 3 $ST2$ values, each question set should align with exactly one of the PCA clusters.
+  • Algorithmic ordering constraints: Node locations must respect computational dependencies.
+Since SV1 = TriAdd(ST1, ST0), a node computing SV1 must be positioned after nodes computing ST1 and ST0.
+Our 6-digit model has over 30 such ordering constraints.
+  • Causal verification through intervention: When we identify a candidate node for a specific subtask (e.g.
+ST2), we can construct paired test questions differing only in that subtask’s value.
+By intervening on the candidate node and swapping its activations between the test pairs, we can predict the model’s output from nearly 2 million possible answers.
+For models with more than 6 digits, the constraints list grows further.
+The subtraction and mixed algorithms each have a more extensive constraints list.
+With 49 models to test, we developed an automated testing framework to:
+  • Systematically search for nodes over each subtask and digit position combination retaining only those that satisfy all the constraints in our algorithm
+  • Compare the list of nodes found with the algorithmic requirements (including intra-node relationships).
+  • Visualize the results.
+4.2.2 Addition model results
+We tested a single 1-layer addition model.
+Our framework found the $SA_{n}$ and $SC_{n}$ (single-digit carry calculated during answer tokens) nodes as described by Quirke and Barez (2024) that give the model its 99% accuracy.
+Testing our fifteen 2-layer addition models, we find all models contain the subtasks required by our addition algorithm, and satisfy the algorithmic requirements.
+This consistent pattern suggests we have identified a robust and generalizable algorithm that naturally arises when transformer models are trained on addition in our setup.
+There is some variability between models:
+  • In a given model, $ST_{N}$ is not as “tidy” as shown in Fig.
+1.
+Different models select different attention heads to calculate say ST2.
+Redundantly, some models have two nodes that both calculate say ST1.
+See Tab.
+1.
+  • Some models combine 2 attention heads (with the same token position and layer) to calculate say SA2. Many models only use one attention head.
+  • For ST nodes, the PCA trigrams have difference appearances in each attention head (- but all have 3 clusters aligned to the ST values. See Fig. 7.
+  • All models have $ST$ nodes, but some models also have and use $SC$ nodes. The $SC$ nodes are redundant: the $ST$ values are equivalent and the $SV$ values are superior! See Tab. 1.
+Figure 7: For a 5-digit addition model, these three sample PCAs each show 1 attention head at a token position tested for 1 answer digit.
+Each dot represents a question, colored by the question’s expected ST value (0, 1, or U).
+While the PCA data varies across cases, the 3 ST classes form visually distinct clusters in each plot, suggesting that the attention head’s activation patterns differ systematically based on ST value and that the model could potentially distinguish between these 3 cases.
+Table 1: For a sample 5-digit addition model, we show a compacted location map for the SASA, SCSC and STST subtasks.
+Interestingly 1) the $ST$ nodes are in a semi-random order 2) the second $ST1$ node is redundant 3) the model uses SCSC nodes which are redundant 4) each $SA$ subtask is shared across two attention heads.
 (P11) (P12) (P13) (P14) (P15) (P16) (P17) (P18) (P19) (P20)
 D’1 D’0 = + A6 A5 A4 A3 A2 A1
-L0H0 A2.ST A3.ST A1.ST A4.ST A4.SC A3.SC A2.SC A1.SC A0.SC
-L0H1 A1.ST A0.ST A5.SA A4.SA A3.SA A2.SA A1.SA A0.SA
-L0H2 A5.ST
-4.3 Investigating n-digit Addition models
-To test whether this algorithm generalises beyond 5-digit addition, we conducted a comprehensive study of larger addition models.
-We trained 15 2-layer transformer models to handle addition problems ranging from 5- to 15-digits.
-These models used varying random seeds and optimizers, and included a sign token in their output format (e.g.
-111111+222222=+0333333).
-Figure 4: The 5- to 13-digit 2-layer 3-head addition models have very low loss and >99.999% accuracy. With more digits the model loss and accuracy is worse. Details in Tab.6.
-Analysis Framework.
-To systematically analyse these models, we first created a declarative method for specifying each subtask within the addition process.
-Each subtask definition includes (1) the input digits it processes, (2) the answer digit it affects, (3) the complexity level it influences, and (4) expected PCA patterns and ablation test results.
-Using this declaration, we identified the nodes performing these subtasks across all models.
-Since our framework lays out the addition algorithm in terms of its required subtasks., we can also verify whether a given model implements this algorithm by checking for the presence and proper sequencing of all required subtasks.
-For example, our addition algorithm specifies that the model must execute $A_{n}.ST$ subtasks for each question digit before the first answer token.
-Table 2 shows how an addition model fully implements all required subtasks in its nodes.
-Using this framework, we found that all fifteen 2-layer addition models, regardless of their size, implemented the same core algorithm we identified in our 5-digit model.
-This consistent pattern suggests we have identified a robust and generalizable algorithm that naturally arises when transformer models are trained on addition in our setup where the digits are tokenized individually.
-4.4 Training n-digit Mixed models
-To explore how transformers can learn multiple arithmetic operations simultaneously, we conducted experiments with models capable of both addition and subtraction.
-We were particularly interested in whether knowledge from addition-specialized models could transfer to this expanded task.
-We trained 14 mixed models with various architectures, including 6- and 10-digit models with 2-3 layers and 3-4 attention heads.
-Each model was initialized using weights from a highly accurate addition-only model, specifically copying the weights into the first two layers and first three heads of the mixed model.
-During training, we used a curriculum of 80% subtraction and 20% addition examples, with deliberate enrichment of rare edge cases for both operations.
-This initialization approach proved effective, with most models achieving high accuracy on both operations.
-We explored alternative training strategies, such as periodically ”freezing” the inserted attention heads or MLPs by restoring their original addition-model weights every 100 training steps.
-However, these more constrained approaches resulted in lower accuracy, suggesting that some flexibility in modifying the initialized weights was necessary for optimal performance.
-The initialized mixed experimental models achieve >99.999% accuracy and contain the same subtasks, implying a common algorithm.
-Figure 5: The 5- to 8-digit initialized mixed models have very low loss and >99.999% accuracy. With more digits the model loss and accuracy is worse. Details in Tab.7.
-4.5 Investigating n-digit Mixed models
-The mixed models revealed fascinating patterns in how they handled different types of arithmetic problems.
-Our hypothesis was that the model handled three distinct question classes: addition, positive-answer subtraction (e.g., 10009-10000=+000009), and negative-answer subtraction (e.g., 10009-20000=-009991).
-Through ablation studies, we found that many nodes specialized in handling specific question classes, while others became ”polysemantic,” capable of processing multiple classes simultaneously (More details in Tab.3).
-Most notably, the majority of nodes initialized from the addition model became polysemantic during training.
+L0H0 ST2 ST3 ST1 ST4 SC4 SC3 SC2 SC1 SC0
+L0H1 ST1 ST0 SA5 SA4 SA3 SA2 SA1 SA0
+L0H2 ST5
+4.2.3 Mixed model results
+Testing our accurate ”mixed” (both addition and subtraction) models, we find all models contain the subtasks required by our addition algorithm, and satisfy the algorithmic requirements.
+The mixed models have the same types of variability that the addition models have.
+Evolution of Polysemantic Nodes. Zhang et al. (2024) identified symmetries between addition and subtraction circuits, with the same attention heads handling both operations. We expand on this finding.
+Consider the mixed models that we initialized with accurate, smaller addition models before training.
+The majority of inherited nodes became polysemantic during training.
 Rather than developing entirely new circuits for subtraction, the models adapted their addition circuits to handle both operations.
+For example, a node that originally only performed Base Add ($SA$ ) often learns to handle three subtasks: addition ($SA$ ), positive-answer subtraction diff ($MD$ ), and negative-answer subtraction diff ($ND$ ).
+(Refer to Tab.
+2.)
+Table 2: For this mixed model, in the last 5 tokens, polysemantic attention heads simultaneously generate outputs for the three question classes (S, M and N).
+Other heads calculate the question class by attending to the question operation (OPR) token and the answer sign (SGN) token.
+We assume the MLP layers then select the output appropriate for the class.
+(P9) (P10) (P11) (P12) (P13) (P14) (P15) (P16) (P17) (P18) (P19) (P20)
+D’3 D’2 D’1 D’0 = A7 A6 A5 A4 A3 A2 A1
+L0H0 MT4 MT3 ST4 SC2 SC1 SC0 OPR
+MT3 MT4 SC4 SC3 NB2 MB1 MB0 SGN
+OPR NB1 NB0
+L0H1 ST4 ST2 ST1 ST3 ST0 SA5
+MT2 MT1 MT0 MD5 SA4 SA3 SA2 SA1 SA0
+L0H2 ST5 SA4 MD4 MD3 MD2 MD1 MD0
+OPR ND5 ND4 ND3 ND2 ND1 ND0
+SGN
+L0H3 OPR OPR OPR OPR OPR OPR
+SGN SGN SGN SGN SGN SGN
+We believe that this adaptation was likely because these three operations share a common structure: each maps a pair of digits (10 possibilities for $D_{n}$× 10 possibilities for $D^{\prime}_{n}$= 100 input cases) to a single-digit result (10 possible outputs).
+Similarly, some models adapted nodes that originally handled carry bits ($ST$) to also process borrow bit operations ($MT$).
 Table 3: Mixed models re-use most inserted addition-model nodes.
@@
 Addition 61 64% 42 88%
-Positive-answer sub 70 73% 40 83%
-Negative-answer sub 53 55% 29 60%
-Defining Subtraction Subtasks.
-Subtraction requires two distinct computational paths based on the relative magnitude of the numbers.
-When D $\geq$ D’, as in 325 - 123 = +202, we perform direct subtraction with borrowing.
-When D $<$ D’, as in 325 - 456 = -131, we must instead compute the magnitude of the difference and mark it negative.
-Hence, this computational difference necessitates separate frameworks for positive-answer (M) and negative-answer (N) subtraction, each with parallel operations to addition’s Base Add ($SA$) and Carry One ($SC$), which we define in Tab.
-4.
-We also define complexity measures for subtraction (see App.
-G and  N).
-Table 4: We define 2 “positive-answer subtraction” and 2 “negative-answer subtraction” subtasks that parallel the addition subtasks, Base Add SA and Carry One SC .
-Name +ve Sub -ve Sub Definition
-Base Diff MD ND $D_{n}$- $D^{\prime}_{n}$10
-Borrow One MB NB $D_{n}$- $D^{\prime}_{n}$$<0$
-Just as addition must handle ”cascading carry” cases (e.g., 99999+1=100000), subtraction must handle ”cascading borrow” cases.
-Consider 100000-1=099999, where a borrow cascades through multiple zeros.
-To handle such cases efficiently, we define a three-state classification $MT$ that parallels addition’s $ST$ :
-$\begin{split}A_{n}.MT&=\underbrace{\text{TriCase}}_{(D_{n},D^{\prime}_{n})}=\\ &\begin{cases}MTN&\text{if }D_{n}<D^{\prime}_{n}\\ MT0&\text{if }D_{n}=D^{\prime}_{n}\\ MT1&\text{if }D_{n}>D^{\prime}_{n}\\ \end{cases}\end{split}$ (5)
-For M questions, MTN is a definite borrow one, MT0 is a possible borrow one (depending on the results of other digit calculations) and MT1 is definitely not a borrow one.
-For N questions, the interpretation is the opposite.
-Finally, paralleling the ”cascading carry one” $An.SV$ operation in addition, we define cascading borrow calculations $A_{n}.MV$ and $A_{n}.NV$ for positive and negative cases.
-Evolution of Polysemantic Nodes.
-During training, the mixed model adapted the nodes it inherited from the addition model.
-Nodes that originally only performed Simple Addition ($SA$) developed the ability to simultaneously handle three distinct operations: addition ($SA$), positive-answer subtraction ($MD$), and negative-answer subtraction ($ND$).
-(Refer to Table 5.) We believe that this adaptation was likely because these three operations share a common structure: each maps a pair of digits (10 possibilities for $D_{n}$× 10 possibilities for $D^{\prime}_{n}$= 100 input cases) to a single-digit result (10 possible outputs).
-The model similarly adapted nodes that originally handled carry bits ($SC$) to also process borrow operations for both positive ($MB$) and negative ($NB$) subtraction cases.
-Table 5: For mixed models, in later tokens, polysemantic attention heads simultaneously generate outputs for the three question classes addition S, M and N.
-Other heads calculate the question class by attending to the question operation (OPR) token and the answer sign (SGN) token.
-The MLP layers then select the output appropriate for the class.
-In this sample map, from P16, the first 3 rows contain many polysemantic nodes, while the 4th row calculates the question class.
-(P9) (P10) (P11) (P12) (P13) (P14) (P15) (P16) (P17) (P18) (P19) (P20)
-D’3 D’2 D’1 D’0 = A7 A6 A5 A4 A3 A2 A1
-L0H0 A4.MT A3.MT A4.ST A2.SC A1.SC A0.SC OPR
-A4.GT A3.MT A4.MT A4.SC A3.SC A2.NB A1.MB A0.MB SGN
-OPR A1.NB A0.NB
-L0H1 A2.ST A1.ST A3.ST A0.ST A5.SA
-A4.ST A2.MT A1.MT A3.GT A0.MT A5.MD
-A2.GT A1.GT A0.GT A4.SA A3.SA A2.SA A1.SA A0.SA
-L0H2 A5.ST A5.SA A4.MD A3.MD A2.MD A1.MD A0.MD
-OPR A5.MD A4.ND A3.ND A2.ND A1.ND A0.ND
-SGN A5.ND
-L0H3 OPR OPR OPR OPR OPR OPR
-SGN SGN SGN SGN SGN SGN
-4.6 Mixed model question class detection
-Mixed models must identify whether each input belongs to one of three classes: addition (S), positive-answer subtraction (M), or negative-answer subtraction (N).
-For addition questions (marked by ”+”), classification is immediate.
-For subtraction questions (marked by ”-”), the model must determine whether D $\geq$ D’ to distinguish between M and N classes.
-This classification must occur by the ”=” token since it determines whether the first answer token will be ”+” or ”-”.
-The model can determine D $\geq$ D’ in two ways.
-First, it can use the $A_{n}.MT$ values already computed for subtraction.
-Alternatively, it can use a dedicated comparison circuit based on a new subtask:
-$A_{n}.GT=\begin{cases}1&\text{if }D_{n}\geq D^{\prime}_{n},\\ 0&\text{otherwise}.\end{cases}$ (6)
-For a 4-digit number, this comparison cascades through the digits: D $\geq$ D’ is true if either $A3.GT=1$, or if $A3.GT=0$ and $A2.GT=1$, and so on through $A0.GT$.
-We verify this circuit through ablation tests that show its disruption causes answer signs to flip.
-We found all mixed models incorporate the $A_{n}.MT$ and $A_{n}.GT$ subtasks to determine whether the answer is postiive or negative.
-Typically (but not always), the same node computes both subtasks (see Table 2).
-Since both the $MT$ and GT approaches can validly distinguish between positive and negative outcomes, the model sometimes chooses one approach for some digits and the other approach for others.
-A single node can implement multiple arithmetic operations (e.g., $A2.SA$, $A2.MD$, $A2.ND$) but has no intrinsic way of knowing whether the current question is addition, positive subtraction, or negative subtraction.
-Hence, we hypothesise that it outputs all three possible results into the residual stream.
-A different node must then select the correct result by identifying the question’s operation and required sign.
-In practice, we observe that one or more attention heads attend to the question operator token (OPR) and the initial answer sign token (SGN), encoding the necessary information for the downstream MLP layers to pick which of the three outputs is correct.
-As Table 2 shows, each answer digit typically has one attention head that focuses on both the OPR and SGN tokens, thereby enabling the MLP to filter out the irrelevant outputs and retain the correct one.
+Positive-answer subtraction 70 73% 40 83%
+Negative-answer subtraction 53 55% 29 60%
 5 Conclusion
-Our analysis of 44 transformer models demonstrates their ability to learn highly accurate mathematical algorithms, achieving over 99.999% accuracy on addition tasks up to 13 digits.
+Our analysis of 49 transformer models demonstrate their ability to learn highly accurate mathematical algorithms, achieving over 99.999% accuracy on addition tasks up to 15 digits.
 By investigating how models implement computations using attention heads and MLPs, we uncovered novel computational approaches distinct from traditional human calculation methods.
@@
 Our research provides insights into how transformers learn precise algorithmic computations across related tasks, developing a reusable library of mechanistic interpretability tools that advances our understanding of neural network computational learning.
-Impact Statement
-Our work aims to explain the inner workings of transformer-based language models, which may have broad implications for a wide range of applications.
-A deeper understanding of generative AI has dual usage.
-While the potential for misuse exists, we discourage it.
-The knowledge gained can be harnessed to safeguard systems, ensuring they operate as intended.
-It is our sincere hope that this research will be directed towards the greater good, enriching our society and preventing detrimental effects.
-We encourage responsible use of AI, aligning with ethical guidelines.
-Appendix A Appendix: Terminology
+6 Limitations and Future Work
+While we identify and test the functional role of each component in the mixed model algorithm, we do not analyze the specific data representations and transformations used by polysemantic nodes, SGN nodes, and OPR nodes in the residual stream, nor the detailed mechanisms by which MLP layers process this information.
+Although our algorithmic-level analysis may be more generalizable across models, this approach leaves gaps in understanding how the identified components actually encode and manipulate information.
+Our automated framework for discovering algorithm subtasks in models, while instrumental in accelerating our research, has limitations.
+Some aspects are specific to our math models.
+Some aspects are generic and have already been used in the published paper by  Harrasse et al.
+(2025).
+Fully reverse engineering neural networks faces significant challenges, including distributed parallel computations difficult to trace to specific components.
+We made progress on an alternative approach: a declarative language describing algorithms in terms of subtasks and a framework to test these descriptions against models.
+However, more work is needed to refine this method and quantity our certainty when comparing evidence gathered from models against an algorithmic description.
+Our models may be useful to researchers investigating the universality of representations / circuits across models.
+Appendix A Appendix: Terminology and Abbreviations
 These terms and abbreviations are used in this paper and the associated Colabs and python code:
@@
   • A : Answer to the question (including answer sign)
-  • An : nth token in the answer. Zero-based. A0 is the units value. The highest token is the “+” or ”-” answer sign
-  • S : Prefix for Addition. Think S for Sum. Aka ADD.
-  • SA : Basic Add. An addition subtask. An.SA is defined as (Dn + D’n) % 10. e.g. 5 + 7 gives 2
-  • SC : Carry One. An addition subtask. An.SC is defined as Dn + D’n >=>=> = 10. e.g. 5 + 7 gives True
-  • SS : Make Sum 9. An addition subtask. An.SS is defined as Dn + D’n == 9. e.g. 5 + 7 gives False
-  • ST : TriCase. An addition subtask. Refer paper 2 for details
-  • ST8, ST9, ST10 : Outputs of the ST TriCase subtask.
+  • An : nth token in the answer. Zero-based. A0 is the units value. The highest token is the “+” or “-” answer sign
+  • S : Prefix for Addition. Think S for Sum.
+  • SA : Base Add. An addition subtask. $SA_{n}$ is defined as (Dn + D’n) % 10. e.g. 5 + 7 gives 2
+  • SC : Carry One. An addition subtask. $SC_{n}$ is defined as Dn + D’n >=>= 10. e.g. 5 + 7 gives True
+  • ST : TriCase. An addition subtask. Refer paper body for details
   • M : Prefix for Subtraction with a positive answer. Think M for Minus. Aka SUB
-  • : Basic Difference. A subtraction subtask. An.MD is defined as (Dn - D’n) % 10. e.g. 3 - 7 gives 6
-  • : Borrow One. A positive-answer subtraction subtask. An.MB is defined as Dn - D’n $<$ 0. e.g. 5 - 7 gives True
-  • MZ : Make Zero. A positive-answer subtraction subtask. An.MZ is defined as Dn - D’n == 0. e.g. 5 - 5 gives True
-  • MT : TriCase. A positive-answer subtraction subtask.
-  • MT1, MT0, MT-1 : Outputs of the MT TriCase subtask.
+  • MD: Basic Difference. A subtraction subtask. $MD_{n}$ is defined as (Dn - D’n) % 10. e.g. 3 - 7 gives 6
+  • MB: Borrow One. A positive-answer subtraction subtask. $MB_{n}$ is defined as Dn - D’n << 0. e.g. 5 - 7 gives True
   • N : Prefix for Subtraction with a negative answer. Think N for Negative. Aka NEG
-  • ND : Basic Difference. A negative-answer subtraction subtask. An.ND is defined as (Dn - D’n) % 10. e.g. 3 - 7 gives 6
-  • NB : Borrow One. A negative-answer subtraction subtask. An.NB is defined as Dn - D’n $<$ 0. e.g. 5 - 7 gives True
-  • NZ : Make Zero. A negative-answer subtraction subtask. An.NZ is defined as Dn - D’n == 0. e.g. 5 - 5 gives True
-  • NT : TriCase. A negative-answer subtraction subtask.
-  • GT : Greater Than. A (positive-answer or negative-answer) subtraction subtask. An.GT is defined as Dn $>$ D’n. e.g. 3 $>$ 5 gives False
+  • ND : Basic Difference. A negative-answer subtraction subtask. $ND_{n}$ is defined as (Dn - D’n) % 10. e.g. 3 - 7 gives 6
+  • NB : Borrow One. A negative-answer subtraction subtask. $NB_{n}$ is defined as Dn - D’n << 0. e.g. 5 - 7 gives True
   • OPR : Operator. A subtask that attends to the + or - token in the question (which determines whether the question is addition or subtraction).
@@
   • EVR : Explained Variance Ratio. In PCA, EVR represents the percentage of variance explained by each of the selected components.
-Appendix B Appendix: Model Configuration
+Appendix B Appendix: Ethics Statement
+Our work aims to explain the inner workings of transformer-based language models, which may have broad implications for a wide range of applications.
+A deeper understanding of generative AI has dual usage.
+While the potential for misuse exists, we discourage it.
+The knowledge gained can be harnessed to safeguard systems, ensuring they operate as intended.
+It is our sincere hope that this research will be directed towards the greater good, enriching our society and preventing detrimental effects.
+We encourage responsible use of AI, aligning with ethical guidelines.
+Appendix C Appendix: Vocabulary
+Each digit is represented as a separate token.
+[Liu and Low, 2023] state that LLaMa’s “remarkable arithmetic ability … is mainly atributed to LLaMA’s consistent tokenization of numbers”.
+The model’s vocabulary contains 14 tokens ( 0, .., 9, +, -, =, *, / ) to enable this and planned future investigations.
+Appendix D Appendix: Training Data
+Training uses a new batch of data each step (aka Infinite Training Data) to minimize memorization.
+Depending on the configuration, each training run processes 1 to 4 million training datums.
+For the 5-digit addition problem there are 100,000 squared (that is 10 billion) possible questions.
+So the training data is much less than 1% of the possible problems.
+Addition and subtraction include rare edge cases.
+For example, these cascades (e.g.
+55555+44446=100001, 54321+45679=1000000, 44450+55550=10000, 1234+8769=10003) are exceedingly rare.
+The data generator was enhanced to increase the frequency of all known edges cases.
+This leads to lower model loss.
+We enriched 60% of training data based on these edge cases (leaving the other 40% of training data random): (1) The initial model failed at cascading carry ones in addition.
+To rectify this, we randomly selected one operand and modified a random subset of digits in that operand to make the selected digit-position sum to 9, increasing the likelihood of a cascading carry one.
+(2) The initial model was worse at the subtraction task when the answer was negative, so we added 1 to each second operand digit (that was 8 or less) to increase the frequency of negative answers.
+(3) When the operands were identical, the initial model predicted -000000 instead of +000000.
+We increased the frequency of this case.
+For example, for 6 digit questions, we increased the frequency from 0.0001% to 0.6%.
+Appendix E Appendix: Surveying LLM Addition Capability
+LLM Gateways provide access to hundreds of LLMs.
+We used the Martian Gateway ( https://app.withmartian.com/ ) to quickly and cheaply access 193 LLMs ( https://withmartian.github.io/llm-adapters/ ) ranging in size from 1 to 405 billion parameters.
+We tested their ability to perform addition accurately on the test questions containing cascading carry ones shown in Tab.
+4.
+A model’s score represents the longest question the model succeeded on before it failed a question.
+The results are shown in Fig.
+2.
+Fig.
+8 shows the size of the models seems uncorrelated to its ability to perform these questions.
+Table 4: Addition prompts used to test LLMs’ ability to handle the cascading carry one use case.
+Test Prompt Correct Answer
+Answer concisely: 6+5= 11
+Answer concisely: 19+87= 106
+Answer concisely: 774+229= 1003
+Answer concisely: 6587+3416= 10003
+Answer concisely: 22605+77398= 100003
+Answer concisely: 532847+467159= 1000006
+Answer concisely: 5613709+4386294= 10000003
+Answer concisely: 72582383+27417619= 100000002
+Answer concisely: 206727644+793272359= 1000000003
+Answer concisely: 7580116456+2419883549= 10000000005
+Answer concisely: 52449010267+47550989737= 100000000004
+Answer concisely: 888522030597+111477969406= 1000000000003
+Figure 8: This scatter plot shows each model’s addition and subtraction score against the model output token cost (an approximate metric for the model size). The maximum possible score is 15.
+At time of writing, 4.76% of models scored 10 or more (specifically nousresearch/hermes-3-llama-3.1-405b, amazon/nova-pro-v1, anthropic/claude-opus-4-1-20250805, mistralai/magistral-medium-2506, openai/gpt-4o-mini-search-preview, openai/gpt-4o-search-preview, openai/o4-mini, qwen/qwen3-coder and tencent/hunyuan-a13b-instruct).
+Some of these models explicitly acknowledge the use of external tools.
+For others it is not obvious whether they use tools.
+The three Chat GPT 5 models tested scored 4, 4 and 7.
+After publication, we will make public the Jupyter notebook used to perform the test.
+It takes less than 20 minutes and costs less than 1 USD to run.
+Appendix F Appendix: Training Setup
 Addition, subtraction and mixed (addition and subtraction) training experiments were done in a Colab notebook.
@@
   • n_heads = 3 or 4: Number of attention heads.
-  • n_digits = 5, 6 or 10: Number of digits in the question.
-Each digit is represented as a separate token.
-(Liu & Low, 2023) state that LLaMa’s “remarkable arithmetic ability … is mainly atributed to LLaMA’s consistent tokenization of numbers”.
-The model’s vocabulary contains 14 tokens ( 0, .., 9, +, -, =, *, / ) to enable this and planned future investigations.
-Training uses a new batch of data each step (aka Infinite Training Data) to minimise memorisation.
-Depending on the configuration, each training run processes 1 to 4 million training datums.
-For the 5-digit addition problem there are 100,000 squared (that is 10 billion) possible questions.
-So the training data is much less than 1% of the possible problems.
-Addition and subtraction include rare edge cases.
-For example, the SS cascades (e.g.
-44445+55555=100000, 54321+45679=1000000, 44450+55550=10000, 1234+8769=10003) are exceedingly rare.
-The data generator was enhanced to increase the frequency of all known edges cases.
-This lead to lower model loss.
-Validation test data covering all edge cases was manually constructed. These test cases are not used during training.
-The Colabs will be made available on publication.
-Appendix C Appendix: Model Loss
-The model defaults to batch size = 64, learning rate = 0.00008 and weight decay = 0.1. The loss function is simple:
-  • Per Digit Loss: For “per digit” graphs and analysis, for a given answer digit, the loss used is negative log likelihood.
-  • All Digits Loss: For “all answer digits” graphs and analysis, the loss used is the mean of the “per digit” loss across all the answer digits.
-In our experimental models, the number of digits in the question varies from 5 to 13, the number of layers varies from 1 to 4, the number of heads varies from 3 to 4.
-Each experimental model’s loss is detailed in Tabs.
-6 and  7.
-Appendix D Appendix: Addition Model Shape
-While we wanted a very low loss addition model, we also wanted to keep the model compact - intuiting that a smaller model would be easier to understand than a large model.
+  • n_digits = 5 .. 15: Number of digits in the question.
+The Colabs are available at https://github.com/PhilipQuirke/quanta_maths
+While we wanted a very low loss models, we also wanted to keep the model compact - intuiting that a smaller model would be easier to understand than a large model.
 Here are the things we tried to reduce loss that didn’t work:
-  • Increasing the frequency of hard (cascading SS ) examples in the training data so the model has more hard examples to learn from. This improved training speed but did not reduce loss.
+  • Increasing the frequency of hard (cascading carry one) examples in the training data so the model has more hard examples to learn from. This improved training speed but did not reduce loss.
   • Increasing the number of attention heads from 3 to 4 or 5 (while still using 1 layer) to provide more computing power.
   • Changing the question format from “12345+22222=” to “12345+22222equals” giving the model more prediction steps after the question is revealed before it needs to state the first answer digit.
+  • Inserting “+” in the answer format (e.g. “12345+22222=034567” becomes “12345+22222=+034567” had no impact on accuracy or the algorithm.
   • With n_layers = 1 increasing the number of attention heads from 3 to 4.
@@
 The smallest model shape that did reduce loss significantly was 2 layers with 3 attention heads.
-Appendix E Appendix: Experimental models
-Forty-four models were trained and analyzed (Refer Tabs.
-6 and  7).
+Appendix G Appendix: Model Loss
+The model defaults to batch size = 64. The loss function is simple:
+  • Per Digit Loss: For “per digit” graphs and analysis, for a given answer digit, the loss used is negative log likelihood.
+  • All Digits Loss: For “all answer digits” graphs and analysis, the loss used is the mean of the “per digit” loss across all the answer digits.
+In our experimental models, the number of digits in the question varies from 5 to 15, the number of layers varies from 1 to 4, the number of heads varies from 3 to 4.
+Each experimental model’s loss is detailed in Tabs.
+5 and  6.
+During training, the models use an AdamW optimizer with a learning rate (LR) of 0.00008, weight decay of 0.1 and betas of (0.9, 0.98).
+During the first 5th of training the LR is linearly increased from 0.01 * LR to LR to help stabilize training.
+During the remainder of training, cosine annealing is used.
+Figure 9: The 5- to 8-digit initialized mixed models have very low loss and >99.999% accuracy. With more digits the model loss and accuracy is worse. Details in Tab.6.
+Appendix H Appendix: Experimental models
+Forty-nine models were trained and analyzed (Refer Tabs.
+5 and  6).
 The models and output of analysis are available on HuggingFace at https://huggingface.co/PhilipQuirke in the quanta_maths_* collections to support further research in AI Safety.
@@
 The number of useful attention heads at token position and useful MLP layers at token position are shown - summarizing the data shown in figures like Fig 11.
-Num Digits Num Layers Num Heads Train Steps Train Seed Train loss Addition Fails / M Heads used MLPs used
-5 1 3 30K 372001 9.4e-2 12621 15 6
-5 2 3 15K 372001 1.6e-8 0 30 16
-5 2 3 40K 372001 2.0e-9 0 22 15
-6 2 3 15K 372001 1.7e-8 2 31 17
-6 2 3 20K 173289 1.5e-8 0 28 17
-6 2 3 20K 572091 7.0e-9 0 35 17
-6 2 3 40K 372001 2.0e-9 0 29 17
-7 2 3 45K 173289 3.0e-9 0 31 20
-8 2 3 45K 173289 3.0e-9 0 35 22
-9 2 3 45K 173289 3.0e-9 0 54 27
-10 2 3 40K 572091 7.0e-9 0 44 28
-11 2 3 50K 173289 8.0e-9 2 56 29
-12 2 3 50K 173289 5.0e-9 3 50 33
-13 2 3 50K 173289 6.3e-8 1 66 31
-14 2 3 60K 173289 5.5e-6 199 68 35
-15 2 3 80K 572091 8.6e-8 10 93 58
+Digits Layers Heads Train Steps Train Seed Train loss Add Fails / M Heads used MLPs used Clopper-Pearson 95% Interval
+5 1 3 30K 372001 9.4e-2 12621 15 6 [1.24e-2, 1.28e-2]
+5 2 3 15K 372001 1.6e-8 0 30 16 [0.00e+0, 3.69e-6]
+5 2 3 40K 372001 2.0e-9 0 22 15 [0.00e+0, 3.69e-6]
+6 2 3 15K 372001 1.7e-8 2 31 17 [2.42e-7, 7.22e-6]
+6 2 3 20K 173289 1.5e-8 0 28 17 [0.00e+0, 3.69e-6]
+6 2 3 20K 572091 7.0e-9 0 35 17 [0.00e+0, 3.69e-6]
+6 2 3 40K 372001 2.0e-9 0 29 17 [0.00e+0, 3.69e-6]
+7 2 3 45K 173289 3.0e-9 0 31 20 [0.00e+0, 3.69e-6]
+8 2 3 45K 173289 3.0e-9 0 35 22 [0.00e+0, 3.69e-6]
+9 2 3 45K 173289 3.0e-9 0 54 27 [0.00e+0, 3.69e-6]
+10 2 3 40K 572091 7.0e-9 0 44 28 [0.00e+0, 3.69e-6]
+11 2 3 50K 173289 8.0e-9 2 56 29 [2.42e-7, 7.22e-6]
+12 2 3 50K 173289 5.0e-9 3 50 33 [4.25e-7, 7.88e-6]
+13 2 3 50K 173289 6.3e-8 1 66 31 [2.53e-8, 5.57e-6]
+14 2 3 60K 173289 5.5e-6 199 68 35 [1.74e-4, 2.00e-4]
+15 2 3 80K 572091 8.6e-8 10 93 58 [4.71e-6, 1.67e-5]
 Table 6: Subtraction-only and mixed models studied.
@@
 The number of useful attention heads at token position and useful MLP layers at token position are shown - summarizing the data shown in figures like Fig 11.
-Num Digits Num Layers Num Heads Train Steps Train Seed Train loss Addition Fails / M Subtract Fails / M Heads used MLPs used
+Digits Layers Heads Train Steps Train Seed Train loss Add Fails / M Sub Fails / M Heads used MLPs used Clopper-Pearson 95% Interval
 Subtraction models
-6 2 3 30K 372001 5.8e-6 N/A 0 40 21
-10 2 3 75K 173289 2.0e-3 N/A 6672 101 37
+5 2 3 30K 372001 1.0e-3 N/A 3689 57 20 [3.57e-3, 3.81e-3]
+6 2 3 30K 372001 5.8e-6 N/A 2 37 21 [2.42e-7, 7.22e-6]
+6 2 3 30K 572091 5.8e-4 N/A 3889 58 21 [3.77e-3, 4.01e-3]
+8 2 3 50K 173289 4.0e-9 N/A 1 65 26 [2.53e-8, 5.57e-6]
+8 2 3 50K 371793 2.5e-5 N/A 487 71 28 [4.45e-4, 5.32e-4]
+10 2 3 75K 173289 2.0e-3 N/A 6672 101 37 [6.50e-3, 6.72e-3]
+12 2 3 75K 371793 3.4e-4 N/A 2175 96 32 [2.08e-3, 2.25e-3]
 Mixed models
-5 3 4 40K 372001 9.0e-9 0 0 45 21
-6 3 4 40K 372001 5.0e-9 1 0 54 26
-7 3 4 50K 372001 2.0e-8 2 6 108 40
-8 3 4 60K 173289 4.7e-8 0 7 123 45
-9 3 4 60K 173289 3.2e-7 1 33 140 46
-10 3 4 75K 173289 1.1e-6 2 295 143 53
-11 3 4 80K 572091 3.9e-8 0 13 138 50
-12 3 4 85K 572091 1.7e-8 2 10 167 55
-13 3 4 685K 572091 9.5e-6 399 4164 197 64
+5 3 4 40K 372001 9.0e-9 0 0 45 21 [0.00e+0, 3.69e-6]
+6 3 4 40K 372001 5.0e-9 1 0 54 26 [2.53e-8, 5.57e-6]
+7 3 4 50K 372001 2.0e-8 2 6 108 40 [3.45e-6, 1.58e-5]
+8 3 4 60K 173289 4.7e-8 0 7 123 45 [2.81e-6, 1.44e-5
+9 3 4 60K 173289 3.2e-7 1 33 140 46 [2.35e-5, 4.75e-5]
+10 3 4 75K 173289 1.1e-6 2 295 143 53 [2.65e-4, 3.53e-4]
+11 3 4 80K 572091 3.9e-8 0 13 138 50 [6.52e-6, 2.33e-5]
+12 3 4 85K 572091 1.7e-8 2 10 167 55 [8.68e-6, 2.72e-5]
+13 3 4 85K 572091 9.5e-6 399 4164 197 64 [4.41e-3, 4.70e-3]
 Mixed models initialized with addition model
-5 2 3 40K 572091 2.5e-7 1 52 49 20
-6 2 3 40K 572091 2.4e-8 0 5 57 21
-6 3 3 40K 572091 1.8e-8 0 3 70 35
-6 3 3 80K 572091 1.6e-8 0 3 75 35
-6 3 4 40K 372001 8.0e-9 0 0 72 26
-6 3 4 40K 173289 1.4e-8 3 2 60 29
-6 3 4 50K 572091 2.9e-8 0 4 79 29
-7 3 4 50K 572091 1.8e-8 4 1 104 38
-8 3 4 70K 572091 4.3e-5 50 1196 116 42
-9 3 4 70K 572091 5.4e-8 1 4 160 50
-10 3 3 50K 572091 6.3e-7 6 7 90 45
-11 3 4 75K 572091 5.9e-5 11066 1120 141 47
+5 2 3 40K 572091 2.5e-7 1 52 49 20 [5.35e-5, 8.48e-5]
+6 2 3 40K 572091 2.4e-8 0 5 57 21 [1.63e-6, 1.10e-5]
+6 3 3 40K 572091 1.8e-8 0 3 70 35 [4.25e-7, 7.88e-6]
+6 3 3 80K 572091 1.6e-8 0 3 75 35 [4.25e-7, 7.88e-6]
+6 3 4 40K 372001 8.0e-9 0 0 72 26 [0.00e+0, 3.69e-6]
+6 3 4 40K 173289 1.4e-8 3 2 60 29 [2.16e-6, 1.18e-5]
+6 3 4 50K 572091 2.9e-8 0 4 79 29 [8.11e-7, 9.28e-6]
+7 3 4 50K 572091 1.8e-8 4 1 104 38 [1.03e-6, 1.08e-5]
+8 3 4 70K 572091 4.3e-5 50 1196 116 42 [1.11e-3, 1.23e-3]
+9 3 4 70K 572091 5.4e-8 1 4 160 50 [8.11e-7, 9.28e-6]
+10 3 3 50K 572091 6.3e-7 6 7 90 45 [5.64e-6, 2.15e-5]
+11 3 4 75K 572091 5.9e-5 11066 1120 141 47 [1.19e-2, 1.22e-2]
 Mixed models initialized with add model. Reset useful heads every 100 steps
-6 4 4 40K 372001 1.7e-8 3 8 51 30
-Mixed models initialized with add model. Reset useful heads & MLPs every 100 steps
-6 4 3 40K 372001 3.0e-4 17 3120 115 53
+6 4 4 40K 372001 1.7e-8 3 8 51 30 [5.64e-6, 2.15e-5]
+Mixed models inited with add model. Reset useful hds & MLPs every 100 steps
+6 4 3 40K 372001 3.0e-4 17 3120 115 53 [3.06e-3, 3.30e-3]
 For each model the ’QuantaMathsTrain’ Colab notebook generates two files:
@@
   • A “behaviors.json” file containing generic “behavior” facts learnt about the model by the Colab e.g. P18L0H0 attends to tokens D3 and D’3
-  • A “features.json” file containing maths-specific ”feature” facts learnt about the model by the Colab e.g. P18L0H0 performs the A3.SC subtask.
-Appendix F Appendix: TriAdd Implementation
-TriAdd transfers data from $A_{n-1}$to $A_{n}$by integrating the values of $A_{n-1}.ST$ and $A_{n}.ST$ This function can be represented as nine bigram mappings with three possible outputs.
-(Refer Tab.8.)
-Table 8: $A_{n}$.TriAdd can be calculated from $A_{n}.ST$ and $A_{n-1}.ST$ through nine bigram mappings and yielding the three distinct outputs ST8, ST9 and ST10
-$A_{n}.ST$ $A_{n}.ST$ $A_{n}.ST$
-$A_{n-1}.ST$ = ST8 = ST9 = ST10
-ST8 ST8 ST9 ST10
-ST9 ST8 ST9 ST10
- ST10 ST8 * ST10 ST10
-Note that in the case $A_{n}.ST$ = ST9 and $A_{n-1}.ST$ = ST10, the answer is indeterminate.
-The result could be ST8 or ST9 but importantly it can not be ST10.
-We choose to use ST8 in our definition, but ST9 would work just as well.
-Appendix G Appendix: Complexity
+  • A “features.json” file containing maths-specific “feature” facts learnt about the model by the Colab e.g. P18L0H0 performs the SC3 subtask.
+Table 7: For a sample model, all nodes used in predictions are shown by token position (horizontally) and model layer (vertically), detailing the answer digits they impact.
+Here, the attention heads in token position P10 labelled A5..3 help predict the answer digits A3, A4 and A5.
+For all addition and mixed models studied, before the “=” token, each node often calculates data used to predict multiple answer digits.
+After the “=” token, all nodes in a given token position are used to predict a single answer digit.
+(P6) (P9) (P10) (P11) (P12) (P13) (P14) (P15) (P16) (P17)
+D’4 D’1 D’0 = + A5 A4 A3 A2 A1
+L0H0 A4 A2 A1
+L0H1 A5 A5..3 A5 A3 A0
+L0H2 A5 A5..1
+L0MLP A5..2 A4 A3 A2 A1 A0
+L1H0
+L1H2 A4 A3 A2
+L1MLP A1 A0
+Appendix I Appendix: Complexity
 To analyze question difficulty, we categorized addition questions by the complexity of the computation required to solve the question, as shown in Tab.
-9.
+8.
 The categories are arranged according to the number of digits that a carry bit has to cascade through.
@@
 Name Contains Example Freq
-S0 SA 11111+12345=23456 $\sim$5%
+S0 SASA 11111+12345=23456 $\sim$5%
 S1 SA,SC 11111+9=22230 $\sim$21%
@@
 S5 SA,SCx5 11111+88889=100000 $\sim$2%
-Appendix H Appendix: Addition Hypothesis 1
-Given the 2-layer attention pattern’s similarity to 1-layer attention pattern, and the above evidence, our first (incorrect) hypothesis was that the 2-layer algorithm:
-  • Is based on the same SA, SC and SS operations as the 1-layer.
-  • Uses the new early positions to (somehow) do the SS calculations with higher accuracy than the 1-layer model.
-  • The long double staircase still finalizes each answer digit’s calculation.
-  • The two attention nodes in the long double staircase positions do the SA and SC calculations and pull in SS information calculated in the early positions.
-If this is correct then the 2-layer algorithm successfully completes these calculations:
-  • A0 = A0.SA
-  • A1 = A1.SA + A0.SC
-  • A2 = A2.SA + (A1.SC or (A1.SS & A0.SC))
-  • A3 = A3.SA + (A2.SC or (A2.SS & A1.SC) or (A2.SS & A1.SS & A0.SC))
-  • A4 = A4.SA + (A3.SC or (A3.SS & A2.SC) or (A3.SS & A2.SS & A1.SC) or (A3.SS & A2.SS & A1.SS & A0.SC))
-  • A5 = A4.SC or (A4.SS & A3.SC) or (A4.SS & A3.SS & A2.SC) or (A4.SS & A3.SS & A2.SS & A1.SC) or (A4.SS & A3.SS & A2.SS & A1.SS & A0.SC)
-Our intuition is that there are not enough useful nodes in positions 8 to 11 to complete the A5 calculation this way. So we abandoned this hypothesis.
-Appendix I Appendix: Addition Hypothesis 2
-Our second (incorrect) hypothesis was that the 2-layer algorithm has a more compact data representation, so it can pack more calculations into each node, allowing it to accurately predict A5 in step 11.
-We claimed the model stores the sum of each digit-pair as a single token in the range “0” to “18” (covering 0+0 to 9+9). We name this operator $A_{n}$.T, where T stands for “token addition”:
-  • $A_{n}$.T = $D_{n}$+ $D^{\prime}_{n}$
-The $A_{n}$.T operation does not understand mathematical addition. Tab. 10 shows how the model implements the T operator as a bigram mapping.
-Table 10: Implementing the T operator as a bigram mapping from 2 input tokens to 1 result token.
-$D_{n}$vs $D^{\prime}_{n}$ 0 1 4 5 8 9
-0 0 1 4 5 8 9
-1 1 2 5 6 9 10
-4 4 5 8 9 12 13
-5 5 6 9 10 13 14
-8 8 9 12 13 16 17
-9 9 10 13 14 17 18
-$A_{n}$.T is a compact way to store data. Tab. 11 show how, if it needs to, the model can convert a $A_{n}$.T value into a one-digit-accuracy SA, SC or SS value.)
-Table 11: Converting a $A_{n}$.T value into a SA , SC or SS value.
-$A_{n}$.T $A_{n}$.SA $A_{n}$.SC $A_{n}$.SS
-0 0 0 0
-1 1 0 0
-8 8 0 0
-9 9 0 1
-10 0 1 0
-17 7 1 0
-18 8 1 0
-Our notation shorthand for one-digit-accuracy these “conversion” bigram mappings is:
-  • $A_{n}$.SA = ($A_{n}$.T % 10) where % is the modulus operator
-  • $A_{n}$.SC = ($A_{n}$.T // 10) where // is the integer division operator
-  • $A_{n}$.SS= ($A_{n}$.T == 9) where == is the equality operator
-The A0.T value is accurate.
-But the other $A_{n}$.T values are not accurate because each is constrained to information from just one digit.
-We define another more accurate operator $A_{n}$.T2 that has “two-digit accuracy”.
-$A_{n}$.T2 is the pair sum for the nth digit plus the carry bit (if any) from the n-1th digit T:
-  • $A_{n}$.T2 = $A_{n}$.T + $A_{n-1}$.SC
-$A_{n}$.T2 is more accurate than $A_{n}$.T.
-The $A_{n}$.T2 value is always in the range “0” to “19” (covering 0+0+0 to 9+9+CarryOne).
-Tab.
-12 show how the model can implement the T2 operator as a mapping.
-$A_{n}$.T $A_{n}$.T2 if $A_{n-1}$.SC==0 $A_{n}$.T2 if $A_{n-1}$.SC==1
-0 0 1
-1 1 2
-9 9 10
-10 10 11
-17 17 18
-18 18 19
-Table 12: Calculating $A_{n}$.T2 from $A_{n}$.T and $A_{n-1}$.T
-Following this pattern, we define operators $A_{n}$.T3, $A_{n}$.T4 and $A_{n}$.T5 with 3, 4 and 5 digit accuracy respectively:
-  • $A_{n}$.T3 = $A_{n}$.T + ( $A_{n-1}$.T2 // 10 ) aka $A_{n}$.T + $A_{n-1}$.SC2
-  • $A_{n}$.T4 = $A_{n}$.T + ( $A_{n-1}$.T3 // 10 ) aka $A_{n}$.T + $A_{n-1}$.SC3
-  • $A_{n}$.T5 = $A_{n}$.T + ( $A_{n-1}$.T4 // 10 ) aka $A_{n}$.T + $A_{n-1}$.SC4
-The value A4.T5 is accurate as it integrates SC and cascading SS data all the way back to and including A0.T.
-The values A1.T2, A2.T3, A3.T4 are also all accurate.
-If the model knows these values it can calculate answer digits accurately:
-  • A1 = A1.T2 % 10
-  • A2 = A2.T3 % 10
-  • A3 = A3.T4 % 10
-  • A4 = A4.T5 % 10
-  • A5 = A4.T5 // 10
-In this hypothesis, all the answer digits are accurately calculated using the nodes in positions 8 to 11.
-This hypothesis 2 is feasible, elegant and compact - reflecting the authors (human) values for good code.
-Experimenation shows the model does not implement this hypothesis. It retains the long staircase SA calculations in positions 11 to 16. Why? Two reasons suggest themselves:
-  • Hypothesis 2 is too compact.
-The model is not optimising for compactness.
-The long staircase is discovered early in training, and it works for simple questions.
-Once the overall algorithm gives low loss consistently it stops optimising.
-  • Hypothesis 2 accurately predicts all answer digits in step 11 - reflecting the authors (human) values for good code.
-The model is not motivated to do this.
-It just needs to accurately predict A5 as 1 or 0 in step 11 and A4 in step 12 - nothing more.
-We abandoned this hypothesis.
-Appendix J Appendix: Addition Hypothesis 3
-The hypothesis 3 pseudo-code was derived iteratively by obtaining experimental results and mapping them to mathematical operations. Some of the experiments and mappings were:
-  • Ablation experiments show that the A5 value is accurately calculated in prediction step 11 using 5 attention heads and 5 MLP layers.
-The pseudo-code accurately calculates A5 while constraining itself to this many steps.
-  • Ablating the nodes one by one shows which answer digit(s) are reliant on each node (Ref Table 1).
-Most interestingly, ablating P10.L0.H1 impacts the answer digits A5, A4, A3, A2 (but not A1 and A0).
-This node is used in the calculation of A5, A4, A3, A2 in prediction steps 11, 12, 13 and 14.
-These relationships are constraints that are all obeyed by the pseudo-code.
-  • The pseudo-code has 4 instances where $A_{n}.ST$ is calculated using TriCase.
-PCA of the corresponding nodes (P8.L0.H1, P9.L0.H1, P11.L0.H2 and P14.L0.H1) shows tri-state output for the specified $D_{n}$.
-(see Figure 9).
-  • The pseudo-code has 4 instances where compound functions using TriCase and TriAdd to generate tri-state outputs.
-PCA of the corresponding nodes (P11.L0.H1, P12.L0.H1 and P13.L0.H1) shows tri-state output for the specified $D_{n}$.
-(see Figure 9).
-  • Activation patching (aka interchange intervention) experiments at attention head level confirmed some aspects of the calculations (see § K for details.
-  • The pseudo code includes calculations like A1.ST which it says is calculated in P9.L0.H1 and P9.L0.MLP.
-Ablation tells us both nodes are necessary.
-For the attention head we use the PCA results for insights.
-We didn’t implement a similar investigative tool for the MLP layer, so in the pseudo-code we attribute the calculation of A1.ST to both nodes.
-  • For P10.L0.H1, the attention head PCA could represent either a bi-state or tri-state output (see Figure 10).
-The MLP layer at P10.L0.MLP could map the attention head output to either a bi-state or tri-state.
-We cannot see which.
-The pseudo-code shows a tri-state calculation at P10.L0.MLP, but with small alterations the pseudo-code would work with a bi-state output.
-  • For P15.L0.H1 the attention head PCA could represent either a bi-state or tri-state output (see Figure 10).
-The pseudo-code shows a bi-state calculation A0.SC at P15.L0.H1, but with small alterations the pseudo-code would work with a tri-state output.
-  • The calculation of A1.ST2 in P14.L0.H1 is a interesting case.
-The model needs A1.ST2 for A2 accuracy.
-The model could simply reuse the accurate A1.ST2 value calculated in P10.
-Activation patching shows that it does not.
-Instead the P14 attention heads calculate A1.ST1 from D1 and D’1 directly, and only relies on the P10.D1.ST2 value in the case where A1.ST2 != A1.ST.
-That is, the calculation is “use P14.A1.ST1 value else use A1.ST2 values”.
-This aligns with the model learning the P10.A1.ST calculation early in training (for 90% accuracy) and later learning that P10.A1.ST2 contains additional information it can use to get to >99.999% accuracy.
 Figure 10: For a sample 5-digit 2-layer 3-head addition model, this map shows a compacted view of all useful token positions (horizontally) and all useful attention heads and MLP layers (vertically) used in predictions as green cells.
-Each cell shows the simpliest (lowest complexity) quanta S0, S1, etc impacted when we ablate each node.
+Each cell shows the simplest (lowest complexity) quanta S0, S1, etc impacted when we ablate each node.
 To answer S0 questions, only the S0 nodes are used.
 To answer S1 questions, S0 and S1 nodes are used, etc.
-The model only uses nodes in nine token positions.
+The model only uses nodes in ten token positions.
 Figure 11: This map shows the % of enriched questions that fail when we ablate each node in a 5-digit 2-layer 3-head addition model.
-The model only uses nodes in token positions P8 to P16 (i.e.
-tokens D’2 to A1).
+The model only uses nodes in token positions P6 to P17 (i.e.
+tokens D’4 to A1).
 Lower percentages correspond to rarer edge cases.
@@
 Figure 12: This map shows the input tokens each attention head attends to at each token position in a 5-digit 2-layer 3-head addition model.
-At token position P11 the model predicts the first answer digit A5.
+At token position P12 the model predicts the first answer digit A5.
 All digit pairs (e.g.
-D2 D’2) are attended to by P11.
-Figure 9: For 5-digit addition, for these attention heads, for exactly 1 answer digit $A_{n}$each, PCA shows these interpretable results.
-The dot colours show the TriCase value of each question.
-The PCA data and TriCase quanta are both tri-state and strongly correlated.
-Figure 10: For 5-digit addition, for these attention heads, for exactly 1 digit $A_{n}$each, PCA shows these interpretable results.
-The dot colours show the TriCase value of each question.
-The PCA and TriCase data are strongly correlated, but the PCA data has 2 states.
-Appendix K Appendix: Addition Interchange Interventions
-To test the hypothesis 3 mapping of the mathematical framework (casual abstraction) to the model attention heads, various “interchange interventions” (aka activation patching) experiments were performed on the model, where
-  • A particular claim about an attention head has selected for testing.
-  • The model predicted answers for sample test questions, and the attention head activations were recorded (stored).
-  • The model then predicted answers for more questions, but this time we intervened during the prediction to override the selected attention head activations with the activations from the previous run.
-Using this approach we obtained the findings in Tab. 13.
-Nodes Claim: Attention head(s) perform … Finding: Attention head(s) perform a function that …
-P8.L0.H1 and MLP A2.ST = TriCase( D2, D’2 ) impacting A4 and A5 accuracy Based on D2 and D’2. Triggers on a A2 carry value. Provides carry bit used in A5 and A4 calculation.
-P9.L0.H1 and MLP A1.ST = TriCase( D1, D’1 ) impacting A5, A4 and A3 accuracy Based on D1 and D’1. Triggers on a A1 carry value. Provides carry bit used in A5, A4 and A3 calculation.
-P10.L0.H1 and MLP A1.ST2 = TriAdd( A1.ST, TriCase(D0, D’0) ) impacting A5, A4, A3 and A2 accuracy Based on D0 and D’0.
-Triggers on a A0 carry value.
-Provides carry bit used in A5, A4, A3 and A2 calculation.
-P11.L0.H1 and MLP A3.ST4 = TriAdd( TriCase( D3, D’3 ), TriAdd( A2.ST, A1.ST2)) impacting A5 accuracy Based on D3 and D’3. Triggers on a A3 carry value. Provides carry bit used in A5 calculations.
-P11.L0.H2 and MLP A4.ST = TriCase(D4, D’4) impacting A5 accuracy A4 Based on D4 and D’4. Triggers on a A4 carry value. Provides carry bit used in A5 calculation.
-P12.L0.H0+H2 and MLP A4.SA = (D4 + D’4) % 10 impacting A4 accuracy A4 Sums D4 and D’4. Impacts A4.
-P13.L0.H0+H2 and MLP A3.SA = (D3 + D’3) % 10 impacting A3 accuracy Sums D3 and D’3. Impacts A3.
-P14.L0.H0+H2 and MLP A2.SA = (D2 + D’2) % 10 impacting A2 accuracy Sums D2 and D’2. Impacts A2.
-P14.L0.H1 and MLP (D1 + D’1) / 10 + P10.A1.ST2 info impacting A2 accuracy Calculates P10.A1.ST1 but add P10.A1.ST2 info when A1.ST != A1.ST2. Impacts A2
-P15.L0.H0+H2 and MLP A1.SA = (D1 + D’1) % 10 impacting A1 accuracy Sums D1 and D’1. Impacts A1.
-P15.L0.H1 and MLP A0.SC = (D0 + D’0) / 10 impacting A1 accuracy Triggers when D0 + D’0 >10. Impacts A1 digit by 1
-P16.L0.H0+H2 and MLP A0.SA = (D0 + D’0) % 10 impacting A0 accuracy Sums D0 and D’0. Impacts A0.
-Table 13: Interchange Interventions experiments used activation patching to test the claims addition hypothesis 3 made for each attention head in a sample mixed model.
-Experimental results are consistent with hypothesis 3 for all nodes.
-+ve Sub Contains -ve Sub Contains Like
-M0 MD N/A N/A S0
-M1 MD,MB N1 ND,NB S1
-M2 MD,MBx2 N2 ND,NBx2 S2
-M3 MD,MBx3 N3 ND,NBx3 S3
-M4 MD,MBx4 N4 ND,NBx4 S4
-Table 14: We define “positive-answer subtraction” and “negative-answer subtraction” calculation complexity quanta that parallel the addition quanta.
-Appendix L Appendix: N-Digit Addition
+D2 D’2) are attended to by P12.
+Appendix J Appendix: N-Digit Addition
 The addition models perform addition accurately. Visualizations that provided insights into the behavior of the model, aiding our interpretation of the algorithm, are below:
@@
   • Per answer digit, some models use the $SC$ calculation, whereas some models optimize it out and rely solely on the $ST$ value (and so are more compact).
-Appendix M Appendix: Mixed Model Initialization
+Appendix K Appendix: Mixed Model Initialization
 We experimented with three approaches to re-using the trained addition model in the “mixed” (addition and subtraction) model:
@@
 Our intuition was that initializing the first layers and heads would be more likely to cause the model to re-use the addition circuits adding interpretability, so we used this approach.
-Appendix N Appendix: N-Digit Subtraction
+Appendix L Appendix: Seed Sensitivity
+An analysis of the sensitivity of 48 models to the initial seed was performed (This analysis excluded one model - the inaccurate 1-layer addition model that reproduces the Quirke and Barez [2024] paper results).
+Fig.
+15 shows the results.
+We conclude:
+  • The Addition models are the most stable - that is they are not sensitive to the seed value.
+  • The other categories (Subtraction, Mixed, Mixed+Init, and Mixed+Reset) show relatively low to moderate sensitivity.
+  • The higher average loss for Subtraction models show that the models find it harder to learn Subtraction is isolation.
+  • The higher average loss for Mixed+Reset models show that the this type of intervention during training makes it harder for the models to learn
+Figure 15: A visualization of the range of training losses across 48 models grouped by the different model categories.
+Appendix M Appendix: N-Digit Subtraction
 The mixed models perform addition and subtraction accurately. Visualizations that provided insights into the behavior of the model, aiding our interpretation of the algorithm, are below:
@@
 The model uses this heuristic even though this probabilistic information is sometimes incorrect and so will work against the model achieving very low loss.
+Appendix N Appendix: Addition Hypothesis 3
+The hypothesis 3 pseudo-code was derived iteratively by obtaining experimental results and mapping them to mathematical operations. Some of the experiments and mappings were:
+  • Ablation experiments show that the A5 value is accurately calculated in prediction step 11 using 5 attention heads and 5 MLP layers.
+The pseudo-code accurately calculates A5 while constraining itself to this many steps.
+  • Ablating the nodes one by one shows which answer digit(s) are reliant on each node (Ref Tab.
+7).
+Most interestingly, ablating P10.L0.H1 impacts the answer digits A5, A4, A3, A2 (but not A1 and A0).
+This node is used in the calculation of A5, A4, A3, A2 in prediction steps 11, 12, 13 and 14.
+These relationships are constraints that are all obeyed by the pseudo-code.
+  • The pseudo-code has 4 instances where $ST_{n}$ is calculated using TriCase.
+PCA of the corresponding nodes (P8.L0.H1, P9.L0.H1, P11.L0.H2 and P14.L0.H1) shows tri-state output for the specified $D_{n}$.
+(see Figure 7).
+  • The pseudo-code has 4 instances where compound functions using TriCase and TriAdd to generate tri-state outputs.
+PCA of the corresponding nodes (P11.L0.H1, P12.L0.H1 and P13.L0.H1) shows tri-state output for the specified $D_{n}$.
+(see Figure 7).
+  • Activation patching (aka interchange intervention) experiments at attention head level confirmed some aspects of the calculations.
+  • The pseudo code includes calculations like ST1 which it says is calculated in P9.L0.H1 and P9.L0.MLP.
+Ablation tells us both nodes are necessary.
+For the attention head we use the PCA results for insights.
+We didn’t implement a similar investigative tool for the MLP layer, so in the pseudo-code we attribute the calculation of ST1 to both nodes.
+  • For P10.L0.H1, the attention head PCA could represent either a bi-state or tri-state output.
+The MLP layer at P10.L0.MLP could map the attention head output to either a bi-state or tri-state.
+We cannot see which.
+The pseudo-code shows a tri-state calculation at P10.L0.MLP, but with small alterations the pseudo-code would work with a bi-state output.
+  • For P15.L0.H1 the attention head PCA could represent either a bi-state or tri-state output.
+The pseudo-code shows a bi-state calculation SC0 at P15.L0.H1, but with small alterations the pseudo-code would work with a tri-state output.
+  • The calculation of ST2 in P14.L0.H1 is a interesting case.
+The model needs ST2 for A2 accuracy.
+The model could simply reuse the accurate ST2 value calculated in P10.
+Activation patching shows that it does not.
+Instead the P14 attention heads calculate ST1 from D1 and D’1 directly, and only relies on the P10.D1.ST2 value in the case where ST2 != ST1.
+That is, the calculation is “use P14.ST1 value else use ST2 value”.
+This aligns with the model learning the P10.ST1 calculation early in training (for 90% accuracy) and later learning that P10.ST2 contains additional information it can use to get to >99.999% accuracy.